### v2

# A vectorial Darboux transformation for integrable matrix versions of the Fokas-Lenells equation

Abstract
Using bidifferential calculus, we derive a vectorial binary Darboux transformation for an integrable matrix version of the first negative flow of the Kaup-Newell hierarchy. A reduction from the latter system to an integrable matrix version of the Fokas-Lenells equation is then shown to inherit a corresponding vectorial Darboux transformation. Matrix soliton solutions are derived from the trivial seed solution. Furthermore, the Darboux transformation is exploited to determine in a systematic way exact solutions of the two-component vector Fokas-Lenells equation on a plane wave background. This comprises breathers, dark solitons, rogue waves and “beating solitons”.
1 Introduction
By a transformation of variables, the original Fokas-Lenells (FL) equation [1, 2] attains the simpler form [3]
$u_{xt}-u-2\mathrm{i}\,|u|^{2}u_{x}=0\,$, (1.1)
where $u$ is a complex function of independent real variables $x$ and $t$, and a subscript $x$ or $t$ indicates a partial derivative with respect to $x$, respectively $t$. In particular, it models propagation of ultrashort light pulses in birefringend optical fibers, taking certain nonlinear effects into account. For some integrability features of (1.1), see [4] and the references cited there. The FL equation (1.1) possesses a generalization to integrable matrix versions
$u_{xt}-u-\mathrm{i}\,(u\,\sigma_{1}u^{\dagger}\sigma_{2}u_{x}+u_{x}\sigma_{1}u^{\dagger}\sigma_{2}u)=0\,.$ (1.2)
Here $u$ is now a matrix of any size, where the components are functions of $x$ and $t$, $u^{\dagger}$ denotes the conjugate transpose of $u$, and $\sigma_{1},\sigma_{2}$ are any constant involutary Hermitian matrices of suitable size. In the case, where $u$ is a vector, (1.2) already appeared in [5] (see (82) therein), also see [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] for the two-component vector case, and [17] for the three-component case. Some results on the $2\times 2$ matrix case have been obtained in [18].
In this work, we present a vectorial Darboux transformation for (1.2), which allows to generate exact solutions from a given “seed” solution $u$. It involves a linear system that depends on $u$ in such a way that its compatibility (or integrability) condition is satisfied if $u$ solves (1.2), and the dependent variables of the linear system are $n$-component vectors, where $n$ may be thought of as corresponding to the number of solitons in the generated class of solutions of (1.2) (though already some $n=1$ solutions can be regarded as being composed of more elementary solutions). For $n=1$, the linear system reduces to a Lax pair, depending on a “spectral parameter”. In our linear system, this parameter is promoted to a “spectral matrix”.
The original Darboux transformation [19] has to be iterated in order to generate multiple soliton solutions of an integrable equation. Moreover, in each step a different spectral parameter value has to be used. Accordingly, important classes of solutions are only obtained indirectly by taking special limits, where a priori different spectral parameters tend to the same value. This concerns “multiple pole solutions” (in the language of the inverse scattering method), which include “positons” of the Korteweg-deVries and related equations, for which V.B. Matveev set up a generalized Darboux transformation [20, 21]. Later a corresponding generalization of the classical Darboux transformation has also been formulated for the Nonlinear Schrödinger (NLS) and related integrable equations in order to generate higher order rogue waves and (multi-) dark soliton solutions (see, e.g., [22, 23], and [7, 8] for applications to the case of the two-component vector FL equation).
A vectorial Darboux transformation does not require iteration in order to generate multi-soliton solutions, but does the job in a single step. Furthermore, it offers a direct way to the abovementioned important classes of solutions. [^1] The “Cauchy matrix approach”, applied in [24] to the scalar FL equation, may be regarded as a vectorial Darboux transformation restricted to the case of trivial seed. Moreover, it allows us, at least to some extent and in a tractable way, to exhaust the various possibilities of exact solutions within this framework. A beautiful aspect of a vectorial Darboux transformation is a simple superposition rule on the level of data (consisting of a “spectral matrix” and the corresponding solution of the linear system) [^2] In case of a vectorial binary Darboux transformation, we are dealing with two spectral matrices and two linear systems. determining a solution of the integrable nonlinear equation for a given seed solution, see Remark 4.4 below.
The usefulness of a vectorial version of the classical Darboux transformation has also been demonstrated by the “non-recursive Darboux transformation” in [25], which has been applied in [11] to construct higher order rogue wave solutions of the two-component vector FL equation. It should be noted, however, that it lacks the, in our opinion important, step of promoting the spectral parameter more consequently to a matrix and thus misses the superposition rule for solutions with the same seed. In any case, the “non-recursive Darboux transformation” is a step towards a more complete vectorial Darboux transformation, as proposed earlier by Mañas [26] for the NLS equation.
A key feature of a vectorial Darboux transformation is the appearance of a Sylvester equation for the so-called Darboux potential $\Omega$. The input for this algebraic matrix equation is the spectral matrix and a matrix built from the solution of the linear system. When the spectral matrix is diagonal and the Sylvester equation has a unique $n\times n$ matrix solution, it is expected that the generated solution of the respective nonlinear integrable equation can also be reached by an $n$-fold iteration of a classical Darboux transformation (using different spectral parameters). Multiple pole solutions and higher order rogue waves are obtained by choosing the spectral matrix as a Jordan block, and the nature of the generated class of solutions depends on (properties of) its eigenvalue.
Important classes of solitons of NLS-type equations are obtained, however, via the complementary case, where the Sylvester equation does not (completely) determine the Darboux potential and imposes constraints on the solution of the respective linear system. [^3] That the solution of the linear system can indeed be reduced in such a way that these constraints are satisfied, appears to be a bit of a miracle. Vectorial Darboux transformations, like those obtained via bidifferential calculus [27] (also see, in particular, [28, 29]), come along with additional differential equations that completely determine the Darboux potential in such a “degenerate case”. Corresponding solution classes will be worked out explicitly for the two-component vector FL equation.
In the present work, we exploit the vectorial Darboux transformation (only) with seed solutions, for which the linear system can be turned into a system of linear PDEs with constant coefficients. Then, generically, the general solution is a linear combination of exponentials of a linear expression in $x$ and $t$. But when the corresponding characteristic equation has a multiple root, the linear system “degenerates” and the general solution also involves a polynomial (or polynomials) of the independent variables. If the seed solution is an exponential describing a plane wave, and if we have a root of maximal order, then the resulting solution of the vector FL equation (or, more generally, some NLS-type equation) is quasi-rational, i.e., a rational expression multiplied by the plane wave exponential. If the norm of such a solution is localized on the uniform background, it represents a rogue wave (or freak wave), a case that attracted enormous interest, since it describes a wave that suddenly appears (“from nowhere”) and then disappears again.
There is another, independent mechanism, how rational expressions can enter the solution stage. This is at work in those cases, where the Sylvester equation does not possess a unique solution, which is true if and only if the spectral matrix does not satisfy the so-called spectrum condition.
Though essentially self-contained, this work is based on [4] (which, in turn, partly draws from [24]), where a vectorial Darboux transformation for the scalar FL equation has been derived, using bidifferential calculus. In Section 2 we show that, by choosing a suitable (realization of the) bidifferential calculus, a simple “Miura transformation” equation (see (2.1)) is essentially turned into a matrix generalization of the first “negative flow” of the Kaup-Newell (KN) hierarchy,
$u_{xt}-u-\mathrm{i}\,(u\,v\,u_{x}+u_{x}\,v\,u)=0\,,\qquad v_{xt}-v+\mathrm{i}\,(v\,u\,v_{x}+v_{x}\,u\,v)=0\,$, (1.3)
of which (1.2) is a reduction, via
$v=\sigma_{1}u^{\dagger}\sigma_{2}\,.$ (1.4)
The crucial point is that the Miura transformation equation possesses a vectorial binary Darboux transformation (bDT). A bDT allows to generate from a given (“seed”) solution a new family of solutions, using solutions of a linear system and also of another (“adjoint”) linear system. The bDT for (1.3) is derived in Section 3.
A difficult task is to find an extension of the reduction (1.4) such that the bDT generates solutions of the matrix FL equation (1.2), i.e., that it preserves the reduction condition (1.4). This problem is solved in Section 4, essentially following the steps in the scalar case [4]. In Section 4 we also derive (multi-) matrix soliton solutions from the trivial solution $u=0$.
Section 5 deals with the vector version of the matrix FL equation, obtained when the matrix $u$ in (1.2) reduces to a vector (also see [5, 6, 7, 8, 9, 17, 11, 14, 16]). In Section 6, we finally concentrate on the two-component vector case. The vectorial Darboux transformation is used to generate all types of soliton solutions on a plane wave background solution in a systematic way.
Finally, Section 7 contains some concluding remarks.
2 From the Miura transformation in bidifferential calculus to matrix versions of the first negative flow of the KN hierarchy
A graded associative algebra is an associative algebra $\boldsymbol{\Omega}=\bigoplus_{r\geq 0}\boldsymbol{\Omega}^{r}$ over a field $\mathbb{K}$ of characteristic zero, where $\mathcal{A}=\boldsymbol{\Omega}^{0}$ is an associative algebra over $\mathbb{K}$ and $\boldsymbol{\Omega}^{r}$, $r\geq 1$, are $\mathcal{A}$-bimodules such that $\boldsymbol{\Omega}^{r}\,\boldsymbol{\Omega}^{s}\subseteq\boldsymbol{\Omega}^{r+s}$. Elements of $\boldsymbol{\Omega}^{r}$ are called $r$-forms. A bidifferential calculus is a unital graded associative algebra $\boldsymbol{\Omega}$, supplied with two $\mathbb{K}$-linear graded derivations $\mathrm{d},\bar{\mathrm{d}}:\boldsymbol{\Omega}\rightarrow\boldsymbol{\Omega}$ of degree one (so that $\mathrm{d}\boldsymbol{\Omega}^{r}\subseteq\boldsymbol{\Omega}^{r+1}$, $\bar{\mathrm{d}}\boldsymbol{\Omega}^{r}\subseteq\boldsymbol{\Omega}^{r+1}$), and such that
$\mathrm{d}^{2}=0\,,\qquad\bar{\mathrm{d}}^{2}=0\,,\qquad\mathrm{d}\bar{\mathrm{d}}+\bar{\mathrm{d}}\mathrm{d}=0\,.$
In the following, we show that the matrix versions (1.3) of the first negative flow of the KN hierarchy arise from the simple equation
$\mathrm{d}\phi=(\bar{\mathrm{d}}g)\,g^{-1}\,$, (2.1)
for a special realization of the bidifferential calculus. The interest in this equation stems from the fact that it possesses a bDT [4], which will be recalled in Section 3. It allows to generate from a given solution $(\phi,g)$ a family of solutions $(\phi^{\prime},g^{\prime})$, depending on a number of parameters.
Let $\mathcal{A}$ now be the associative algebra of smooth functions of independent real variables $x$ and $t$. We choose $\boldsymbol{\Omega}=\mathrm{Mat}(\mathcal{A})\otimes\bigwedge(\mathbb{C}^{2})$ and a basis $\xi_{1},\xi_{2}$ of the Grassmann algebra $\bigwedge(\mathbb{C}^{2})$. $\mathrm{Mat}(\mathcal{A})$ denotes the algebra of all matrices over $\mathcal{A}$, where the product of two matrices is set to zero if the sizes do not match. Defining $\mathrm{d}$ and $\bar{\mathrm{d}}$ on $\mathrm{Mat}(\mathcal{A})$, they extend in an obvious way to $\Omega$ (treating $\xi_{1},\xi_{2}$ as “constants” with respect to $\mathrm{d}$ and $\bar{\mathrm{d}}$).
For each $n>1$, let $J_{n}$ be a constant $n\times n$ matrix, $J_{n}\neq I_{n}$, where $I_{n}$ denotes the $n\times n$ identity matrix. For an $m\times n$ matrix $F$ over $\mathcal{A}$, we set
$\mathrm{d}F=F_{x}\,\xi_{1}+\frac{1}{2}(J_{m}F-FJ_{n})\,\xi_{2}\,,\qquad\bar{\mathrm{d}}F=\frac{1}{2}(J_{m}F-FJ_{n})\,\xi_{1}+F_{t}\,\xi_{2}$
(also see [30, 4]). In the following, we fix $m>1$. Let $\phi$ and $g$ be $m\times m$ matrices over $\mathcal{A}$, $g$ invertible. We write $J$ instead of $J_{m}$. Then the Miura (transformation) equation (2.1) becomes the “Miura system”
$\phi_{x}=\frac{1}{2}\,[J,g]\,g^{-1}\,,\qquad g_{t}g^{-1}=\frac{1}{2}\,[J,\phi]\,.$ (2.2)
Proposition 2.1.
(2.2) implies that $\det(g)$ does not depend on $t$.
Proof.
In terms of $\tilde{g}=\det(g)^{-1/m}\,g$, so that $\det(\tilde{g})=1$, (2.2) becomes
$\phi_{x}=\frac{1}{2}\,[J,\tilde{g}]\,\tilde{g}^{-1}\,,\qquad\tilde{g}_{t}\tilde{g}^{-1}+\frac{1}{m}(\log\det(g))_{t}\,I_{m}=\frac{1}{2}\,[J,\phi]\,.$
By taking the trace of the second equation, we obtain $\det(g)_{t}=0$. ∎
Choosing
$J=\mbox{block-diagonal}(I_{m_{1}},-I_{m_{2}})\,,\qquad m=m_{1}+m_{2}\,$,
and decomposing the matrices $\phi$ and $g$ accordingly,
$\phi=\left(\begin{array}[]{cc}\phi_{11}&\phi_{12}\\ \phi_{21}&\phi_{22}\end{array}\right)\,,\qquad g=\left(\begin{array}[]{cc}g_{11}&g_{12}\\ g_{21}&g_{22}\end{array}\right)\,$,
assuming invertibility of $g_{22}$ and also of the Schur complement of $g_{22}$ in $g$,
$S=g_{11}-g_{12}\,g_{22}^{-1}g_{21}\,$,
we have
$g^{-1}=\left(\begin{array}[]{cc}S^{-1}&-S^{-1}g_{12}\,g_{22}^{-1}\\ -g_{22}^{-1}g_{21}S^{-1}&g_{22}^{-1}\big(I_{m_{2}}+g_{21}S^{-1}g_{12}\,g_{22}^{-1}\big)\end{array}\right)\,$,
and the Miura system (2.2) takes the form
$\left(\begin{array}[]{cc}g_{11}&g_{12}\\ g_{21}&g_{22}\end{array}\right)_{t}=\left(\begin{array}[]{cc}\phi_{12}\,g_{21}&\phi_{12}\,g_{22}\\ -\phi_{21}\,g_{11}&-\phi_{21}\,g_{12}\end{array}\right)\,$, (2.9)
$\left(\begin{array}[]{cc}\phi_{11}&\phi_{12}\\ \phi_{21}&\phi_{22}\end{array}\right)_{x}=\left(\begin{array}[]{cc}-g_{12}\,g_{22}^{-1}\,g_{21}S^{-1}&g_{12}\,g_{22}^{-1}\big(I_{m_{2}}+g_{21}S^{-1}g_{12}\,g_{22}^{-1}\big)\\ -g_{21}S^{-1}&g_{21}S^{-1}g_{12}\,g_{22}^{-1}\end{array}\right)\,.$ (2.14)
Remark 2.2.
The Miura system, (2.9) and (2.14), is invariant under $g\mapsto f\,g$ with any nowhere vanishing function $f(x)$. Any physical equation derived from it, should therefore only involve quotients of components of $g$. According to Proposition 2.1, for each solution we can arrange that $\det(g)=1$ by a scaling. $\Box$
Let us introduce
$u=\phi_{21}\,,\qquad v=\mathrm{i}\,g_{12}\,g_{22}^{-1}\,$, (2.15)
which are matrices of size $m_{2}\times m_{1}$, respectively $m_{1}\times m_{2}$. The following proposition generalizes a result of [24, 4] from the scalar case to the arbitrary size matrix case.
Proposition 2.3.
As a consequence of (2.9) and (2.14), $u$ and $v$ satisfy the corresponding matrix version (1.3) of the first negative flow of the KN hierarchy.
Proof.
The equation for $\phi_{21x}$ in (2.14) reads
$u_{x}=-g_{21}S^{-1}\,.$ (2.16)
With its help, we can write $S=g_{11}-\mathrm{i}\,v\,u_{x}\,S$, so that
$g_{11}=(I_{m_{1}}+\mathrm{i}\,v\,u_{x})\,S\,.$
Using (2.9), we obtain
$S_{t}=-\mathrm{i}\,v\,u\,S\,.$
Now we have
$u_{xt}=-g_{21t}S^{-1}+g_{21}S^{-1}S_{t}S^{-1}=u\,g_{11}S^{-1}+\mathrm{i}\,u_{x}\,v\,u=u+\mathrm{i}\,\big(u\,v\,u_{x}+u_{x}\,v\,u\big)\,$,
and
$v_{tx}$ $=$ $\mathrm{i}\,\big(g_{12t}\,g_{22}^{-1}-g_{12}g_{22}^{-1}g_{22t}\,g_{22}^{-1}\big)_{x}=\mathrm{i}\,\big(\phi_{12}-\mathrm{i}\,v\,u\,g_{12}\,g_{22}^{-1}\big)_{x}=\mathrm{i}\,\big(\phi_{12}-v\,u\,v\big)_{x}$
$=$ $v-\mathrm{i}\,(v_{x}\,u\,v+v\,u\,v_{x}\big)\,$,
where we used
$\phi_{12x}=g_{12}\,g_{22}^{-1}\big(I_{m_{2}}+g_{21}S^{-1}g_{12}\,g_{22}^{-1}\big)=-\mathrm{i}\,v\,(I_{m_{2}}+\mathrm{i}\,u_{x}v)\,.$
Moreover, any solution of (1.3) determines a solution of the Miura system.
Lemma 2.4.
Let $(u,v)$ be a solution [^4] A solution may not be defined everywhere on $\mathbb{R}^{2}$. of (1.3). Then
$g_{22t}=\mathrm{i}\,u\,v\,g_{22}\,,\qquad g_{12}=-\mathrm{i}\,v\,g_{22}\,,\qquad g_{21}=-u_{x}S\,,\qquad g_{11}=(I_{m_{1}}+\mathrm{i}\,v\,u_{x})\,S\,$,
$\phi_{21}=u\,,\qquad\phi_{12}=-\mathrm{i}\,v_{t}+v\,u\,v\,,\qquad\phi_{11}=-\mathrm{i}\,\int v\,u_{x}\,dx\,,\qquad\phi_{22}=\mathrm{i}\,\int u_{x}v\,dx\,$,
where $S$ is an invertible solution of
$S_{t}=-\mathrm{i}\,v\,u\,S\,$,
determines a solution of the Miura system, (2.9) and (2.14).
Proof.
The linear first order matrix ODEs for $g_{22}$ and $S$ possess invertible solutions (see [31], for example), wherever $u$ and $v$ are smooth. The statement of the lemma is then easily verified. ∎
3 A vectorial binary Darboux transformation for matrix versions of the first negative flow of the KN hierarchy
We recall the following result of bidifferential calculus [4].
Theorem 3.1.
Given a bidifferential calculus with maps $\mathrm{d},\bar{\mathrm{d}}$, let 0-forms $\Delta,\Gamma$ and 1-forms $\kappa,\lambda$ satisfy
$\bar{\mathrm{d}}\Delta+[\lambda,\Delta]=(\mathrm{d}\Delta)\,\Delta\,,\qquad\bar{\mathrm{d}}\lambda+\lambda^{2}=(\mathrm{d}\lambda)\,\Delta\,$,
$\bar{\mathrm{d}}\Gamma-[\kappa,\Gamma]=\Gamma\,\mathrm{d}\Gamma\,,\qquad\bar{\mathrm{d}}\kappa-\kappa^{2}=\Gamma\,\mathrm{d}\kappa\,.$ (3.1)
Let 0-forms $\phi$ and $g$, the latter assumed to be invertible, be a solution of the Miura equation (2.1) and $\theta,\eta$ solutions of the linear equations
$\bar{\mathrm{d}}\theta=(\mathrm{d}\phi)\,\theta+(\mathrm{d}\theta)\,\Delta+\theta\,\lambda\,,\qquad\bar{\mathrm{d}}\eta=-\eta\,\mathrm{d}\phi+\Gamma\,\mathrm{d}\eta+\kappa\,\eta\,.$ (3.2)
Furthermore, let $\Omega$ be an invertible solution of the linear equations
$\Gamma\,\Omega-\Omega\,\Delta=\eta\,\theta\,$, (3.3)
$\bar{\mathrm{d}}\Omega=(\mathrm{d}\Omega)\,\Delta-(\mathrm{d}\Gamma)\,\Omega+\kappa\,\Omega+\Omega\,\lambda+(\mathrm{d}\eta)\,\theta\,.$ (3.4)
Then, if $\Omega$ is invertible,
$\phi^{\prime}=\phi-\theta\,\Omega^{-1}\eta+C\,,\qquad g^{\prime}=g-\theta\,\Omega^{-1}\Gamma^{-1}\eta\,g\,$, (3.5)
where $\mathrm{d}C=0$, also solve the Miura equation (2.1). $\Box$
Elaborating this result for the choice of bidifferential calculus in Section 2, leads to the following theorem. [^5] Here we restrict our considerations to a framework of finite-dimensional matrices. But under suitable technical assumptions, we may let $n$ in Theorem 3.2 to be infinite, or better generalize the matrices involving the matrix size $n$ to become (bounded) operators on a Banach space. Also see [26], for example.
Theorem 3.2.
Let $\Delta$ and $\Gamma$ be invertible constant $n\times n$ matrices. Let $(u,v)$ (where $u$ and $v$ have matrix size $m_{2}\times m_{1}$, respectively $m_{1}\times m_{2}$) be a solution of the matrix version (1.3) of the first negative flow of the KN hierarchy on an open set $\mathcal{U}\subset\mathbb{R}^{2}$. Furthermore, let $\theta_{1},\theta_{2}$ be $m_{1}\times n$, respectively $m_{2}\times n$, matrix solutions of the linear system
$\theta_{1x}\,\Delta-(\frac{1}{2}I_{m_{1}}+\mathrm{i}\,vu_{x})\,\theta_{1}\,-\mathrm{i}\,v\,(I_{m_{2}}+\mathrm{i}\,u_{x}v)\,\theta_{2}=0\,,\qquad\theta_{2x}\,\Delta+(\frac{1}{2}I_{m_{2}}+\mathrm{i}\,u_{x}v)\,\theta_{2}\,+u_{x}\,\theta_{1}=0\,$,
$\theta_{1t}-\frac{1}{2}\theta_{1}\Delta+(\mathrm{i}\,v_{t}-vuv)\,\theta_{2}=0\,,\qquad\theta_{2t}+\frac{1}{2}\theta_{2}\Delta+u\,\theta_{1}=0\,$, (3.6)
and $\eta_{1},\tilde{\eta}_{2}$ matrix solutions of size $n\times m_{1}$, respectively $n\times m_{2}$ of the adjoint linear system
$\Gamma\,\eta_{1x}+\frac{1}{2}\eta_{1}-\tilde{\eta}_{2}\,u_{x}=0\,,\qquad\Gamma\,\tilde{\eta}_{2x}-\frac{1}{2}\tilde{\eta}_{2}+\mathrm{i}\,\Gamma\eta_{1}v_{x}=0\,$,
$\eta_{1t}+\frac{1}{2}\Gamma\,\eta_{1}-\mathrm{i}\,\eta_{1}\,v\,u-\tilde{\eta}_{2}\,u=0\,,\qquad\tilde{\eta}_{2t}-\frac{1}{2}\Gamma\,\tilde{\eta}_{2}+\mathrm{i}\,\tilde{\eta}_{2}\,u\,v-\mathrm{i}\,\Gamma\eta_{1}v=0\,.$ (3.7)
Furthermore, let $\Omega$ be an $n\times n$ matrix solution of the Sylvester equation
$\Gamma\,\Omega-\Omega\,\Delta=\eta_{1}\,(\theta_{1}+\mathrm{i}\,v\,\theta_{2})+\tilde{\eta}_{2}\,\theta_{2}\,$, (3.8)
and the linear differential equations
$\Omega_{x}\Delta=-\eta_{1x}(\theta_{1}+\mathrm{i}\,v\,\theta_{2})-(\mathrm{i}\,\eta_{1}v_{x}+\tilde{\eta}_{2x})\,\theta_{2}\,,\qquad\Omega_{t}=-\frac{1}{2}\eta_{1}(\theta_{1}-\mathrm{i}\,v\,\theta_{2})+\frac{1}{2}\tilde{\eta}_{2}\,\theta_{2}\,.$ (3.9)
Then, in any open subset of $\mathcal{U}$ where $\Omega$ is invertible,
$u^{\prime}=u-\theta_{2}\,\Omega^{-1}\eta_{1}\,$, (3.10)
$v^{\prime}=\big(v-\mathrm{i}\,\theta_{1}(\Gamma\Omega)^{-1}\tilde{\eta}_{2}\big)\big(I_{m_{2}}-\theta_{2}(\Gamma\Omega)^{-1}\tilde{\eta}_{2}\big)^{-1}\,$, (3.11)
solve (1.3).
Proof.
Writing $\kappa=\kappa_{1}\xi_{1}+\kappa_{2}\xi_{2}$, $\lambda=\lambda_{1}\xi_{1}+\lambda_{2}\xi_{2}$, choosing $\kappa_{1}=\frac{1}{2}J_{n}$, $\kappa_{2}=-\frac{1}{2}\Gamma J_{n}$, $\lambda_{1}=-\frac{1}{2}J_{n}$, $\lambda_{2}=\frac{1}{2}J_{n}\Delta$, and using
$\mathrm{d}\phi=\phi_{x}\,\xi_{1}+\frac{1}{2}[J,\phi]\,\xi_{2}\,,\qquad\frac{1}{2}[J,\phi]=\left(\begin{array}[]{cc}0&\phi_{12}\\ -\phi_{21}&0\end{array}\right)\,$,
(3.2) takes the form
$\frac{1}{2}J\,\theta=\phi_{x}\,\theta+\theta_{x}\Delta\,,\qquad\theta_{t}=\frac{1}{2}[J,\phi]\,\theta+\frac{1}{2}J\,\theta\Delta\,$,
$\frac{1}{2}\eta\,J=\eta\,\phi_{x}-\Gamma\eta_{x}\,,\qquad\eta_{t}=-\frac{1}{2}\eta\,[J,\phi]-\frac{1}{2}\Gamma\eta\,J\,.$
Writing further
$\theta=\left(\begin{array}[]{c}\theta_{1}\\ \theta_{2}\end{array}\right)\,,\qquad\eta=\left(\begin{array}[]{cc}\eta_{1}&\eta_{2}\end{array}\right)\,$,
with an $m_{1}\times n$ matrix $\theta_{1}$, an $m_{2}\times n$ matrix $\theta_{2}$, an $n\times m_{1}$ matrix $\eta_{1}$ and an $n\times m_{2}$ matrix $\eta_{2}$, and using the solution of the Miura system corresponding to $(u,v)$ according to Lemma 2.4, the last system becomes (3.6) and
$\Gamma\,\eta_{1x}+\eta_{1}\,(\frac{1}{2}I_{m_{1}}+\mathrm{i}\,vu_{x})-\eta_{2}u_{x}=0\,,\qquad\Gamma\,\eta_{2x}-\eta_{2}\,(\frac{1}{2}I_{m_{2}}+\mathrm{i}\,u_{x}v)+\mathrm{i}\,\eta_{1}\,v\,(I_{m_{2}}+\mathrm{i}\,u_{x}v)=0\,$,
$\eta_{1t}+\frac{1}{2}\Gamma\,\eta_{1}-\eta_{2}\,u=0\,,\qquad\eta_{2t}-\frac{1}{2}\Gamma\,\eta_{2}-\eta_{1}(\mathrm{i}\,v_{t}-vuv)=0\,.$ (3.14)
Introducing
$\tilde{\eta}_{2}=\eta_{2}-\mathrm{i}\,\eta_{1}v\,$,
this is turned into (3.7). The conditions in (3.1) amount to $\Delta$ and $\Gamma$ being constant. (3.4) takes the form
$\Omega_{x}\Delta=-\eta_{x}\theta\,,\qquad\Omega_{t}=-\frac{1}{2}\eta\,J\,\theta\,$,
which is (3.9). To derive the second equation, we used the Sylvester equation. The statement of the theorem now follows from Theorem 3.1, which says that
$\phi^{\prime}=\phi-\theta\,\Omega^{-1}\eta+C\,,\qquad g^{\prime}=(I_{m}-\theta\,\Omega^{-1}\Gamma^{-1}\eta)\,g\,$,
where $C$ is any $\mathrm{d}$-constant $2\times 2$ block-matrix, solve the Miura equation, which via Proposition 2.3 implies that $u^{\prime}=\phi^{\prime}_{21}$ and
$v^{\prime}=\mathrm{i}\,g^{\prime}_{12}\,{g^{\prime}}_{22}^{-1}$
$=$ $\mathrm{i}\,\big((I_{m_{1}}-\theta_{1}\Omega^{-1}\Gamma^{-1}\eta_{1})\,g_{12}-\theta_{1}\Omega^{-1}\Gamma^{-1}\eta_{2}\,g_{22}\big)\big((I_{m_{2}}-\theta_{2}\Omega^{-1}\Gamma^{-1}\eta_{2})\,g_{22}-\theta_{2}\Omega^{-1}\Gamma^{-1}\eta_{1}\,g_{12}\big)^{-1}$
solve (1.3). Noting that $C$ is $\mathrm{d}$-constant if and only if $C=\mbox{block-diag}(c_{1},c_{2})$ with $c_{ix}=0$, $i=1,2$, and using (2.15), we arrive at (3.10) and (3.11). ∎
Remark $3.$
(3.6) and (3.7) are invariant under
$\theta_{i}\mapsto\theta_{i}S\,,\qquad\eta_{i}\mapsto T\eta_{i}\qquad i=1,2\,$,
$\Delta\mapsto S^{-1}\Delta S\,,\qquad\Gamma\mapsto T\Gamma T^{-1}\,$,
with any constant invertible $n\times n$ matrices $S$ and $T$. As a consequence, without restriction of generality, we may assume that $\Delta$ and $\Gamma$ are in (possibly different, upper or lower triangular) Jordan normal form. $\Box$
4 A vectorial Darboux transformation for the matrix FL equation
As mentioned in the introduction, (1.3) reduces to the matrix FL equation (1.2) by imposing the reduction condition (1.4). In this section, we extend this reduction to the vectorial binary Darboux transformation, expressed in Theorem 3.2, in such a way that a vectorial Darboux transformation for the matrix FL equation (1.2) is obtained.
Theorem 4.1.
Let $\Gamma$ be an invertible constant $n\times n$ matrix, such that $\Gamma$ satifies the “spectrum condition”, i.e., $\Gamma$ and $\Gamma^{\dagger}$ have no common eigenvalue. Let $u$ be a solution of the $m_{2}\times m_{1}$ matrix Fokas-Lenells equation (1.2) on an open set $\mathcal{U}\subset\mathbb{R}^{2}$. Furthermore, let $\eta_{1}$ and $\tilde{\eta}_{2}$ be $n\times m_{1}$, respectively $n\times m_{2}$, matrix solutions of the linear system
$\Gamma\,\eta_{1x}+\frac{1}{2}\eta_{1}-\tilde{\eta}_{2}\,u_{x}=0\,,\qquad\quad\eta_{1t}+\frac{1}{2}\Gamma\eta_{1}-\mathrm{i}\,\eta_{1}\,\sigma_{1}u^{\dagger}\sigma_{2}u-\tilde{\eta}_{2}\,u=0\,$,
$\Gamma\tilde{\eta}_{2x}-\frac{1}{2}\tilde{\eta}_{2}+\mathrm{i}\,\Gamma\eta_{1}\,\sigma_{1}u^{\dagger}_{x}\sigma_{2}=0\,,\qquad\tilde{\eta}_{2t}-\frac{1}{2}\Gamma\tilde{\eta}_{2}+\mathrm{i}\,\tilde{\eta}_{2}\,u\,\sigma_{1}u^{\dagger}\sigma_{2}-\mathrm{i}\,\Gamma\eta_{1}\,\sigma_{1}u^{\dagger}\sigma_{2}=0\,.$ (4.1)
Furthermore, let $\Omega$ be an $n\times n$ matrix solution of the Lyapunov equation
$\Gamma\Omega+\Omega\Gamma^{\dagger}=\mathrm{i}\,\eta_{1}\sigma_{1}\eta_{1}^{\dagger}\Gamma^{\dagger}+\tilde{\eta}_{2}\,\sigma_{2}\,\tilde{\eta}_{2}^{\dagger}\,.$ (4.2)
Then, in any open subset of $\mathcal{U}$ where $\Omega$ is invertible,
$u^{\prime}=u-\sigma_{2}\,\tilde{\eta}_{2}^{\dagger}\,\Omega^{-1}\eta_{1}$ (4.3)
solves the $m_{2}\times m_{1}$ matrix FL equation (1.2).
Proof.
In Theorem 3.2, we impose the conditions
$\Delta=-\Gamma^{\dagger}\,,\quad\theta_{1}=\mathrm{i}\,\sigma_{1}(\eta_{1}^{\dagger}\,\Gamma^{\dagger}-u^{\dagger}\,\tilde{\eta}_{2}^{\dagger})\,,\quad\theta_{2}=\sigma_{2}\,\tilde{\eta}_{2}^{\dagger}\,$, (4.4)
where now
$\tilde{\eta}_{2}=\eta_{2}-\mathrm{i}\,\eta_{1}\sigma_{1}u^{\dagger}\sigma_{2}\,.$
Then, by using (1.4), the four equations (3.6) turn out to be consequences of (3.7), and the latter system can be equivalently replaced by (4.1). The Sylvester equation (3.8) becomes the Lyapunov equation (4.2). We have $\Omega=\Omega_{1}+\Omega_{2}$, where
$\Gamma\Omega_{1}+\Omega_{1}\Gamma^{\dagger}=\mathrm{i}\,\eta_{1}\sigma_{1}\eta_{1}^{\dagger}\Gamma^{\dagger}\,,\qquad\Gamma\Omega_{2}+\Omega_{2}\Gamma^{\dagger}=\tilde{\eta}_{2}\,\sigma_{2}\,\tilde{\eta}_{2}^{\dagger}\,.$
Since we assume the spectrum condition for $\Gamma$, the Lyapunov equation has a unique solution. It follows that $\Omega_{2}^{\dagger}=\Omega_{2}$, and thus
$\tilde{\eta}_{2}\,\sigma_{2}\,\tilde{\eta}_{2}^{\dagger}=\Gamma\Omega+\Omega^{\dagger}\Gamma^{\dagger}-(\Gamma\Omega_{1}+\Omega_{1}^{\dagger}\Gamma^{\dagger})\,.$
The last bracket vanishes since the Lyapunov equation for $\Omega_{1}$ implies $(\Gamma\Omega_{1})^{\dagger}=-\Gamma\Omega_{1}$. Hence
$\tilde{\eta}_{2}\,\sigma_{2}\,\tilde{\eta}_{2}^{\dagger}=\Gamma\Omega+\Omega^{\dagger}\Gamma^{\dagger}\,.$ (4.5)
$u^{\prime}$ in (3.10) takes the form (4.3), and $v^{\prime}$ can be written as
$v^{\prime}=G\,H^{-1}\,$,
with
$G=\sigma_{1}u^{\dagger}\sigma_{2}+\sigma_{1}\big(\Gamma\eta_{1}-\tilde{\eta}_{2}\,u\big)^{\dagger}\,(\Gamma\Omega)^{-1}\tilde{\eta}_{2}\,,\qquad H=I_{m_{2}}-\sigma_{2}\,\tilde{\eta}_{2}^{\dagger}\,(\Gamma\Omega)^{-1}\tilde{\eta}_{2}\,.$
Now
$\sigma_{1}u^{\prime\dagger}\sigma_{2}\,H$ $=$ $\big(\sigma_{1}u^{\dagger}\sigma_{2}-\sigma_{1}\,\eta_{1}^{\dagger}\,\Omega^{\dagger-1}\tilde{\eta}_{2}\big)\big(I_{m_{2}}-\sigma_{2}\,\tilde{\eta}_{2}^{\dagger}\,(\Gamma\Omega)^{-1}\tilde{\eta}_{2}\big)$
$=$ $\sigma_{1}u^{\dagger}\sigma_{2}-\sigma_{1}u^{\dagger}\,\tilde{\eta}_{2}^{\dagger}\,(\Gamma\Omega)^{-1}\tilde{\eta}_{2}-\sigma_{1}\eta_{1}^{\dagger}\,\Omega^{\dagger-1}\tilde{\eta}_{2}+\sigma_{1}\eta_{1}^{\dagger}\,\Omega^{\dagger-1}\tilde{\eta}_{2}\sigma_{2}\tilde{\eta}_{2}^{\dagger}\,(\Gamma\Omega)^{-1}\tilde{\eta}_{2}=G\,$,
where we used (4.5) in the last step, shows that $v^{\prime}=\sigma_{1}u^{\prime\dagger}\sigma_{2}$. Since, according to Theorem 3.2, $(u^{\prime},v^{\prime})$ solves (1.3), it follows that $u^{\prime}$ solves the matrix Fokas-Lenells equation (1.2). ∎
Corollary 4.2.
The spectrum condition for $\Gamma$ can be dropped in Theorem 4.1 if the assumptions there are supplemented by
$\Omega_{x}=\mathrm{i}\,\eta_{1x}\,\sigma_{1}\,\eta_{1}^{\dagger}+(\tilde{\eta}_{2x}\,\sigma_{2}+\mathrm{i}\,\eta_{1}\sigma_{1}\,u^{\dagger}_{x})\,\tilde{\eta}_{2}^{\dagger}\,\Gamma^{\dagger-1}\,$,
$\Omega_{t}=\frac{1}{2}\Big(-\mathrm{i}\,\eta_{1}\sigma_{1}\eta_{1}^{\dagger}\Gamma^{\dagger}+2\mathrm{i}\,\eta_{1}\sigma_{1}u^{\dagger}\tilde{\eta}_{2}^{\dagger}+\tilde{\eta}_{2}\sigma_{2}\tilde{\eta}_{2}^{\dagger}\Big)\,$, (4.6)
and if in addition (4.5) is imposed.
Proof.
(4.6) results from (3.9), using (4.4). According to Theorem 3.2, taking a seed that solves the matrix FL equation, the resulting $u^{\prime}$ and $v^{\prime}$ then solve (1.3). Now (4.5) guarantees that $u^{\prime}$ actually solves the matrix FL equation.
As a consequence of (4.6) and the linear system (4.1), it follows that
$\big(\Gamma\Omega+\Omega^{\dagger}\Gamma^{\dagger}-\tilde{\eta}_{2}\,\sigma_{2}\tilde{\eta}_{2}^{\dagger}\big)_{x}=0\,,\qquad\big(\Gamma\Omega+\Omega^{\dagger}\Gamma^{\dagger}-\tilde{\eta}_{2}\,\sigma_{2}\tilde{\eta}_{2}^{\dagger}\big)_{t}=0\,$,
so that (4.5) and (4.6) are indeed compatible. ∎
Remark 4.3.
If the spectrum condition for $\Gamma$ holds, then (4.5) and (4.6) are automatically satisfied as a consequence of the Lyapunov equation (4.2). $\Box$
Remark 4.4.
A crucial feature of the vectorial Darboux transformation is the following “superposition property”. Let $(\Gamma^{(i)},\eta^{(i)}_{1},\tilde{\eta}^{(i)}_{2},\Omega^{(i)})$, $i=1,2$, be data determining solutions ${u^{\prime}}^{(i)}$ of the matrix FL equation, for the same seed solution $u$. Let
$\Gamma=\left(\begin{array}[]{cc}\Gamma^{(1)}&0\\ 0&\Gamma^{(2)}\end{array}\right)\,,\quad\eta_{1}=\left(\begin{array}[]{c}\eta^{(1)}_{1}\\ \eta^{(2)}_{1}\end{array}\right)\,,\quad\tilde{\eta}_{2}=\left(\begin{array}[]{c}\tilde{\eta}^{(1)}_{2}\\ \tilde{\eta}^{(2)}_{2}\end{array}\right)\,,\quad\Omega=\left(\begin{array}[]{cc}\Omega^{(1)}&\Omega^{(12)}\\ \Omega^{(21)}&\Omega^{(2)}\end{array}\right)\,.$
The corresponding linear system (4.1) is then solved. Essentially, it only remains to determine the matrices $\Omega^{(12)}$ and $\Omega^{(21)}$ in order to get a new solution $u^{\prime}$, which is then a kind of nonlinear superposition of ${u^{\prime}}^{(1)}$ and ${u^{\prime}}^{(2)}$. The Lyapunov equation (4.2) reduces to the two Sylvester equations
$\Gamma^{(1)}\Omega^{(12)}+\Omega^{(12)}\Gamma^{(2)\dagger}=\mathrm{i}\,\eta^{(1)}_{1}\sigma_{1}\,\eta^{(2)\dagger}_{1}\Gamma^{(2)\dagger}+\tilde{\eta}^{(1)}_{2}\sigma_{2}\,\tilde{\eta}^{(2)\dagger}_{2}\,$,
$\Gamma^{(2)}\Omega^{(21)}+\Omega^{(21)}\Gamma^{(1)\dagger}=\mathrm{i}\,\eta^{(2)}_{1}\sigma_{1}\,\eta^{(1)\dagger}_{1}\Gamma^{(1)\dagger}+\tilde{\eta}^{(2)}_{2}\sigma_{2}\,\tilde{\eta}^{(1)\dagger}_{2}\,$.
If $\Gamma^{(1)}$ and $\Gamma^{(2)\dagger}$ have no common eigenvalue, these equations possess unique solutions $\Omega^{(12)}$ and $\Omega^{(21)}$. Otherwise additional equations, obtained from (4.5) and (4.6), have to be solved. $\Box$
4.1 Zero seed solutions
For $u=0$ (zero seed), (4.1) is completely solved by
$\eta_{1}=e^{-\frac{1}{2}(\Gamma^{-1}x+\Gamma\,t)}\,a_{1}\,,\qquad\tilde{\eta}_{2}=e^{\frac{1}{2}(\Gamma^{-1}x+\Gamma\,t)}\,a_{2}\,$,
with constant $n\times m_{1}$, respectively $n\times m_{2}$, matrices $a_{1}$ and $a_{2}$. The Lyapunov equation reads
$\Gamma\,\Omega+\Omega\,\Gamma^{\dagger}=\mathrm{i}\,e^{-\frac{1}{2}(\Gamma^{-1}\,x+\Gamma\,t)}\,\boldsymbol{a}\,\Gamma^{\dagger}\,e^{-\frac{1}{2}({\Gamma^{\dagger}}^{-1}\,x+\Gamma^{\dagger}\,t)}+e^{\frac{1}{2}(\Gamma^{-1}\,x+\Gamma\,t)}\,\boldsymbol{b}\,e^{\frac{1}{2}({\Gamma^{\dagger}}^{-1}\,x+\Gamma^{\dagger}\,t)}\,$,
where we introduced the Hermitian $n\times n$ matrices
$\boldsymbol{a}=a_{1}\,\sigma_{1}\,a_{1}^{\dagger}\,,\qquad\boldsymbol{b}=a_{2}\,\sigma_{2}\,a_{2}^{\dagger}\,.$
If the spectrum condition for $\Gamma$ holds, the Lyapunov equation possesses a unique solution $\Omega$. If not, then the Lyapunov equation leads to constraints and we also have to use the equations in Corollary 4.2 to determine $\Omega$. According to Theorem 4.1 and Corollary 4.2,
$u^{\prime}=-\sigma_{2}\,\tilde{\eta}_{2}^{\dagger}\,\Omega^{-1}\eta_{1}=-\sigma_{2}\,a_{2}^{\dagger}\,e^{\frac{1}{2}({\Gamma^{\dagger}}^{-1}x+\Gamma^{\dagger}\,t)}\,\Omega^{-1}e^{-\frac{1}{2}(\Gamma^{-1}x+\Gamma\,t)}\,a_{1}$ (4.8)
solves the matrix FL equation (1.2). In the following, we set
$\varphi(\Gamma)=\frac{1}{2}(\Gamma^{-1}\,x+\Gamma\,t)\,$,
and work out some classes of solutions.
Example 4.5.
Ordinary matrix multi-soliton solutions are obtained if $\Gamma=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n})$, with $\gamma_{i}\neq-\gamma_{j}^{\ast}$ for $i,j=1,\ldots,n$. Then $\Omega$ is given by the Cauchy-like matrix with components
$\Omega_{ij}=\frac{1}{\gamma_{i}+\gamma_{j}^{\ast}}\,\Big(\mathrm{i}\,\boldsymbol{a}_{ij}\,\gamma_{j}^{\ast}\,e^{-\varphi(\gamma_{i})-\varphi(\gamma_{j}^{\ast})}+\boldsymbol{b}_{ij}\,e^{\varphi(\gamma_{i})+\varphi(\gamma_{j}^{\ast})}\Big)\,.$
Now (4.8) yields an $n$-soliton solution of the $m_{2}\times m_{1}$ matrix FL equation.
For $n=1$, excluding $\boldsymbol{a}=\boldsymbol{b}=0$, we get the matrix single soliton solution
$u^{\prime}=-\frac{2\,\mathrm{Re}(\gamma)\,e^{-2\,\mathrm{i}\,\mathrm{Im}(\varphi)}}{\big(\mathrm{i}\,\mathrm{Re}(\gamma)+\mathrm{Im}(\gamma)\big)\,\boldsymbol{a}\,e^{-2\mathrm{Re}(\varphi)}+\boldsymbol{b}\,e^{2\mathrm{Re}(\varphi)}}\,\sigma_{2}\,a_{2}^{\dagger}\,a_{1}\,$,
where $a_{2}^{\dagger}\,a_{1}$ is a Kronecker product. Since $\mathrm{Re}(\gamma)\neq 0$ and $\boldsymbol{a},\boldsymbol{b}\in\mathbb{R}$, this is regular and vanishes as $x\to\pm\infty$ or $t\to\pm\infty$, except on the lines where $x+|\gamma|^{2}t$ is constant. $\Box$
Example 4.6.
Let $\Gamma=-\mathrm{i}\,\mathrm{diag}(k_{1},\ldots,k_{n})$, $k_{i}\in\mathbb{R}\setminus\{0\}$ and $k_{i}\neq k_{j}$, $i,j=1,\ldots,n$. Then the Lyapunov equation becomes the constraint
$\boldsymbol{b}_{ii}=k_{i}\,\boldsymbol{a}_{ii}\,,\qquad i=1,\ldots,n\,.$
The differential equation for $\Omega_{ii}$ reads
$\Omega_{iix}=\frac{\boldsymbol{a}_{ii}}{k_{i}}\,,\qquad\Omega_{iit}=k_{i}\,\boldsymbol{a}_{ii}\,$,
so that
$\Omega_{ii}=\boldsymbol{a}_{ii}(k_{i}^{-1}x+k_{i}\,t)+c_{i}\,.$
with a constant $c_{i}$, for which (4.5) requires
$\mathrm{Im}(c_{i})=\frac{1}{2}\boldsymbol{a}_{ii}\,.$
For $i\neq j$, we obtain
$\Omega_{ij}=\frac{1}{\mathrm{i}\,(k_{i}-k_{j})}\,\Big(k_{j}\,\boldsymbol{a}_{ij}\,e^{-\frac{\mathrm{i}}{2}(k_{i}^{-1}-k_{j}^{-1})(x+k_{i}k_{j}\,t)}-\boldsymbol{b}_{ij}\,e^{\frac{\mathrm{i}}{2}(k_{i}^{-1}-k_{j}^{-1})(x+k_{i}k_{j}\,t)}\Big)\,.$
Inserting our results in (4.8), we obtain another $n$-soliton solution of the $m_{2}\times m_{1}$ matrix FL equation. For $n=1$, this is the quasi-rational matrix single soliton
$u^{\prime}=-\frac{e^{-\mathrm{i}\,\big(k^{-1}x-k\,t\big)}}{\boldsymbol{a}\,(k^{-1}x+k\,t)+\mathrm{Re}(c_{1})+\frac{1}{2}\mathrm{i}\,\boldsymbol{a}}\,\sigma_{2}\,a_{2}^{\dagger}\,a_{1}\,$,
which is regular. The absolute values of the components fall off toward zero away from the line $x=-k^{2}t-\mathrm{Re}(c_{1})\,k/\boldsymbol{a}$. In this sense, the solution is concentrated along this line. $\Box$
Example 4.7.
Let $\Gamma=-\mathrm{i}\,k\,I_{n}$, where $k\in\mathbb{R}\setminus\{0\}$. Then the Lyapunov equation becomes the constraint $\boldsymbol{b}=k\,\boldsymbol{a}$, and (4.6) leads to $\Omega=(k^{-1}x+k\,t)\,\boldsymbol{a}+\boldsymbol{c}$, with a constant $n\times n$ matrix $\boldsymbol{c}$, for which (4.5) requires $\boldsymbol{a}+\mathrm{i}\,(\boldsymbol{c}-\boldsymbol{c}^{\dagger})=0$. The resulting solution is $u^{\prime}=-e^{-\mathrm{i}\,(k^{-1}x-k\,t)}\,\sigma_{2}\,a_{2}^{\dagger}\,\Omega^{-1}a_{1}$. $\Box$
Example 4.8.
Let $n=2$ and $\Gamma=\mathrm{diag}(\gamma,-\gamma^{\ast})$ with $\mathrm{Re}(\gamma)\neq 0$. Then we have
$\eta_{1}=\left(\begin{array}[]{cc}e^{-\varphi(\gamma)}&0\\ 0&e^{\varphi(\gamma^{\ast})}\end{array}\right)\,a_{1}\,,\qquad\tilde{\eta}_{2}=\left(\begin{array}[]{cc}e^{\varphi(\gamma)}&0\\ 0&e^{-\varphi(\gamma^{\ast})}\end{array}\right)\,a_{2}\,.$
Hermiticity of the matrices $\boldsymbol{a}$ and $\boldsymbol{b}$ means that the components $\boldsymbol{a}_{11},\boldsymbol{a}_{22},\boldsymbol{b}_{11},\boldsymbol{b}_{22}$ are real and $\boldsymbol{a}_{21}=\boldsymbol{a}_{12}^{\ast},\boldsymbol{b}_{21}=\boldsymbol{b}_{12}^{\ast}$. The off-diagonal parts of the Lyapunov equation amount to the constraint
$\boldsymbol{b}_{12}=\mathrm{i}\,\gamma\,\boldsymbol{a}_{12}\,$,
and the diagonal parts determine the diagonal components of $\Omega$,
$\Omega_{11}$ $=$ $\frac{1}{2\mathrm{Re}(\gamma)}\big(\mathrm{i}\,\gamma^{\ast}\boldsymbol{a}_{11}\,e^{-2\mathrm{Re}(\varphi(\gamma))}+\boldsymbol{b}_{11}\,e^{2\mathrm{Re}(\varphi(\gamma))}\big)\,$,
$\Omega_{22}$ $=$ $\frac{1}{2\mathrm{Re}(\gamma)}\big(\mathrm{i}\,\gamma\,\boldsymbol{a}_{22}\,e^{2\mathrm{Re}(\varphi(\gamma))}-\boldsymbol{b}_{22}\,e^{-2\mathrm{Re}(\varphi(\gamma))}\big)\,.$
The differential equations for $\Omega$ are now equivalent to
$\Omega_{12x}=-\frac{\mathrm{i}\,\boldsymbol{a}_{12}}{\gamma}\,,\quad\Omega_{21x}=\frac{\mathrm{i}\,\boldsymbol{a}_{12}^{\ast}}{\gamma^{\ast}}\,,\quad\Omega_{12t}=\mathrm{i}\,\gamma\,\boldsymbol{a}_{12}\,,\quad\Omega_{21t}=-\mathrm{i}\,\gamma^{\ast}\boldsymbol{a}_{12}^{\ast}\,$,
which integrates to
$\Omega_{12}=-\mathrm{i}\,\boldsymbol{a}_{12}\,(\gamma^{-1}x-\gamma t)+c_{12}\,,\qquad\Omega_{21}=\mathrm{i}\,\boldsymbol{a}_{12}^{\ast}\,(\gamma^{\ast-1}x-\gamma^{\ast}\,t)+c_{21}\,$,
with complex constants $c_{12}$ and $c_{21}$. The remaining condition (4.5) becomes
$c_{21}=c_{12}^{\ast}+\mathrm{i}\,\boldsymbol{a}_{12}^{\ast}\,.$
Inserting the results in (4.8) yields the following solution of the $m_{2}\times m_{1}$ matrix FL equation,
$u^{\prime}=-\frac{1}{\mathrm{det}(\Omega)}\,\sigma_{2}\,a_{2}^{\dagger}\left(\begin{array}[]{cc}e^{-2\mathrm{i}\,\mathrm{Im}(\varphi(\gamma))}\,\Omega_{22}&-e^{2\varphi(\gamma^{\ast})}\,\Omega_{12}\\ -e^{-2\varphi(\gamma)}\,\Omega_{21}&e^{-2\mathrm{i}\mathrm{Im}(\varphi(\gamma))}\,\Omega_{11}\end{array}\right)a_{1}\,$,
where
$\mathrm{det}(\Omega)$ $=$ $\frac{1}{4\mathrm{Re}(\gamma)^{2}}\Big(\mathrm{i}\,\gamma\,\boldsymbol{a}_{22}\boldsymbol{b}_{11}\,e^{4\mathrm{Re}(\varphi(\gamma))}-\mathrm{i}\,\gamma^{\ast}\boldsymbol{a}_{11}\boldsymbol{b}_{22}\,e^{-4\mathrm{Re}(\varphi(\gamma))}-|\gamma|^{2}\boldsymbol{a}_{11}\boldsymbol{a}_{22}-\boldsymbol{b}_{11}\boldsymbol{b}_{22}\Big)$
$-\big|\mathrm{i}\,\boldsymbol{a}_{12}(\gamma^{-1}x-\gamma t)-c_{12}\big|^{2}-|\boldsymbol{a}_{12}|^{2}\rho-\mathrm{i}\,\boldsymbol{a}_{12}^{\ast}c_{12}\,.$
Besides the exponential dependence on the independent variables via $e^{\varphi(\gamma)}$, there is also a rational dependence on $\gamma^{-1}x-\gamma t$ and its complex conjugate. Fig. 1 shows plots of the absolute values for two solutions of the scalar FL equation ($m_{1}=m_{2}=1$) from this class.
Figure 1: Plots of the absolute value of solutions of the scalar FL equation (with $\sigma_{1}=\sigma_{2}=1$) obtained with an anti-conjugate pair of spectral parameters from the family in Example 4.8. We set $\gamma=1+\mathrm{i}$, $a_{1}=(1,1)^{T}$, $a_{2}=(\mathrm{i}\,\gamma,1)^{T}$ and $c_{12}=1$ (left plot), respectively $c_{12}=50$ (right plot).
$\Box$
Example 4.9.
Let $n=2$ and $\Gamma$ be the $2\times 2$ Jordan block
$\Gamma=\left(\begin{array}[]{cc}\gamma&0\\ 1&\gamma\end{array}\right)\,$,
with eigenvalue $\gamma$. Then we have
$\eta_{1}=e^{-\varphi}\,\left(\begin{array}[]{cc}1&0\\ \rho&1\end{array}\right)\,a_{1}\,,\qquad\tilde{\eta}_{2}=e^{\varphi}\,\left(\begin{array}[]{cc}1&0\\ -\rho&1\end{array}\right)\,a_{2}\,$,
where $\varphi$ stands for $\varphi(\gamma)$, and
$\rho=\frac{1}{2}\big(\gamma^{-2}x-t\big)\,$,
which introduces the same dependence on the independent variables that we met in the preceding example. After having determined
$\Omega=\left(\begin{array}[]{cc}\Omega_{11}&\Omega_{12}\\ \Omega_{21}&\Omega_{22}\end{array}\right)\,,\qquad\Omega^{-1}=\frac{1}{\mathrm{det}(\Omega)}\left(\begin{array}[]{cc}\Omega_{22}&-\Omega_{12}\\ -\Omega_{21}&\Omega_{11}\end{array}\right)\,$,
from (4.8) we obtain the solution
$u^{\prime}=-\frac{e^{-2\,\mathrm{i}\,\mathrm{Im}(\varphi)}}{\det(\Omega)}\,\sigma_{2}\,a_{2}^{\dagger}\left(\begin{array}[]{cc}\Omega_{22}-\rho\,\Omega_{12}+\rho^{\ast}\,\Omega_{21}-|\rho|^{2}\,\Omega_{11}&-\Omega_{12}-\rho^{\ast}\Omega_{11}\\ -\Omega_{21}+\rho\,\Omega_{11}&\Omega_{11}\end{array}\right)\,a_{1}\,.$ (4.16)
If $\kappa:=2\mathrm{Re}(\gamma)\neq 0$, the solution of the Lyapunov equation is found from Example A.2 in Appendix Appendix A: Some facts about the Lyapunov equation to be
$\Omega=\mathrm{i}\,\tilde{\Omega}_{1}\Gamma^{\dagger}+\Omega_{2}$
with
$\tilde{\Omega}_{1}$ $=$ $\frac{1}{\kappa}\left(\begin{array}[]{cc}A_{11}&A_{12}+\rho_{-}^{\ast}\,A_{11}\\ A_{12}^{\ast}+\rho_{-}\,A_{11}&(|\rho_{-}|^{2}+\kappa^{-2})\,A_{11}+\rho_{-}\,A_{12}+\rho_{-}^{\ast}A_{12}^{\ast}+A_{22}\end{array}\right)\,$,
$\Omega_{2}$ $=$ $\frac{1}{\kappa}\left(\begin{array}[]{cc}B_{11}&B_{12}-\rho_{+}^{\ast}\,B_{11}\\ B_{12}^{\ast}-\rho_{+}\,B_{11}&(|\rho_{+}|^{2}+\kappa^{-2})\,B_{11}-\rho_{+}\,B_{12}-\rho_{+}^{\ast}B_{12}^{\ast}+B_{22}\end{array}\right)\,$,
where
$A_{ij}=\boldsymbol{a}_{ij}\,e^{-2\,\mathrm{Re}(\varphi)}\,,\qquad B_{ij}=\boldsymbol{b}_{ij}\,e^{2\,\mathrm{Re}(\varphi)}\,,\qquad\rho_{\pm}=\rho\pm\kappa^{-1}\,.$
Hence, $\Omega$ has the components
$\Omega_{11}$ $=$ $\kappa^{-1}(\mathrm{i}\,\gamma^{\ast}\,A_{11}+B_{11})\,$,
$\Omega_{12}$ $=$ $\kappa^{-1}\Big(\mathrm{i}\,\big(A_{11}+\gamma^{\ast}(A_{12}+\rho_{-}^{\ast}A_{11})\big)+B_{12}-\rho_{+}^{\ast}B_{11}\Big)\,$,
$\Omega_{21}$ $=$ $\kappa^{-1}\big(\mathrm{i}\,\gamma^{\ast}(A_{12}^{\ast}+\rho_{-}\,A_{11})+B_{12}^{\ast}-\rho_{+}\,B_{11}\big)\,$,
$\Omega_{22}$ $=$ $\kappa^{-1}\Big(\mathrm{i}\,(A_{12}^{\ast}+\rho_{-}\,A_{11})+\mathrm{i}\,\gamma^{\ast}\big((|\rho_{-}|^{2}+\kappa^{-2})\,A_{11}+\rho_{-}\,A_{12}+\rho_{-}^{\ast}A_{12}^{\ast}+A_{22}\big)$
$+(|\rho_{+}|^{2}+\kappa^{-2})\,B_{11}-\rho_{+}\,B_{12}-\rho_{+}^{\ast}B_{12}^{\ast}+B_{22}\Big)\,.$
Besides the exponential dependence via $A_{ij}$ and $B_{ij}$, the solution (4.16) also depends on the independent variables $x$ and $t$ through the linear expression $\rho$. Generically, the determinant of $\Omega$ has non-zero real and imaginary parts, so that $u^{\prime}$ is regular. We obtain matrix generalizations of “double pole solutions”, see [4] for the scalar case.
If $\gamma$ is imaginary, i.e., $\gamma=-\mathrm{i}\,k$ with $k\in\mathbb{R}\setminus\{0\}$, then we have $\rho=-\frac{1}{2}(x/k^{2}+t)=\rho^{\ast}$ and the Lyapunov equation requires (cf. Example A.4 in Appendix Appendix A: Some facts about the Lyapunov equation)
$k\,\boldsymbol{a}_{11}-\boldsymbol{b}_{11}=0\,$, (4.19)
$\Omega_{11}=-\rho\,(k\,\boldsymbol{a}_{11}+\boldsymbol{b}_{11})+\mathrm{i}\,\boldsymbol{a}_{11}-k\,\boldsymbol{a}_{12}+\boldsymbol{b}_{12}=-\rho\,(k\,\boldsymbol{a}_{11}+\boldsymbol{b}_{11})-k\,\boldsymbol{a}_{12}^{\ast}+\boldsymbol{b}_{12}^{\ast}\,$,
$\Omega_{12}+\Omega_{21}=-\rho^{2}(k\,\boldsymbol{a}_{11}-\boldsymbol{b}_{11})+\rho\,(\mathrm{i}\,\boldsymbol{a}_{11}-k\,\boldsymbol{a}_{12}-\boldsymbol{b}_{12})-\rho\,(k\,\boldsymbol{a}_{12}^{\ast}+\boldsymbol{b}_{12}^{\ast})$
$\hskip 62.59596pt+\mathrm{i}\,\boldsymbol{a}_{12}^{\ast}-k\,\boldsymbol{a}_{22}+\boldsymbol{b}_{22}\,.$
Using the first equation to eliminate $\boldsymbol{b}_{11}$ from the others, leads to
$\Omega_{11}=-2\,k\,\rho\,\boldsymbol{a}_{11}+\mathrm{i}\,\boldsymbol{a}_{11}-k\,\boldsymbol{a}_{12}+\boldsymbol{b}_{12}=-2\,k\,\rho\,\boldsymbol{a}_{11}-k\,\boldsymbol{a}_{12}^{\ast}+\boldsymbol{b}_{12}^{\ast}\,$,
$\Omega_{12}+\Omega_{21}=\rho\,(\mathrm{i}\,\boldsymbol{a}_{11}-k\,\boldsymbol{a}_{12}-\boldsymbol{b}_{12})-\rho\,(k\,\boldsymbol{a}_{12}^{\ast}+\boldsymbol{b}_{12}^{\ast})+\mathrm{i}\,\boldsymbol{a}_{12}^{\ast}-k\,\boldsymbol{a}_{22}+\boldsymbol{b}_{22}\,.$
Consistency of the two expressions for $\Omega_{11}$ requires
$\boldsymbol{a}_{11}=2\,\mathrm{Im}(k\,\boldsymbol{a}_{12}-\boldsymbol{b}_{12})\,$,
which is
$\mathrm{Im}(a_{21}\,a_{22}^{\ast})-\sigma_{1}\sigma_{2}\,\Big(\frac{1}{2}|a_{11}|^{2}-k\,\mathrm{Im}(a_{11}\,a_{12}^{\ast})\Big)=0\,.$ (4.20)
The remaining components of $\Omega$ are determined by
$\Omega_{12x}$ $=$ $\frac{1}{2}k^{-2}\big(\mathrm{i}\,\boldsymbol{a}_{11}+k\,\boldsymbol{a}_{12}+\boldsymbol{b}_{12}\big)\,,\qquad\Omega_{12t}=-\frac{1}{2}\big(\mathrm{i}\,\boldsymbol{a}_{11}-k\,\boldsymbol{a}_{12}-\boldsymbol{b}_{12}\big)\,$,
$\Omega_{21x}$ $=$ $-\frac{1}{2}k^{-2}\big(2\mathrm{i}\,\boldsymbol{a}_{11}-k\,\boldsymbol{a}_{12}^{\ast}-\boldsymbol{b}_{12}^{\ast}\big)\,,\qquad\Omega_{21t}=\frac{1}{2}(k\,\boldsymbol{a}_{12}^{\ast}+\boldsymbol{b}_{12}^{\ast})\,$,
$\Omega_{22x}$ $=$ $\frac{1}{2}k^{-3}\Big(2k^{2}\rho^{2}\,\boldsymbol{a}_{11}-k\,\rho\,\big(-k\,(\boldsymbol{a}_{12}+\boldsymbol{a}_{12}^{\ast})+\mathrm{i}\,\boldsymbol{a}_{11}+\boldsymbol{b}_{12}+\boldsymbol{b}_{12}^{\ast}\big)+k^{2}\boldsymbol{a}_{22}-k\,(\mathrm{i}\,\boldsymbol{a}_{12}-\boldsymbol{b}_{22})$
$+\boldsymbol{a}_{11}-\mathrm{i}\,(\boldsymbol{b}_{12}-\boldsymbol{b}_{12}^{\ast})\Big)\,$,
$\Omega_{22t}$ $=$ $\frac{1}{2}k^{-1}\Big(2k^{2}\rho^{2}\,\boldsymbol{a}_{11}-k\,\rho\,\big(-k\,(\boldsymbol{a}_{12}+\boldsymbol{a}_{12}^{\ast})+\mathrm{i}\,\boldsymbol{a}_{11}+\boldsymbol{b}_{12}+\boldsymbol{b}_{12}^{\ast}\big)+k^{2}\boldsymbol{a}_{22}-k\,(\mathrm{i}\,\boldsymbol{a}_{12}^{\ast}-\boldsymbol{b}_{22})\Big)\,.$
This system is solved by
$\Omega_{12}$ $=$ $-\rho\,(k\,\boldsymbol{a}_{12}+\boldsymbol{b}_{12}+\mathrm{i}\,\boldsymbol{a}_{11})-\mathrm{i}\,\boldsymbol{a}_{11}\,t+c_{12}\,$,
$\Omega_{21}$ $=$ $-\rho\,(k\,\boldsymbol{a}_{12}^{\ast}+\boldsymbol{b}_{12}^{\ast}-2\mathrm{i}\,\boldsymbol{a}_{11})+\mathrm{i}\,\boldsymbol{a}_{11}\,t+c_{21}\,$,
$\Omega_{22}$ $=$ $-\frac{2}{3}k\,\rho^{3}\boldsymbol{a}_{11}+\frac{1}{2}\rho^{2}\big(-k\,(\boldsymbol{a}_{12}+\boldsymbol{a}_{12}^{\ast})+\mathrm{i}\,\boldsymbol{a}_{11}+\boldsymbol{b}_{12}+\boldsymbol{b}_{12}^{\ast}\big)-\rho\,(k\,\boldsymbol{a}_{22}+\boldsymbol{b}_{22}-2\mathrm{i}\,\boldsymbol{a}_{12}+\mathrm{i}\,\boldsymbol{a}_{12}^{\ast})$
$+\mathrm{i}\,(\boldsymbol{a}_{12}-\boldsymbol{a}_{12}^{\ast})\,t+c_{22}\,$,
with constants $c_{ij}$, for which (4.5) and the above equation for $\Omega_{12}+\Omega_{21}$ require
$\mathrm{Re}(c_{12})=\frac{1}{2}(-k\,\boldsymbol{a}_{22}+\boldsymbol{b}_{22})\,,\qquad\mathrm{Re}(c_{21})=\frac{1}{2}\big(-k\,\boldsymbol{a}_{22}+2\,\mathrm{Im}(\boldsymbol{a}_{12})+\boldsymbol{b}_{22}\big)\,$,
$\mathrm{Im}(c_{21})=\mathrm{Re}(\boldsymbol{a}_{12})-\mathrm{Im}(c_{12})\,,\qquad\mathrm{Im}(c_{22})=\frac{1}{2}\boldsymbol{a}_{22}\,$,
so that $\mathrm{Re}(c_{22})$ and $\mathrm{Im}(c_{12})$ remain as free parameters. Let $k$ be determined by (4.19). The constants $a_{ij}$, $i,j=1,2$, are still constrained by (4.20). Then (4.16) is a quasi-rational solution of the $m_{2}\times m_{1}$ matrix FL equation. If $\boldsymbol{a}_{11}\neq 0$, for constant $t$ (respectively $x$) the numerators of components of $u^{\prime}$ grow at most with $x^{3}$ (respectively $t^{3}$), whereas $\det(\Omega)$ grows with the fourth power. On a line, along which $\rho$ and thus $x+k^{2}t$ is constant, $|u^{\prime}|\sim 1/|t|$ as $t\to\pm\infty$. As a consequence, the solution must be localized along a worldline, which is not a straight line. In particular, our results supplement the treatment of the scalar FL equation ($m_{1}=m_{2}=1$) in Section 4.1 of [4], also see Fig. 2 for two examples.
Figure 2: Plots of the absolute value of solutions of the scalar FL equation (with $\sigma_{1}=\sigma_{2}=1$) obtained with a $2\times 2$ Jordan block $\Gamma$ with eigenvalue $\gamma=-\mathrm{i}$, and $a_{1}=(1,-\frac{1}{2}\mathrm{i})^{T}$, $a_{2}=(1,0)^{T}$, $\mathrm{Im}(c_{12})=0$, and $\mathrm{Re}(c_{22})=0$, respectively =50=50. Also see Fig. 3 for details on the first example.
Figure 3: Plots of the absolute value of the first solution of the scalar FL equation displayed in Fig. 2, at t=2,1,0,1,3t=-2,-1,0,1,3. Below is a plot of $\mathrm{Re}(u^{\prime})$ and $\mathrm{Im}(u^{\prime})$ at $t=0$. The apparent cusp of $\mathrm{Im}(u^{\prime})$ at $t=0$ is just a sharp maximum.
$\Box$
The preceding examples exhaust the possibilities of solutions up to $n=2$. Mostly, they also present corresponding extensions to arbitrary $n$, but in the case, where $\Gamma$ is a Jordan block, we restricted our computations to $n=2$. The step to higher $n$ is straightforward, but the resulting expressions lengthy. According to Remark 4.4, we can superpose any number of solutions from the classes elaborated in the above examples.
5 The vector FL equation
Choosing $m_{1}=1$, $u=(u_{1},\ldots,u_{m_{2}})^{T}$ is an $m_{2}$-component column vector. Without restriction of generality, we may set $\sigma_{1}=1$. Renaming $\sigma_{2}$ to $\boldsymbol{\sigma}$, (1.2) reads
$u_{xt}=(1+\mathrm{i}\,u^{\dagger}\boldsymbol{\sigma}\,u_{x})\,u+\mathrm{i}\,(u^{\dagger}\,\boldsymbol{\sigma}\,u)\,u_{x}\,$, (5.1)
which first appeared in [5], also see [6, 7, 8, 9, 17, 11, 14, 16].
Setting all components but one to zero, (5.1) reduces to the scalar FL equation for the remaining component. If all components of $u$ are equal, then $u_{\mu}/\sqrt{m_{2}}$ satisfies the scalar FL equation.
Now we specialize Theorem 4.1 to the vector FL case. In this case, the linear system (4.1) can be decoupled and determines $\eta_{1}$ in terms of $\tilde{\eta}_{2}$.
Theorem 5.1.
Let $\Gamma$ be an invertible constant $n\times n$ matrix, such that $\Gamma$ and $\Gamma^{\dagger}$ have no common eigenvalue (spectrum condition). Let $u$ be a solution of the vector FL equation (5.1) on an open set $\mathcal{U}\subset\mathbb{R}^{2}$, where $u^{\dagger}\boldsymbol{\sigma}u\neq 0$ and $u_{x}^{\dagger}\boldsymbol{\sigma}u\neq 0$. Furthermore, let $\tilde{\eta}$ be an $n\times m_{2}$ matrix solution of the linear system
$\hskip-22.76228pt\Big(\tilde{\eta}_{t}-\frac{1}{2}\Gamma\tilde{\eta}+\frac{u^{\dagger}\boldsymbol{\sigma}u}{u_{x}^{\dagger}\boldsymbol{\sigma}u}\big(\Gamma\tilde{\eta}_{x}-\frac{1}{2}\tilde{\eta}\big)+\mathrm{i}\,u^{\dagger}\boldsymbol{\sigma}u\,\tilde{\eta}\Big)\,u=0\,$, (5.2)
$\hskip-22.76228pt\tilde{\eta}_{xx}\,u+\tilde{\eta}_{x}\Big(u_{x}-\frac{(u_{x}^{\dagger}\boldsymbol{\sigma}u)_{x}}{u_{x}^{\dagger}\boldsymbol{\sigma}u}\,u\Big)+\frac{1}{2}\Big(\frac{(u_{x}^{\dagger}\boldsymbol{\sigma}u)_{x}}{u_{x}^{\dagger}\boldsymbol{\sigma}u}I_{n}-\frac{1}{2}\Gamma^{-1}\Big)\Gamma^{-1}\tilde{\eta}\,u+(\mathrm{i}\,u_{x}^{\dagger}\boldsymbol{\sigma}u-\frac{1}{2})\,\Gamma^{-1}\tilde{\eta}\,u_{x}=0\,$$\qquad$ (5.3)
and
$(\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\tilde{\eta})\,P=0\,,\qquad(\tilde{\eta}_{t}-\frac{1}{2}\Gamma\,\tilde{\eta})\,Q=0\,$, (5.4)
where
$P=I_{m_{2}}-\frac{u\,u^{\dagger}_{x}\boldsymbol{\sigma}}{u_{x}^{\dagger}\boldsymbol{\sigma}u}\,,\qquad Q=I_{m_{2}}-\frac{u\,u^{\dagger}\boldsymbol{\sigma}}{u^{\dagger}\boldsymbol{\sigma}u}\,$,
are projections (both with kernel spanned by $u$). Furthermore, let $\Omega$ be an $n\times n$ matrix solution of the Lyapunov equation
$\Gamma\Omega+\Omega\Gamma^{\dagger}=\frac{\mathrm{i}}{|u_{x}^{\dagger}\boldsymbol{\sigma}u|^{2}}\big(\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\tilde{\eta}\big)\,u\,\big[\big(\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\tilde{\eta}\big)u\big]^{\dagger}\,\Gamma^{\dagger}+\tilde{\eta}\,\boldsymbol{\sigma}\,\tilde{\eta}^{\dagger}\,$. (5.5)
Then, in any open subset of $\mathcal{U}$ where $\Omega$ is invertible,
$u^{\prime}=u-\frac{\mathrm{i}}{u_{x}^{\dagger}\boldsymbol{\sigma}u}\,\boldsymbol{\sigma}\,\tilde{\eta}^{\dagger}\,\Omega^{-1}\big(\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\tilde{\eta}\big)\,u$ (5.6)
solves the vector FL equation (5.1).
Proof.
Writing $\tilde{\eta}$ instead of $\tilde{\eta}_{2}$, the linear system (4.1) reads
$\Gamma\,\eta_{1x}+\frac{1}{2}\eta_{1}-\tilde{\eta}\,u_{x}=0\,,\qquad\quad\eta_{1t}+\big(\frac{1}{2}\Gamma-\mathrm{i}\,u^{\dagger}\boldsymbol{\sigma}u\,I_{n}\big)\,\eta_{1}-\tilde{\eta}\,u=0\,$,
$\Gamma\tilde{\eta}_{x}-\frac{1}{2}\tilde{\eta}+\mathrm{i}\,\Gamma\eta_{1}\,u^{\dagger}_{x}\boldsymbol{\sigma}=0\,,\qquad\tilde{\eta}_{t}-\frac{1}{2}\Gamma\tilde{\eta}+\mathrm{i}\,\tilde{\eta}\,uu^{\dagger}\boldsymbol{\sigma}-\mathrm{i}\,\Gamma\eta_{1}\,u^{\dagger}\boldsymbol{\sigma}=0\,.$
Multiplying the last two equations by $u$ from the right, and solving them for $\eta_{1}$, we get
$\frac{\mathrm{i}}{u_{x}^{\dagger}\boldsymbol{\sigma}u}\big(\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\tilde{\eta}\big)\,u=\eta_{1}=-\frac{\mathrm{i}}{u^{\dagger}\boldsymbol{\sigma}u}\big(\Gamma^{-1}\tilde{\eta}_{t}-\frac{1}{2}\tilde{\eta}\big)u+\Gamma^{-1}\tilde{\eta}\,u\,.$
Inserting this back in the last two equations of the linear system, they are turned into (5.4). Using the expressions for $\eta_{1}$ in the first two equations of the linear system, yields (5.3) and
$\Big(\big(\tilde{\eta}_{t}-\frac{1}{2}\Gamma\,\tilde{\eta}\big)\,u\Big)_{t}+\Big(\frac{1}{2}\Gamma-\big(\mathrm{i}\,u^{\dagger}\boldsymbol{\sigma}u+\frac{(u^{\dagger}\boldsymbol{\sigma}u)_{t}}{u^{\dagger}\boldsymbol{\sigma}u}\big)I_{n}\Big)\big(\tilde{\eta}_{t}-\frac{1}{2}\Gamma\tilde{\eta}\big)\,u+\mathrm{i}\,u^{\dagger}\boldsymbol{\sigma}u\,(\tilde{\eta}u)_{t}$
$-u^{\dagger}\boldsymbol{\sigma}u\,\big(\frac{1}{2}\mathrm{i}\,\Gamma-u^{\dagger}\boldsymbol{\sigma}u\,I_{n}\big)\,\tilde{\eta}\,u=0\,.$
The two expressions for $\eta_{1}$ imply (5.2), and the last equation can be shown to be a consequence of (5.2) and (5.3). (5.5) and (5.6) result from (4.2) and (4.3), respectively, by using the first expression for $\eta_{1}$ above. ∎
Corollary 5.2.
The spectrum condition for $\Gamma$ can be dropped in Theorem 5.1 if the assumptions there are supplemented by
$\Gamma\Omega+\Omega^{\dagger}\Gamma^{\dagger}=\tilde{\eta}\,\boldsymbol{\sigma}\,\tilde{\eta}^{\dagger}\,$, (5.7)
and
$\Omega_{x}=\frac{\mathrm{i}}{u^{\dagger}\boldsymbol{\sigma}u_{x}}\Big(\frac{1}{u_{x}^{\dagger}\boldsymbol{\sigma}u}\Xi\Big)_{x}\,\Xi^{\dagger}+\Big(\tilde{\eta}_{x}\,\boldsymbol{\sigma}-\frac{1}{u_{x}^{\dagger}\boldsymbol{\sigma}u}\Xi\,u_{x}^{\dagger}\Big)\,\tilde{\eta}^{\dagger}\,\Gamma^{\dagger-1}\,$,
$\Omega_{t}=-\frac{\mathrm{i}}{2|u_{x}^{\dagger}\boldsymbol{\sigma}u|^{2}}\Xi\,\Xi^{\dagger}\,\Gamma^{\dagger}+\big(\frac{1}{2}\tilde{\eta}\,\boldsymbol{\sigma}-\frac{1}{u_{x}^{\dagger}\boldsymbol{\sigma}u}\Xi\,u^{\dagger}\big)\,\tilde{\eta}^{\dagger}\,$, (5.8)
where we introduced
$\Xi=(\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\tilde{\eta})\,u\,.$ (5.9)
Proof.
This follows directly from Corollary 4.2, using intermediate results in the proof of Theorem 4.1. ∎
Remark 5.3.
Using a familiar formula for the determinant of a $2\times 2$ block matrix, the new solution (5.6) can be expressed as follows in terms of a quotient of determinants,
$u^{\prime}_{\mu}$ $=$ $u_{\mu}-\frac{\mathrm{i}}{(u_{x}^{\dagger}\boldsymbol{\sigma}u)\,\det(\Omega)}\,\det\left(\begin{array}[]{cc}\Omega&\big(\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\tilde{\eta}\big)\,u\\ \boldsymbol{\sigma}\,\tilde{\eta}_{\mu}^{\dagger}&0\end{array}\right)$
$=$ $\frac{1}{\det(\Omega)}\,\det\left(\begin{array}[]{cc}\Omega&\big(\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\tilde{\eta}\big)\,u\\ \mathrm{i}\,(u_{x}^{\dagger}\boldsymbol{\sigma}u)^{-1}\boldsymbol{\sigma}\,\tilde{\eta}_{\mu}^{\dagger}&u_{\mu}\end{array}\right)\qquad\quad\mu=1,\ldots,m_{2}\,.$
In the cited references, where solutions have been generated via the classical (iterative) Darboux transformation method, they are mostly presented in such a form, with a specialized matrix $\Omega$. $\Box$
Remark 5.4.
For $m_{2}=1$, the equations (5.4) are identically satisfied. For $m_{2}>1$ and if $\boldsymbol{\sigma}$ is positive (or negative) definite, at each point of $\mathbb{R}^{2}$, where $u$ is defined and $u\neq 0$, the vector $u$ determines a splitting of the $m_{2}$-dimensional complex vector space, $\mathbb{C}^{m_{2}}=\mathrm{span}(u)\oplus V$, where $V$ is the orthogonal complement of the linear span of $u$ with respect to the inner product $\langle w|w^{\prime}\rangle=w^{\dagger}\boldsymbol{\sigma}w^{\prime}$. Then the second of (5.4) implies that $(\tilde{\eta}_{t}-\frac{1}{2}\Gamma\tilde{\eta})\,w=0$ for all $w\in V$. Whereas $Q$ is an orthogonal projection, $P$ is not. We have
$P(v)=-\frac{\langle u_{x}|v\rangle}{\langle u_{x}|u\rangle}\,u+v\qquad\forall v\perp u\,.$
The first equation of (5.4) says that $\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\tilde{\eta}=0$ holds on the orthogonal complement of $\mathrm{span}(u_{x})$. In particular, we have $\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\tilde{\eta}=0$ and $\tilde{\eta}_{t}-\frac{1}{2}\Gamma\tilde{\eta}=0$ on the orthogonal complement of $\mathrm{span}(u,u_{x})$. $\Box$
Proposition 5.5.
Let $\boldsymbol{U}_{0}$ be any constant $\boldsymbol{\sigma}$-unitary matrix (i.e., a constant $m_{2}\times m_{2}$ matrix satisfying $\boldsymbol{U}_{0}^{\dagger}\,\boldsymbol{\sigma}\,\boldsymbol{U}_{0}=\boldsymbol{\sigma}$). Then $u\mapsto\boldsymbol{U}_{0}\,u$ is a symmetry of the vector FL equation (5.1) and also a symmetry of the Darboux transformation in Theorem 5.1 and Corollary 5.2.
Proof.
It is easily verified that, if $u$ is a solution of (5.1), then also $\boldsymbol{U}_{0}\,u$. Furthermore, we note that
$u\mapsto\boldsymbol{U}_{0}\,u\,,\qquad\tilde{\eta}\mapsto\tilde{\eta}\,\boldsymbol{U}_{0}^{-1}\,$,
is a symmetry of the linear system (5.2) - (5.4), since $P\mapsto\boldsymbol{U}_{0}\,P\,\boldsymbol{U}_{0}^{-1}$ and $Q\mapsto\boldsymbol{U}_{0}\,Q\,\boldsymbol{U}_{0}^{-1}$. $\Omega$ is invariant, hence it follows that $u^{\prime}\mapsto\boldsymbol{U}_{0}\,u^{\prime}$. We used that $\boldsymbol{\sigma}$ is Hermitian and involutary. ∎
Remark 5.6.
We specialize the important “superposition property”, formulated in Remark 4.4, to the case of the vector FL equation. Let $(\Gamma^{(i)},\tilde{\eta}^{(i)},\Omega^{(i)})$, $i=1,2$, be data determining solutions ${u^{\prime}}^{(i)}$ of the vector FL equation via the vectorial Darboux transformation, using the same seed solution $u$. Let
$\Gamma=\left(\begin{array}[]{cc}\Gamma^{(1)}&0\\ 0&\Gamma^{(2)}\end{array}\right)\,,\quad\tilde{\eta}=\left(\begin{array}[]{c}\tilde{\eta}^{(1)}\\ \tilde{\eta}^{(2)}\end{array}\right)\,,\quad\Omega=\left(\begin{array}[]{cc}\Omega^{(1)}&\Omega^{(12)}\\ \Omega^{(21)}&\Omega^{(2)}\end{array}\right)\,.$
The linear system (5.2)-(5.4) is then solved, and it only remains to determine $\Omega^{(12)}$ and $\Omega^{(21)}$ in order to get a new solution $u^{\prime}$, which then represents a nonlinear superposition of ${u^{\prime}}^{(1)}$ and ${u^{\prime}}^{(2)}$. These non-diagonal blocks of $\Omega$ are subject to the Sylvester equations
$\Gamma^{(1)}\Omega^{(12)}+\Omega^{(12)}\Gamma^{(2)\dagger}=\frac{\mathrm{i}}{|u_{x}^{\dagger}\boldsymbol{\sigma}u|^{2}}\Xi^{(1)}\,\Xi^{(2)\dagger}\,\Gamma^{(2)\dagger}+\tilde{\eta}^{(1)}\,\boldsymbol{\sigma}\,\tilde{\eta}^{(2)\dagger}\,$,
$\Gamma^{(2)}\Omega^{(21)}+\Omega^{(21)}\Gamma^{(1)\dagger}=\frac{\mathrm{i}}{|u_{x}^{\dagger}\boldsymbol{\sigma}u|^{2}}\Xi^{(2)}\,\Xi^{(1)\dagger}\,\Gamma^{(1)\dagger}+\tilde{\eta}^{(2)}\,\boldsymbol{\sigma}\,\tilde{\eta}^{(1)\dagger}\,$,
where
$\Xi^{(i)}=\Big(\tilde{\eta}_{x}^{(i)}-\frac{1}{2}\Gamma^{(i)-1}\tilde{\eta}^{(i)}\Big)\,u\,,\qquad i=1,2\,.$
If $\Gamma^{(1)}$ and $\Gamma^{(2)\dagger}$ have no common eigenvalue, these equations possess unique solutions. Otherwise additional equations, obtained from those in Corollary 5.2, have to be solved.
In the simplest case of a superposition of two $n=1$ solutions, setting $\Gamma^{(1)}=\gamma_{1}$ and $\Gamma^{(2)}=\gamma_{2}$, and assuming $\gamma_{1}\neq-\gamma_{2}^{\ast}$, we have
$\Omega^{(12)}=\frac{1}{\gamma_{1}+\gamma_{2}^{\ast}}\Big(\frac{\mathrm{i}\,\gamma_{2}^{\ast}}{|u_{x}^{\dagger}\boldsymbol{\sigma}u|^{2}}\,\Xi^{(1)}\,\Xi^{(2)\ast}+\tilde{\eta}^{(1)}\,\boldsymbol{\sigma}\,\tilde{\eta}^{(2)\dagger}\Big)\,$,
$\Omega^{(21)}=\frac{1}{\gamma_{2}+\gamma_{1}^{\ast}}\Big(\frac{\mathrm{i}\,\gamma_{1}^{\ast}}{|u_{x}^{\dagger}\boldsymbol{\sigma}u|^{2}}\,\Xi^{(2)}\,\Xi^{(1)\ast}+\tilde{\eta}^{(2)}\,\boldsymbol{\sigma}\,\tilde{\eta}^{(1)\dagger}\Big)\,$,
where the $\Xi^{(i)}$ are scalars. $\Box$
In the following subsection and the next section we concentrate on a plane wave seed solution in the above Darboux transformation for the vector FL equation. Inserting the plane wave ansatz
$u=e^{\mathrm{i}\,(\boldsymbol{\alpha}x+\boldsymbol{\beta}t)}\,A$
in (5.1), with a constant column vector $A$ and commuting constant Hermitian matrices $\boldsymbol{\alpha},\boldsymbol{\beta}$, leads to
$\boldsymbol{\alpha}\,\boldsymbol{\beta}=(u^{\dagger}\boldsymbol{\sigma}\boldsymbol{\alpha}\,u-1)\,I_{m_{2}}+(u^{\dagger}\boldsymbol{\sigma}u)\,\boldsymbol{\alpha}\,.$
Assuming that $\boldsymbol{\alpha}$ is invertible, this yields
$\boldsymbol{\beta}=(u^{\dagger}\boldsymbol{\sigma}\boldsymbol{\alpha}\,u-1)\,\boldsymbol{\alpha}^{-1}+(u^{\dagger}\boldsymbol{\sigma}u)\,I_{m_{2}}\,$,
which requires that the scalars $u^{\dagger}\boldsymbol{\sigma}u$ and $u^{\dagger}\boldsymbol{\sigma}\boldsymbol{\alpha}\,u$ are constant. This is guaranteed is $\boldsymbol{\alpha}$ commutes with $\boldsymbol{\sigma}$. The assumptions $u^{\dagger}\boldsymbol{\sigma}u\neq 0$ and $u_{x}^{\dagger}\boldsymbol{\sigma}u\neq 0$ in Theorem 5.1 now read
$A^{\dagger}\boldsymbol{\sigma}A\neq 0\,,\qquad A^{\dagger}\boldsymbol{\alpha}\,\boldsymbol{\sigma}A\neq 0\,.$
Throughout, they will be assumed in the following.
5.1 Plane wave seed
We restrict our considerations to the case where $\boldsymbol{\sigma}=I_{m_{2}}$ and $\boldsymbol{\alpha}=\mathrm{diag}(\alpha_{1},\ldots,\alpha_{m_{2}})$, with $\alpha_{\mu}\in\mathbb{R}$. The plane wave solution of the vector FL equation is then given by
$u_{\mu}=A_{\mu}\,e^{\mathrm{i}\,\varphi_{\mu}}\,,\qquad\varphi_{\mu}=\alpha_{\mu}x+\beta_{\mu}t\,,\qquad\mu=1,\ldots,m_{2}\,$,
with $A_{\mu}\in\mathbb{C}$ and $\alpha_{\mu}\in\mathbb{R}$. The absolute value of $\alpha_{\mu}$ is the wave number of the respective mode. Assuming $\alpha_{\mu}\neq 0$ for all $\mu$, we have $\boldsymbol{\beta}=\mathrm{diag}(\beta_{1},\ldots,\beta_{m_{2}})$ with
$\beta_{\mu}=\sum_{\nu=1}^{m_{2}}\Big(1+\frac{\alpha_{\nu}}{\alpha_{\mu}}\Big)\,|A_{\nu}|^{2}-\frac{1}{\alpha_{\mu}}\qquad\quad\mu=1,\ldots,m_{2}\,.$
This solution will be taken as the seed to generate further exact solutions via Theorem 5.1. We have
$u_{x}=\mathrm{i}\,\boldsymbol{\alpha}\,u\,,\qquad u_{t}=\mathrm{i}\,\boldsymbol{\beta}\,u\,.$
Let $\boldsymbol{U}$ be a unitary $m_{2}\times m_{2}$ matrix such that
$u=e^{\mathrm{i}\,(\boldsymbol{\alpha}x+\boldsymbol{\beta}t)}\,A=\boldsymbol{U}\left(\begin{array}[]{c}\|A\|\\ 0\\ \vdots\\ 0\end{array}\right)\,,\qquad A=\left(\begin{array}[]{c}A_{1}\\ \vdots\\ A_{m_{2}}\end{array}\right)\,$,
where $\|A\|$ is the standard norm of the complex vector $A$. This requires $\boldsymbol{U}_{\mu 1}=e^{\mathrm{i}\,\varphi_{\mu}}A_{\mu}/\|A\|$, $\mu=1,\ldots,m_{2}$. We set
$\tilde{\eta}=\chi\,\boldsymbol{U}^{\dagger}\,,\qquad\chi=(\chi_{1},\ldots,\chi_{m_{2}})\,$,
with $n$-component column vectors $\chi_{\mu}$, $\mu=1,\ldots,m_{2}$. Since we assume $A^{\dagger}\boldsymbol{\alpha}A\neq 0$, (5.2) and the second of equations (5.4) become
$\chi_{1t}-\frac{1}{2}\Gamma\,\chi_{1}+\mathrm{i}\,\|A\|^{2}\chi_{1}-\mathrm{i}\,\chi\boldsymbol{U}^{\dagger}\boldsymbol{\beta}\,\boldsymbol{U}\,(1,0,\ldots,0)^{T}$
$\qquad+\mathrm{i}\,\frac{\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\big(\Gamma(\chi_{1x}-\mathrm{i}\,\chi\boldsymbol{U}^{\dagger}\boldsymbol{\alpha}\,\boldsymbol{U}\,(1,0,\ldots,0)^{T})-\frac{1}{2}\chi_{1}\big)=0\,$,
$\chi_{\mu t}=\frac{1}{2}\Gamma\,\chi_{\mu}-\sum_{\nu=1}^{m_{2}}\chi_{\nu}\,(\boldsymbol{U}_{t}^{\dagger}\boldsymbol{U})_{\nu\mu}\qquad\mu=2,\ldots,m_{2}\,.$ (5.14)
Here we used $\boldsymbol{U}_{x}^{\dagger}u=-\boldsymbol{U}^{\dagger}u_{x}=-\mathrm{i}\,\boldsymbol{U}^{\dagger}\boldsymbol{\alpha}\,u$ and $\boldsymbol{U}_{t}^{\dagger}u=-\boldsymbol{U}^{\dagger}u_{t}=-\mathrm{i}\,\boldsymbol{U}^{\dagger}\boldsymbol{\beta}\,u$. The first of (5.4) tells us that we have $\tilde{\eta}_{x}-\frac{1}{2}\Gamma^{-1}\,\chi=0$ on the orthogonal complement of $u_{x}$. A vector (field) $w$ is orthogonal to $u_{x}$ if and only if $\tilde{w}=\boldsymbol{U}^{\dagger}w$ is orthogonal to $\boldsymbol{U}^{\dagger}u_{x}=\mathrm{i}\,\boldsymbol{U}^{\dagger}\boldsymbol{\alpha}\,u=\mathrm{i}\,\boldsymbol{U}^{\dagger}\boldsymbol{\alpha}\,\boldsymbol{U}(\|A\|,0,\ldots,0)^{T}$, i.e.,
$0=(1,0,\ldots,0)\,\boldsymbol{U}^{\dagger}\boldsymbol{\alpha}\,\boldsymbol{U}\,\tilde{w}=\sum_{\nu=1}^{m_{2}}(\boldsymbol{U}^{\dagger}\boldsymbol{\alpha}\,\boldsymbol{U})_{1\nu}\,\tilde{w}_{\nu}\,.$
Then (5.4) reads
$(\chi_{x}+\chi\,\boldsymbol{U}_{x}^{\dagger}\boldsymbol{U}-\frac{1}{2}\Gamma^{-1}\,\chi)\,\tilde{w}=0\,$, (5.15)
for all such $\tilde{w}$. The remaining equation (5.3) of the linear system becomes
$\chi_{1xx}-\frac{1}{4}\Gamma^{-2}\chi_{1}+\Big(2\chi_{x}\boldsymbol{U}_{x}^{\dagger}\boldsymbol{U}+\chi\,\boldsymbol{U}_{xx}^{\dagger}\boldsymbol{U}+\mathrm{i}\,(\chi_{x}+\chi\,\boldsymbol{U}_{x}^{\dagger}\boldsymbol{U})\,\boldsymbol{U}^{\dagger}\boldsymbol{\alpha}\,\boldsymbol{U}$
$+\mathrm{i}\,(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2})\,\Gamma^{-1}\chi\,\boldsymbol{U}^{\dagger}\boldsymbol{\alpha}\,\boldsymbol{U}\Big)\,(1,0,\ldots,0)^{T}=0\,.$ (5.16)
The Lyapunov equation (5.5) reads
$\Gamma\,\Omega+\Omega\,\Gamma^{\dagger}=\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\,\Xi\,\Xi^{\dagger}\Gamma^{\dagger}+\sum_{\mu=1}^{m_{2}}\chi_{\mu}\,\chi_{\mu}^{\dagger}\,$, (5.17)
now with
$\Xi=\|A\|\,\big(\chi_{1x}-\frac{1}{2}\Gamma^{-1}\chi_{1}-\mathrm{i}\,\chi\,\boldsymbol{U}^{\dagger}\boldsymbol{\alpha}\,\boldsymbol{U}\,(1,0,\ldots,0)^{T}\big)\,.$
If $\Gamma$ satisfies the spectrum condition, then (5.17) has a unique solution $\Omega$. If $\Gamma$ does not satisfy the spectrum condition, then the additional equations in Corollary 5.2 have to be taken into account in order to determine $\Omega$. (5.6) takes the form
$u^{\prime}$ $=$ $u+\frac{1}{A^{\dagger}\boldsymbol{\alpha}A}\,\boldsymbol{U}\,\chi^{\dagger}\,\Omega^{-1}\Xi$
$=$ $\boldsymbol{U}\left[\left(\begin{array}[]{c}\|A\|\\ 0\\ \vdots\\ 0\end{array}\right)+\frac{1}{A^{\dagger}\boldsymbol{\alpha}A}\,\left(\begin{array}[]{c}\chi_{1}^{\dagger}\,\Omega^{-1}\,\Xi\\ \vdots\\ \chi_{m_{2}}^{\dagger}\,\Omega^{-1}\,\Xi\end{array}\right)\right]\,.$
In the following section, we will concentrate on the two-component vector FL equation.
6 Solutions of a two-component vector FL equation
In this section, we elaborate the set of solutions of the two-component vector FL equation, i.e, (5.1) with $m_{2}=2$, generated via the Darboux transformation obtained in the preceding section. In order to somewhat reduce the complexity, we restrict our considerations to the case where $\boldsymbol{\sigma}=I_{2}$ and $\boldsymbol{\alpha}=\mathrm{diag}(\alpha_{1},\alpha_{2})$, with $\alpha_{\mu}\in\mathbb{R}$, $\mu=1,2$. Then we can choose
$\boldsymbol{U}=\left(\begin{array}[]{cc}e^{\mathrm{i}\,\varphi_{1}}&0\\ 0&e^{\mathrm{i}\,\varphi_{2}}\end{array}\right)\,\boldsymbol{U}_{0}\,,\qquad\boldsymbol{U}_{0}=\frac{1}{\|A\|}\left(\begin{array}[]{cc}A_{1}&-A_{2}^{\ast}\\ A_{2}&A_{1}^{\ast}\end{array}\right)\,$, (6.5)
so that $\boldsymbol{U}_{x}=\mathrm{i}\,\boldsymbol{\alpha}\,\boldsymbol{U},\boldsymbol{U}_{t}=\mathrm{i}\,\boldsymbol{\beta}\,\boldsymbol{U}$ and
$\boldsymbol{U}^{\dagger}\boldsymbol{\alpha}\,\boldsymbol{U}=\boldsymbol{U}_{0}^{\dagger}\,\boldsymbol{\alpha}\,\boldsymbol{U}_{0}=\frac{1}{\|A\|^{2}}\left(\begin{array}[]{cc}\alpha_{1}|A_{1}|^{2}+\alpha_{2}|A_{2}|^{2}&(\alpha_{2}-\alpha_{1})\,A_{1}^{\ast}A_{2}^{\ast}\\ (\alpha_{2}-\alpha_{1})\,A_{1}A_{2}&\alpha_{2}|A_{1}|^{2}+\alpha_{1}|A_{2}|^{2}\end{array}\right)\,.$
From (5.15) we obtain
$\chi_{2x}-\Big(\frac{1}{2}\Gamma^{-1}+\mathrm{i}\,\frac{\alpha_{1}\alpha_{2}\,\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\,I_{n}\Big)\,\chi_{2}=\frac{(\alpha_{2}-\alpha_{1})A_{1}^{\ast}A_{2}^{\ast}}{A^{\dagger}\boldsymbol{\alpha}A}\big(\chi_{1x}-\frac{1}{2}\Gamma^{-1}\chi_{1}\big)\,$, (6.7)
and (5.16) becomes
$\chi_{1xx}-\frac{1}{4}\Gamma^{-2}\chi_{1}-\frac{\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}}\big(\chi_{1x}+(\frac{1}{2}-A^{\dagger}\boldsymbol{\alpha}A)\,\Gamma^{-1}\chi_{1}\big)$
$-\frac{\mathrm{i}\,A_{1}A_{2}}{\|A\|^{2}}(\alpha_{2}-\alpha_{1})\,\big(\chi_{2x}+(\frac{1}{2}-A^{\dagger}\boldsymbol{\alpha}A)\,\Gamma^{-1}\chi_{2}\big)=0\,.$
Eliminating $\chi_{2x}$ with the help of the preceding equation, we get
$\chi_{1xx}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}\,\chi_{1x}-\frac{1}{4}\Gamma^{-2}\chi_{1}$
$+\frac{\mathrm{i}}{2\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\Big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\Big)\,\Gamma^{-1}\chi_{1}$
$-(\alpha_{2}-\alpha_{1})\frac{A_{1}A_{2}}{\|A\|^{2}}\Big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,\Gamma^{-1}-\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}I_{n}\Big)\,\chi_{2}=0\,.$ (6.8)
We note that
$\Xi$ $=$ $\|A\|\,\big(\chi_{1x}-\frac{1}{2}\Gamma^{-1}\chi_{1}-\mathrm{i}\,\chi\,\boldsymbol{U}_{0}^{\dagger}\,\boldsymbol{\alpha}\,\boldsymbol{U}_{0}\,(1,0)^{T}\big)$ (6.9)
$=$ $\|A\|\,\Big(\chi_{1x}-\big(\frac{1}{2}\Gamma^{-1}+\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}}\,I_{n}\big)\,\chi_{1}\Big)-\mathrm{i}\,(\alpha_{2}-\alpha_{1})\,\frac{A_{1}A_{2}}{\|A\|}\,\chi_{2}\,.$
The Lyapunov equation (5.17) now reads
$\Gamma\,\Omega+\Omega\,\Gamma^{\dagger}=\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Xi\,\Xi^{\dagger}\Gamma^{\dagger}+\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger}\,.$ (6.10)
From this point on, we have to treat the cases $\alpha_{1}=\alpha_{2}$ and $\alpha_{1}\neq\alpha_{2}$ separately. This is done in Sections 6.1 and 6.2, respectively. Although the first case is non-generic, it exhibits interesting features. Moreover, it is considerably simpler than the second, generic case and provides a convenient first step into the plethora of solutions generated via Theorem 5.1.
The corresponding analysis is different, depending on whether the spectrum condition for the matrix $\Gamma$ in the Lyapunov equation holds or not. It is necessary to distinguish several cases. Moreover, for special choices of $\Gamma$, the linear system, which here is equivalent to a constant coefficient linear system, “degenerates” and admits quasi-rational solutions. In Sections 6.1 and 6.2, we will organize the various possibilities into a collection of examples.
Remark 6.1.
If $\boldsymbol{\sigma}=\mathrm{diag}(1,\sigma)$, with $\sigma\in\{\pm 1\}$, we choose
$\boldsymbol{U}=\frac{1}{\sqrt{A^{\dagger}\boldsymbol{\sigma}A}}\left(\begin{array}[]{cc}e^{\mathrm{i}\,\varphi_{1}}&0\\ 0&e^{\mathrm{i}\,\varphi_{2}}\end{array}\right)\left(\begin{array}[]{cc}A_{1}&-\sigma\,A_{2}^{\ast}\\ A_{2}&A_{1}^{\ast}\end{array}\right)\,$,
which satisfies $(A_{1}e^{\mathrm{i}\,\varphi_{1}},A_{2}e^{\mathrm{i}\,\varphi_{2}})^{T}=\boldsymbol{U}\,(\sqrt{A^{\dagger}\boldsymbol{\sigma}A},0)^{T}$ and the $\boldsymbol{\sigma}$-unitarity condition $\boldsymbol{U}^{\dagger}\boldsymbol{\sigma}\boldsymbol{U}=\boldsymbol{\sigma}$. If $\sigma=-1$, so that $\boldsymbol{\sigma}$ is indefinite, we have to restrict $A$ such that $|A_{1}|\neq|A_{2}|$. $\Box$
Remark 6.2.
The form of the two-component vector FL equation in [7, 16] and corresponding solutions are obtained from ours via $x\mapsto 2x$ and $t\mapsto-t/2$. In [10, 11] the two-component vector FL equation is taken in the form
$\mathrm{i}\,D_{\tau}\frac{\partial v_{1}}{\partial\xi}+\frac{\sigma}{2}\frac{\partial^{2}v_{1}}{\partial\tau^{2}}+(2|v_{1}|^{2}+|v_{2}|^{2})\,D_{\tau}v_{1}+v_{1}v_{2}^{\ast}D_{\tau}v_{2}=0\,$,
$\mathrm{i}\,D_{\tau}\frac{\partial v_{2}}{\partial\xi}+\frac{\sigma}{2}\frac{\partial^{2}v_{2}}{\partial\tau^{2}}+(2|v_{2}|^{2}+|v_{1}|^{2})\,D_{\tau}v_{2}+v_{2}v_{1}^{\ast}D_{\tau}v_{1}=0\,$,
where $D_{\tau}=1+\mathrm{i}\,\epsilon\,\partial/\partial\tau$, $\xi$ and $\tau$ are independent variables, and $\epsilon\neq 0$ and $\sigma$ are real constants. Then
$v_{j}(\tau,\xi)=e^{\mathrm{i}\,(\epsilon^{-1}\tau+\sigma\epsilon^{-2}\xi)}\,u_{j}(x,t)\qquad j=1,2\,$,
with $x=-\frac{1}{4}\sigma\epsilon^{-4}\,(\sigma\,\xi+\epsilon\,\tau)$ and $t=\xi$, maps (solutions of) the latter system to (solutions of) our form of the two-component vector FL equation, and vice versa (also see [16]). $\Box$
6.1 The special plane wave case $\alpha_{1}=\alpha_{2}$
If $\alpha_{1}=\alpha_{2}=:\alpha$, we have
$\beta:=\beta_{1}=\beta_{2}=2\|A\|^{2}-1/\alpha\,.$
Writing
$\chi_{1}=e^{\frac{\mathrm{i}}{2}\varphi}\,\tilde{\chi}_{1}\,,\qquad\chi_{2}=e^{\mathrm{i}\,\varphi}\,\tilde{\chi}_{2}\,$,
with
$\varphi=\alpha\,x+\beta\,t=\alpha\,x+(2\|A\|^{2}-1/\alpha)\,t\,$,
the linear system becomes
$\tilde{\chi}_{1xx}-R^{2}\,\tilde{\chi}_{1}=0\,,\quad\tilde{\chi}_{1t}+\frac{\mathrm{i}}{\alpha}\Gamma\,\tilde{\chi}_{1x}=0\,$, (6.12)
where [^6] If the right hand side does not admit a matrix root, we shall simply regard $R^{2}$ as an abbreviation for the right hand side of (6.13).
$R^{2}=\frac{1}{4}\Gamma^{-2}-\frac{\mathrm{i}}{2}\,\alpha^{2}\,\beta\,\Gamma^{-1}-\frac{\alpha^{2}}{4}I_{n}\,$, (6.13)
and
$\tilde{\chi}_{2x}-\frac{1}{2}\Gamma^{-1}\tilde{\chi}_{2}=0\,,\quad\tilde{\chi}_{2t}-\frac{1}{2}\Gamma\,\tilde{\chi}_{2}=0\,$,
which is solved by
$\tilde{\chi}_{2}=e^{\frac{1}{2}(\Gamma^{-1}x+\Gamma\,t)}\,c_{0}\,$,
where $c_{0}$ is a constant $n$-component column vector.
After determining the solution of the linear system and then $\Omega$, finally,
$u^{\prime}$ $=$ $e^{\mathrm{i}\,\varphi}\,\left(\begin{array}[]{cc}A_{1}+(\alpha\,\|A\|^{3})^{-1}(A_{1}\chi_{1}^{\dagger}-A_{2}^{\ast}\chi_{2}^{\dagger})\,\Omega^{-1}\Xi\\[4.0pt] A_{2}+(\alpha\,\|A\|^{3})^{-1}(A_{2}\chi_{1}^{\dagger}+A_{1}^{\ast}\chi_{2}^{\dagger})\,\Omega^{-1}\Xi\end{array}\right)$ (6.16)
$=$ $e^{\mathrm{i}\,\varphi}\,\boldsymbol{U}_{0}\,\left[\left(\begin{array}[]{c}\|A\|\\ 0\end{array}\right)+\frac{1}{\alpha\,\|A\|^{2}}\left(\begin{array}[]{c}\chi_{1}^{\dagger}\,\Omega^{-1}\Xi\\ \chi_{2}^{\dagger}\,\Omega^{-1}\Xi\end{array}\right)\right]\,$, (6.21)
with
$\Xi=\|A\|\Big(\chi_{1x}-\big(\frac{1}{2}\Gamma^{-1}+\mathrm{i}\,\alpha\,I_{n}\big)\,\chi_{1}\Big)\,$,
is a solution of the two-component vector FL equation. Here $\boldsymbol{U}_{0}$ is the constant unitary matrix in (6.5). According to Proposition 5.5, it yields a symmetry of the vector FL equation and of its Darboux transformation. It is therefore sufficient to elaborate (6.21) with $\boldsymbol{U}_{0}$ replaced by the identity matrix, since all other solutions are obtained by application of a constant unitary matrix. This particular solution corresponds to having a plane wave background only in the first component ($A_{1}=\|A\|$) and zero background in the second ($A_{2}=0$).
We have to distinguish the following two cases.
(1) If the spectrum condition for $\Gamma$ holds, the Lyapunov equation (6.10) possesses a unique solution. In particular, if $\Gamma=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n})$, with $\gamma_{i}\neq-\gamma_{j}^{\ast}$, $i,j=1,\ldots,n$, then $\Omega$ is given by the Cauchy-like matrix with components
$\Omega_{ij}=\frac{1}{\gamma_{i}+\gamma_{j}^{\ast}}\Big(\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\,\Xi_{i}\Xi_{j}^{\ast}\gamma_{j}^{\ast}+\chi_{1i}\chi_{1j}^{\ast}+\chi_{2i}\chi_{2j}^{\ast}\Big)\,.$
Given an $n=1$ soliton solution, generated with some value of $\gamma$, an “$n$-th order” extension of it is determined by an $n\times n$ (lower triangular) Jordan block with eigenvalue $\gamma$.
(2) If the spectrum condition does not hold, the Lyapunov equation (6.10) imposes a constraint on the solution of the linear system and determines at most part of $\Omega$. According to Corollary 5.2, we also have to solve the equation (5.7), which is now
$\Gamma\Omega+\Omega^{\dagger}\Gamma^{\dagger}=\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger}\,$, (6.22)
and the compatible differential equations
$\Omega_{x}$ $=$ $\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\Xi_{x}\Xi^{\dagger}+\frac{1}{2}\Gamma^{-1}\,\chi_{1}\chi_{1}^{\dagger}\,\Gamma^{\dagger-1}+\Big(\chi_{2x}-\mathrm{i}\,\alpha\,\chi_{2}\Big)\chi_{2}^{\dagger}\Gamma^{\dagger-1}\,$,
$\Omega_{t}$ $=$ $-\frac{\mathrm{i}}{2\alpha^{2}\|A\|^{4}}\,\Xi\,\Xi^{\dagger}\,\Gamma^{\dagger}-\frac{\mathrm{i}}{\alpha\,\|A\|}\Xi\,\chi_{1}^{\dagger}+\frac{1}{2}(\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger})\,.$ (6.23)
If the expression on the right hand side of (6.13) is invertible, then it indeed possesses a matrix root $R$. (6.12) is then solved by
$\chi_{1}=e^{\frac{\mathrm{i}}{2}\varphi}\,\sum_{\kappa}e^{R_{\kappa}(I_{n}x-\frac{\mathrm{i}}{\alpha}\Gamma\,t)}\,c_{\kappa}\,$, (6.24)
where $c_{\kappa}$ are constant $n$-component column vectors and $\{R_{\kappa}\}$ is a set of different roots of (6.13).
If the right hand side of (6.13) is not invertible, so that typically no square root exists (and $R^{2}$ in (6.12) should then be replaced by the right hand side of (6.13)), we have to step back to (6.12) to determine its solution. This is actually an important case, since it leads in particular to rogue wave solutions.
Let us recall that a solution $u^{\prime}$ is called quasi-rational if it has the form $u^{\prime}=e^{\mathrm{i}\,\varphi}\,(v_{1},v_{2})^{T}$ with (in $x$ and $t$) rational expressions $v_{1},v_{2}$. If $u_{1}$ is a solution of the scalar FL equation, then $u=U_{0}\,(u_{1},0)^{T}$, with a constant unitary matrix $U_{0}$ solves the two-component vector FL equation. In this case we say that $u$ is equivalent to the solution $u_{1}$ of the scalar FL equation.
Proposition 6.3.
Any quasi-rational solution of the two-component vector FL equation, obtained by the vectorial Darboux transformation from the special plane wave seed with $\alpha_{1}=\alpha_{2}$, is equivalent to a quasi-rational solution of the scalar FL equation.
Proof.
In order for (6.21) to be quasi-rational, it is necessary that $\chi_{2}=0$. But then the second component of $u^{\prime}$ in the last expression in (6.21) vanishes. ∎
Remark 6.4.
In [8, 6, 12, 17, 13] only the special ($\alpha_{1}=\alpha_{2}$) plane wave background case has been used as a seed solution, applying a generalized Darboux transformation, with a limit procedure that takes different spectral parameters to the same value, in order to obtain higher order solutions, in particular higher order rogue waves. In our framework, the corresponding $n$-th order solution is obtained by choosing $\Gamma$ to be an $n\times n$ Jordan block. Proposition 6.3 indicates that all quasi-rational solutions presented in the mentioned publications are equivalent to quasi-rational solutions of the scalar FL equation, which is indeed the case. This is a useful insight within the vectorial Darboux transformation framework. The publication [7] also contains solutions with the special plane wave background, but steps into the richer case of different wave numbers ($\alpha_{1}\neq\alpha_{2}$, see Section 6.2), though without treating rogue waves. $\Box$
6.1.1 $n=1$
For $n=1$, the roots of the quadratic characteristic equation (6.13) are $r_{\pm}=\pm w/(2\gamma)$ with
$w=\sqrt{1-2\mathrm{i}\,\alpha^{2}(2\|A\|^{2}-\alpha^{-1})\,\gamma-\alpha^{2}\gamma^{2}}\,.$ (6.25)
Throughout this subsection, $\chi_{2}$ is always given by
$\chi_{2}=c_{0}\,e^{\mathrm{i}\,\varphi}\,e^{\frac{1}{2}(\gamma^{-1}x+\gamma t)}\,.$
Example 6.5.
Let $\mathrm{Re}(\gamma)\neq 0$ and also $w\neq 0$. Then (6.24) yields the general solution of the linear system for $\chi_{1}$,
$\chi_{1}=e^{\frac{1}{2}\mathrm{i}\,\varphi}\,\big(c_{1}\,e^{\rho}+c_{2}\,e^{-\rho}\big)\,$,
where
$\rho=\frac{1}{2}w\,\big(\frac{x}{\gamma}-\mathrm{i}\,\frac{t}{\alpha}\big)=\frac{1}{2}\big(\mathrm{Re}(w/\gamma)\,x+\alpha^{-1}\mathrm{Im}(w)\,t\big)+\frac{1}{2}\mathrm{i}\,\big(\mathrm{Im}(w/\gamma)\,x-\alpha^{-1}\mathrm{Re}(w)\,t\big)\,.$
We obtain
$\Xi$ $=$ $-\frac{\|A\|}{2\gamma}\,e^{\frac{1}{2}\mathrm{i}\,\varphi}\,W\,,\qquad W=c_{1}\,(1+\mathrm{i}\,\alpha\gamma-w)\,e^{\rho}+c_{2}\,(1+\mathrm{i}\,\alpha\gamma+w)\,e^{-\rho}\,$,
$\Omega$ $=$ $\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}\,\gamma^{\ast}}{\alpha^{2}\|A\|^{4}}|\Xi|^{2}+|\chi_{1}|^{2}+|\chi_{2}|^{2}\Big)$
$=$ $\frac{1}{2\mathrm{Re}(\gamma)}\,\Big(|c_{0}|^{2}\,e^{\mathrm{Re}(\gamma)(x/|\gamma|^{2}+t)}+\big|c_{1}\,e^{\rho}+c_{2}\,e^{-\rho}\big|^{2}+\frac{\mathrm{i}}{4\alpha^{2}\,\|A\|^{2}\,\gamma}\,|W|^{2}\Big)\,.$
As mentioned before, it is sufficient to elaborate the particular case of (6.21) where $\boldsymbol{U}_{0}$ is replaced by the identity matrix. Inserting the above expression for $\Omega$, then yields the following solution of the two-component vector FL equation,
$u_{1}^{\prime}=\|A\|\,e^{\mathrm{i}\,\varphi}\,\Big(1-\frac{W\,(c_{1}^{\ast}\,e^{\rho^{\ast}}+c_{2}^{\ast}\,e^{-\rho^{\ast}})}{2\alpha\,\gamma\,\|A\|^{2}\,\Omega}\Big)\,,\qquad u_{2}^{\prime}=-\frac{c_{0}^{\ast}\,W}{2\alpha\gamma\,\|A\|\,\Omega}e^{\frac{1}{2}\big(\mathrm{i}\,\varphi+\gamma^{\ast-1}x+\gamma^{\ast}t\big)}\,.$
If $c_{0}=0$ and either $c_{1}\neq 0$ or $c_{2}\neq 0$, we have $u_{2}^{\prime}=0$ and $u_{1}^{\prime}$ is then a solution of the scalar FL equation. Therefore we shall now assume $c_{0}\neq 0$. If $c_{1}=c_{2}=0$, $u_{1}^{\prime}$ reduces to the plane wave background (seed) solution and $u_{2}^{\prime}=0$. Therefore, besides $c_{0}\neq 0$, at least one of the constants $c_{1},c_{2}$ should be non-zero.
If $c_{2}=0$, the above solution reduces to
$u_{1}^{\prime}$ $=$ $\|A\|\,e^{\mathrm{i}\,\varphi}\,\Big(1-\mathrm{Re}(\gamma)\,\frac{\gamma^{\ast}(1-w)+\mathrm{i}\,\alpha\,|\gamma|^{2}}{2\alpha\,|\gamma|^{2}\,\|A\|^{2}\,e^{2\delta}}\,\big[1+\mathrm{tanh}\big(\vartheta+\mathrm{i}\,\mathrm{Im}(\delta)\big)\big]\Big)\,$,
$u_{2}^{\prime}$ $=$ $\mathrm{Re}(\gamma)\,\frac{\gamma^{\ast}(1-w)+\mathrm{i}\,\alpha\,|\gamma|^{2}}{2\alpha\,|\gamma|^{2}\,\|A\|\,e^{\mathrm{Re}(\delta)}}\,e^{\mathrm{i}\,\varphi}\,e^{-\frac{1}{2}\mathrm{i}\,\psi}\,\mathrm{sech}\big(\vartheta+\mathrm{i}\,\mathrm{Im}(\delta)\big)\,$,
after removing a constant, $\mathrm{Re}(\delta)+\ln(|c_{1}/c_{0}|)$, in the argument of the hyperbolic functions via a translation in $x$ or $t$, and using the unitary symmetry of the vector FL equation to eliminate constant phase factors. Here we set
$\vartheta$ $=$ $\frac{1}{2|\gamma|^{2}}\mathrm{Re}\big(\gamma\,(w^{\ast}-1)\big)\,x+\frac{1}{2}\big(\alpha^{-1}\mathrm{Im}(w)-\mathrm{Re}(\gamma)\big)\,t\,$,
$\psi$ $=$ $\big(\alpha-\mathrm{Im}(\gamma)/|\gamma|^{2}-\mathrm{Im}(w/\gamma)\big)\,x+\big(2\|A\|^{2}+\mathrm{Im}(\gamma)+\alpha^{-1}(\mathrm{Re}(w)-1)\big)\,t\,$,
$e^{2\delta}$ $=$ $1+\mathrm{i}\,\gamma^{\ast}\frac{|1+\mathrm{i}\,\alpha\gamma-w|^{2}}{4\alpha^{2}|\gamma|^{2}\,\|A\|^{2}}=1+\big(\mathrm{Im}(\gamma)+\mathrm{i}\,\mathrm{Re}(\gamma)\big)\frac{|1+\mathrm{i}\,\alpha\gamma-w|^{2}}{4\alpha^{2}|\gamma|^{2}\,\|A\|^{2}}\,.$
The above expressions also cover the alternative case $c_{1}=0$ (and $c_{2}\neq 0$) if we allow for $w$ also the negative of the square root in (6.25). If $\mathrm{Re}\big(\gamma\,(w^{\ast}-1)\big)=0$, then both components of $u^{\prime}$ are functions of $t$ only, multiplied by a phase factor with an exponent linear in $x$. See Appendices Appendix B: Special examples from the class of solutions in Example 6.5 and Appendix C: An almost trivial class of solutions of the two-component vector FL equation for this case. If $\mathrm{Re}\big(\gamma\,(w^{\ast}-1)\big)\neq 0$, then the wave, described by $u^{\prime}$, travels with constant velocity
$v=|\gamma|^{2}\,\frac{\mathrm{Re}(\gamma)-\alpha^{-1}\mathrm{Im}(w)}{\mathrm{Re}\big(\gamma\,(w^{\ast}-1)\big)}\,$,
along the $x$-axis, while $u_{2}^{\prime}$ oscillates (i.e., “beats”) with period
$T=4\pi\,\big|2\|A\|^{2}-\alpha^{-1}+\mathrm{Im}(\gamma)+\alpha^{-1}\mathrm{Re}(w)\big|^{-1}\,$,
disregarding the phase factor $e^{\mathrm{i}\,\varphi}$, which causes a modulation and is common to both components. In the comoving coordinate $x^{\prime}=x-v\,t$, we have
$\psi$ $=$ $\Big(\big(\alpha-\mathrm{Im}(\gamma)/|\gamma|^{2}-\mathrm{Im}(w/\gamma)\big)\,v+2\|A\|^{2}+\mathrm{Im}(\gamma)+\alpha^{-1}(\mathrm{Re}(w)-1)\Big)\,t$
$+\big(\alpha-\mathrm{Im}(\gamma)/|\gamma|^{2}-\mathrm{Im}(w/\gamma)\big)\,x^{\prime}\,$,
so that $u_{2}^{\prime}$ exhibits, at constant $x^{\prime}$, the period
$T^{\prime}=4\pi\,\Big|\big(\alpha-\mathrm{Im}(\gamma)/|\gamma|^{2}-\mathrm{Im}(w/\gamma)\big)\,v+2\|A\|^{2}+\mathrm{Im}(\gamma)+\alpha^{-1}(\mathrm{Re}(w)-1)\Big|^{-1}\,.$
The absolute squares of the components of $u^{\prime}$ are found to be given by
$|u_{1}^{\prime}|^{2}=\|A\|^{2}\,\Big|1+\frac{2\,\mathrm{Re}(\gamma)\,\mathrm{Re}(\gamma\,(w^{\ast}-1))}{\alpha\,|\gamma|^{2}\,\|A\|^{2}\,|e^{\vartheta+2\delta}+e^{-\vartheta}|^{2}}\Big|\,,\qquad|u_{2}^{\prime}|^{2}=\mathrm{Re}(\gamma)^{2}\,\frac{|1+\mathrm{i}\,\alpha\gamma-w|^{2}}{\alpha^{2}\,|\gamma|^{2}\,\|A\|^{2}\,|e^{\vartheta+2\delta}+e^{-\vartheta}|^{2}}\,.$
As a consequence, $u_{1}^{\prime}$ represents a dark soliton if $\alpha\,\mathrm{Re}(\gamma)\,\mathrm{Re}(\gamma\,(w^{\ast}-1))<0$. If the maximum of the absolute value in the expression for $|u_{1}^{\prime}|^{2}$ is smaller than 1, we have a single dark soliton. An example is shown in Fig. 4.
Figure 4: Plots of the absolute values of the first, and the real part (the imaginary part is similar) of the second component for an $n=1$ solution from the class in Example 6.5. The parameters are $\alpha=\|A\|=2$, $\gamma=1+\mathrm{i}$, $c_{0}=c_{1}=1$, $c_{2}=0$, but we replaced $w$ by w-w. The second plot shows the “beating”. The velocity is $v\approx 0.3$, the beating period $T^{\prime}\approx 1$.
If the maximum is greater than 1, there will be two dark soliton parts on both sides of an “anti-dark” one, i.e., a bright soliton on the plane wave background.
When $\alpha\,\mathrm{Re}(\gamma)\,\mathrm{Re}(\gamma\,(w^{\ast}-1))>0$, then $u_{1}^{\prime}$ represents an anti-dark soliton, since its amplitude is then greater than the background density value and $|u_{1}^{\prime}|$ falls off toward it as $x\to\pm\infty$.
In each case, $u_{2}^{\prime}$ describes a bright soliton (on vanishing background). Solutions of the kind described here already appeared in [7] (see Section 5.2 therein). $\Box$
Example 6.6.
Applying $\boldsymbol{U}_{0}$, given in (6.5), with $A_{2}=A_{1}\in\mathbb{R}$, to the solution with $c_{2}=0$ in the preceding example, the resulting new solution,
$u^{\prime}_{\mathrm{new}}=\frac{1}{2}\left(\begin{array}[]{c}u_{1}^{\prime}-u_{2}^{\prime}\\ u_{1}^{\prime}+u_{2}^{\prime}\end{array}\right)\,$, (6.28)
is generated from the uniform plane wave seed $u=A_{1}e^{\mathrm{i}\,\varphi}(1,1)^{T}$ by the Darboux transformation. Now we have $\|A\|^{2}=2A_{1}^{2}$. The structure of the solution displayed in (6.28), and with a relative oscillation of $u_{2}^{\prime}$, is characteristic for what has been called “beating solitons” in the NLS literature (cf. [32] and references cited there). We have
$u^{\prime}_{\mathrm{new}}(x^{\prime},t+T^{\prime}/2)=\frac{1}{2}\left(\begin{array}[]{c}u_{1}^{\prime}(x^{\prime},t)+u_{2}^{\prime}(x^{\prime},t)\\ u_{1}^{\prime}(x^{\prime},t)-u_{2}^{\prime}(x^{\prime},t)\end{array}\right)\,$,
which means an exchange of the two components, and $u^{\prime}_{\mathrm{new}}(x^{\prime},t+T^{\prime})=u^{\prime}_{\mathrm{new}}(x^{\prime},t)$.
An extreme example from the resulting class of solutions is presented in Fig. 5. Here the parameters are chosen in such a way that the coefficient of $t$ in the expression for $\psi$ vanishes, i.e., $4A_{1}^{2}-\alpha^{-1}+\mathrm{Im}(\gamma)+\alpha^{-1}\mathrm{Re}(w)=0$, so that $T$ is infinite while $T^{\prime}$ is finite.
Figure 5: Plots of the absolute value of the components of an $n=1$ beating soliton solution from the class in Example 6.6 at t=19.2t=-19.2 (left), $t=0$ (middle) and $t=16.9$ (right). Here we set $A_{2}=A_{1}$, $\alpha=1/(4A_{1}^{2})$ (so that $\beta=0$ and $w=\sqrt{1-\gamma^{2}/(16A_{1}^{4})}$), $A_{1}=c_{0}=c_{1}=1$, $c_{2}=0$ and $\gamma=10$. Comparing the first and the third plot, we observe an exchange of the profiles of the absolute values of the components $u_{1}^{\prime}$ and $u_{2}^{\prime}$.
Now
$x\mapsto x-2\pi/\big(\alpha-\mathrm{Im}(\gamma)/|\gamma|^{2}-\mathrm{Im}(w/\gamma)\big)$
exchanges the signs in the expression for the components of $u^{\prime}_{\mathrm{new}}$. Together with
$t\mapsto t-2\pi\alpha\,\frac{\mathrm{Re}(\gamma)/|\gamma|^{2}-\mathrm{Re}(w/\gamma)}{\big(\alpha-\mathrm{Im}(\gamma)/|\gamma|^{2}-\mathrm{Im}(w/\gamma)\big)\big(\mathrm{Im}(w)-\alpha\,\mathrm{Re}(\gamma)\big)}\,$,
it maps $|u^{\prime}_{\mathrm{new},1}|$ into $|u^{\prime}_{\mathrm{new},2}|$, and vice versa. Applying this map twice, $|u^{\prime}_{\mathrm{new},i}|$ at the shifted values of the independent variables coincides with $|u^{\prime}_{\mathrm{new},i}|$ itself, for $i=1,2$. $\Box$
Next we turn to the case $\mathrm{Re}(\gamma)=0$, where the spectrum condition for the Lyapunov equation does not hold.
Example 6.7.
Let $\gamma=-\mathrm{i}\,k$ with $k\in\mathbb{R}\setminus\{0\}$. Then the two roots of the characteristic equation (6.13) are
$r_{\pm}=\pm\frac{1}{2k}\sqrt{2\,\alpha^{2}(2\|A\|^{2}-\alpha^{-1})\,k-\alpha^{2}k^{2}-1}\,.$
The discriminant is real. If it is negative (positive), then $r_{\pm}$ is imaginary (respectively, real). The Lyapunov equation can only be satisfied if just one of the constants $c_{\kappa}$ is different from zero and, moreover, $c_{0}=0$. We thus restrict the solution of the linear system to
$\chi_{1}=e^{\frac{\mathrm{i}}{2}\varphi}\,e^{r(x-\frac{k}{\alpha}\,t)}\,c_{1}\,,\qquad\chi_{2}=0\,$,
where $c_{1}\neq 0$ and $r$ now stands for any of the roots $r_{\pm}$. As a consequence of $\chi_{2}=0$, the resulting solution is equivalent to a solution of the scalar FL equation. The Lyapunov equation reduces to the constraint
$k\,\big|r-\frac{\mathrm{i}}{2}(\alpha+\frac{1}{k})\big|^{2}=\alpha^{2}\|A\|^{2}\,$,
which requires $k>0$. This is only compatible with the expression for $r$ if it is real, in which case the constraint is automatically satisfied. In the following, we therefore assume that $r$ is real. The differential equations for $\Omega$ read
$\Omega_{x}=\frac{|c_{1}|^{2}}{2\,k^{2}}\big(1+2\,\mathrm{i}\,k\,(r+\mathrm{i}\,\frac{\alpha}{2})\big)\,e^{2\theta}\,,\qquad\Omega_{t}=-\frac{k}{\alpha}\,\Omega_{x}\,$,
where
$\theta=r\,\big(x-\frac{k}{\alpha}\,t\big)\,.$
(i) $r\neq 0$. We obtain
$\Omega=\frac{\mathrm{i}\,|c_{1}|^{2}\,(1+\mathrm{i}\,k\,(2r+\mathrm{i}\,\alpha))}{2\,k^{2}\,(\alpha+\mathrm{i}\,(2r+\mathrm{i}\,\alpha))}\,e^{2\theta}+c\,$,
with a constant $c$, which (6.22) requires to be real. The generated solution of the two-component vector FL equation is
$u_{\mu}^{\prime}=A_{\mu}\,e^{\mathrm{i}\,\varphi}\,\Big(1+\frac{k\,|c_{1}|^{2}\big(k\,(2r+\mathrm{i}\,\alpha)-\mathrm{i}\,(1+2\,\alpha\,k)\big)\,(\alpha+\mathrm{i}\,(2r+\mathrm{i}\,\alpha))\,e^{2\theta}}{\alpha\|A\|^{2}\big(\mathrm{i}\,|c_{1}|^{2}(1+\mathrm{i}\,k\,(2r+\mathrm{i}\,\alpha))\,e^{2\theta}+2\,k^{2}\,c\,(\alpha+\mathrm{i}\,(2r+\mathrm{i}\,\alpha))\big)}\Big)\,,\quad\mu=1,2\,.$
(ii) $r=0$. Then we have
$k=2\,\|A\|^{2}-\alpha^{-1}\pm 2\sqrt{\|A\|^{2}(\|A\|^{2}-\alpha^{-1})}\,$,
which is real if and only if $\|A\|^{2}\geq\alpha^{-1}$. In case of the plus sign, $k$ is then positive. In case of the minus sign, $k$ is only positive if $\alpha<0$. Since now $\theta=0$, we obtain
$\Omega=\frac{|c_{1}|^{2}}{2k}(1-k\alpha)\,\big(\frac{x}{k}-\frac{t}{\alpha}\big)+c\,$,
with a constant $c$, for which (6.22) requires $\mathrm{Im}(c)=|c_{1}|^{2}/(2k)$. Obviously, $\mathrm{Re}(c)$ can be absorbed by a shift (respectively redefinition) of one of the independent variables and multiplication with a constant phase factor (which is a special case of the unitary symmetry of the vector FL equation). We obtain the regular quasi-rational solution
$u^{\prime}$ $=$ $e^{\mathrm{i}\,\varphi}\,\Big(1-\frac{\mathrm{i}\,|c_{1}|^{2}}{2\alpha\,\|A\|^{2}\,\Omega}\big(\alpha+\frac{1}{k}\big)\Big)\left(\begin{array}[]{cc}A_{1}\\ A_{2}\end{array}\right)$ (6.32)
$=$ $e^{\mathrm{i}\,\varphi}\,\Big(1-\frac{\mathrm{i}\,k\,(1+k\alpha)}{\|A\|^{2}\,\big((1-k\alpha)\,(\alpha\,x-k\,t)+\mathrm{i}\,k\alpha\big)}\Big)\left(\begin{array}[]{cc}A_{1}\\ A_{2}\end{array}\right)\,$, (6.35)
which represents a soliton rationally concentrated along the line $x=(k/\alpha)\,t$ (hence with velocity $k/\alpha$), where $|u^{\prime}_{\mu}|=|A_{\mu}|\,|1-(\alpha^{-1}+k)/\|A\|^{2}|$. This shows that we also obtain dark solitons in this way. An example is displayed in Fig. 6.777This solution is in the “repeated real root” class, obtained via the “limit technique” in [7], page 206.
Figure 6: Plot of the absolute square of the first component of an $n=1$ quasi-rational solution of the two-component vector FL equation, from the class given by (6.35). Here we chose $\alpha=1$, $A_{1}=1$, $A_{2}=1/\sqrt{3}$ and $k=1/3$. The second component just differs in the amplitude.
$\Box$
In case of a double root, i.e., $r_{+}=r_{-}$, (6.24) does not constitute the complete solution of the corresponding part of the linear system. Then we have to find the missing solution, which is done in the next example. This leads, in particular, to quasi-rational solutions.
Example 6.8.
We have a double root if and only if $r=0$, so that
$\gamma=-\mathrm{i}\,\beta\pm 2\sqrt{\|A\|^{2}(\alpha^{-1}-\|A\|^{2})}\,.$ (6.36)
Then the general solution of (6.12) is
$\chi_{1}=(c_{1}+c_{2}\,\rho)\,e^{\frac{\mathrm{i}}{2}\,\varphi}\,,\qquad\rho=x-\mathrm{i}\,\frac{\gamma}{\alpha}\,t\,$,
with complex constants $c_{1},c_{2}$. If the real discriminant in the expression for $\gamma$ is non-positive, then $\gamma$ is imaginary and the constraint resulting from the Lyapunov equation enforces $c_{2}=0$, which takes us back to the case covered under (ii) in Example 6.7. In the remaining case, where $\alpha^{-1}>\|A\|^{2}$, we have $\mathrm{Re}(\gamma)\neq 0$. We obtain
$\Xi=-\frac{\|A\|}{2\gamma}\,\Big((1+\mathrm{i}\,\alpha\gamma)(c_{1}+c_{2}\,\rho)-2c_{2}\gamma\Big)\,e^{\frac{\mathrm{i}}{2}\,\varphi}\,$,
and the Lyapunov equation has the unique solution
$\Omega=\frac{1}{2\,\mathrm{Re}(\gamma)}\big(\frac{\mathrm{i}\,\gamma^{\ast}}{\alpha^{2}\|A\|^{4}}|\Xi|^{2}+|\chi_{1}|^{2}+|\chi_{2}|^{2}\big)\,.$
According to (6.21), the generated solution of the two-component FL equation is now given by
$u^{\prime}=e^{\mathrm{i}\,\varphi}\,\boldsymbol{U}_{0}\,\left(\begin{array}[]{c}\|A\|+(\alpha\,\|A\|^{2}\,\Omega)^{-1}\chi_{1}^{\ast}\,\Xi\\ (\alpha\,\|A\|^{2}\,\Omega)^{-1}\,\chi_{2}^{\ast}\,\Xi\end{array}\right)\,.$
This solution is quasi-rational if $c_{0}=0$ and then models a rogue wave, rationally localized in space and time and equivalent to a corresponding solution of the scalar FL equation. Fig. 7 displays an example.
Figure 7: Plots of the absolute value of the first component of an $n=1$ quasi-rational solution from the class in Example 6.8. We chose $\alpha=1/10$, $\gamma=8+6\mathrm{i}$, $A_{1}=A_{2}=c_{1}=c_{2}=1$ and $c_{0}=0$.
If $c_{0}\neq 0$, the solution is not equivalent to one of the scalar equation. In this case, we have a superposition with a breather, see Fig. 8 for an example. The double root case treated in this example has already been considered in [12].
Figure 8: Plots of the absolute values of the two components of an $n=1$ solution from the class in Example 6.8. We chose $\alpha=1/3$, $A_{1}=A_{2}=c_{0}=c_{2}=1$ and $c_{1}=10$.
$\Box$
Appendix Appendix D: Superposition of two solutions from the class in Example 6.5 presents some examples of superpositions of two solutions from the class in Example 6.5.
6.1.2 $n=2$
If $\Gamma$ is diagonal and either satisfies the spectrum condition or has imaginary eigenvalues (which is one of the cases where the spectrum condition does not hold), then the corresponding $n=2$ solution is easily obtained by use of results in Section 6.1.1 and an application of the superposition rule, formulated in Remark 5.6. Next we address another case, where $\Gamma$ is diagonal and the spectrum condition is violated, and then the case, where $\Gamma$ is a Jordan block.
Example 6.9.
Let $\Gamma=\mathrm{diag}(\gamma,-\gamma^{\ast})$ with $\mathrm{Re}(\gamma)\neq 0$. Then $R=\mathrm{diag}(\pm r,\pm r^{\ast})$ with
$r=\frac{1}{2}\sqrt{\gamma^{-2}-2\mathrm{i}\,\alpha^{2}\beta\,\gamma^{-1}-\alpha^{2}}\,.$
Although the matrix equation for $R$ thus has four different matrix roots, it is plausible that already two of them are sufficient to obtain the complete solution of the linear system. Assuming $r\neq 0$, this is
$\chi_{1}=e^{\frac{\mathrm{i}}{2}\varphi}\left(\begin{array}[]{c}c_{11}\,e^{\psi}+c_{12}\,e^{-\psi}\\ c_{21}\,e^{\psi^{\ast}}+c_{22}\,e^{-\psi^{\ast}}\end{array}\right)\,,\quad\chi_{2}=e^{\mathrm{i}\,\varphi}\left(\begin{array}[]{c}c_{01}\,e^{\frac{1}{2}(\gamma^{-1}x+\gamma\,t)}\\ c_{02}\,e^{-\frac{1}{2}(\gamma^{\ast-1}x+\gamma^{\ast}\,t)}\end{array}\right)\,$,
where
$\psi=r\,(x-\mathrm{i}\,\alpha^{-1}\gamma\,t)\,.$
We further obtain
$\Xi=\frac{\|A\|}{2}\,e^{\frac{\mathrm{i}}{2}\varphi}\left(\begin{array}[]{c}c_{11}\,(2r-\gamma^{-1}-\mathrm{i}\,\alpha)\,e^{\psi}-c_{12}\,(2r+\gamma^{-1}+\mathrm{i}\,\alpha)\,e^{-\psi}\\ c_{21}\,(2r^{\ast}+\gamma^{\ast-1}-\mathrm{i}\,\alpha)\,e^{\psi^{\ast}}-c_{22}\,(2r^{\ast}-\gamma^{\ast-1}+\mathrm{i}\,\alpha)\,e^{-\psi^{\ast}}\end{array}\right)\,.$
The Lyapunov equation determines the diagonal entries of $\Omega$,
$\Omega_{11}$ $=$ $\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\Xi\Xi^{\dagger}\,\Gamma^{\dagger}+\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger}\Big)_{11}$
$=$ $\frac{1}{2\mathrm{Re}(\gamma)}\Big(\big(\frac{\mathrm{i}\,\gamma^{\ast}}{4\alpha^{2}\|A\|^{2}}|2r-\gamma^{-1}-\mathrm{i}\,\alpha|^{2}+1\big)\,|c_{11}|^{2}e^{2\mathrm{Re}(\psi)}$
$-\big(\frac{\mathrm{i}\,\gamma^{\ast}}{4\alpha^{2}\|A\|^{2}}(2r-\gamma^{-1}-\mathrm{i}\,\alpha)\,(2r^{\ast}+\gamma^{\ast-1}-\mathrm{i}\,\alpha)-1\big)\,c_{11}c_{12}^{\ast}e^{2\mathrm{i}\,\mathrm{Im}(\psi)}$
$-\big(\frac{\mathrm{i}\,\gamma^{\ast}}{4\alpha^{2}\|A\|^{2}}(2r+\gamma^{-1}+\mathrm{i}\,\alpha)\,(2r^{\ast}-\gamma^{\ast-1}+\mathrm{i}\,\alpha)-1\big)\,c_{11}^{\ast}c_{12}e^{-2\mathrm{i}\,\mathrm{Im}(\psi)}$
$+\big(\frac{\mathrm{i}\,\gamma^{\ast}}{4\alpha^{2}\|A\|^{2}}|2r+\gamma^{-1}+\mathrm{i}\,\alpha|^{2}+1\big)\,|c_{12}|^{2}e^{-2\mathrm{Re}(\psi)}+|c_{01}|^{2}e^{\mathrm{Re}(\gamma^{-1}x+\gamma t)}\Big)\,$,
$\Omega_{22}$ $=$ $-\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\Xi\Xi^{\dagger}\,\Gamma^{\dagger}+\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger}\Big)_{22}$
$=$ $\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}\,\gamma}{4\alpha^{2}\|A\|^{2}}\big(|2r+\gamma^{-1}+\mathrm{i}\,\alpha|^{2}-1\big)\,|c_{21}|^{2}e^{2\mathrm{Re}(\psi)}$
$-\big(\frac{\mathrm{i}\,\gamma}{4\alpha^{2}\|A\|^{2}}(2r+\gamma^{-1}+\mathrm{i}\,\alpha)\,(2r^{\ast}-\gamma^{\ast-1}+\mathrm{i}\,\alpha)+1\big)\,c_{21}^{\ast}c_{22}e^{2\mathrm{i}\,\mathrm{Im}(\psi)}$
$-\big(\frac{\mathrm{i}\,\gamma}{4\alpha^{2}\|A\|^{2}}(2r-\gamma^{-1}-\mathrm{i}\,\alpha)\,(2r^{\ast}+\gamma^{\ast-1}-\mathrm{i}\,\alpha)+1\big)\,c_{21}c_{22}^{\ast}e^{-2\mathrm{i}\,\mathrm{Im}(\psi)}$
$+\frac{\mathrm{i}\,\gamma}{4\alpha^{2}\|A\|^{2}}\big(|2r-\gamma^{-1}-\mathrm{i}\,\alpha|^{2}-1\big)\,|c_{22}|^{2}e^{-2\mathrm{Re}(\psi)}-|c_{02}|^{2}e^{-\mathrm{Re}(\gamma^{-1}x+\gamma t)}\Big)\,$,
and imposes the constraints
$\Big(\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\Xi\Xi^{\dagger}\,\Gamma^{\dagger}+\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger}\Big)_{ij}=0\qquad\mbox{for}\quad i\neq j\,.$
The latter amount to
$(c_{11}c_{22}^{\ast}+c_{12}c_{21}^{\ast})\,\big(4\alpha^{2}\|A\|^{2}\gamma+\mathrm{i}\,(1+\mathrm{i}\,\alpha\gamma)^{2}\big)-2\mathrm{i}\,\gamma\,r\,(c_{11}c_{22}^{\ast}-c_{12}c_{21}^{\ast})\,(1+\mathrm{i}\,\alpha\gamma)+2c_{01}c_{02}^{\ast}\,\alpha^{2}\|A\|^{2}\gamma=0\,.$
The differential equations for $\Omega$ lead to
$\Omega_{12x}$ $=$ $\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\big(\Xi_{1x}-\frac{1}{2}\gamma^{-1}\Xi_{1}\big)\,\Xi_{2}^{\ast}\,,\qquad\Omega_{12t}=-\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\big(\alpha\|A\|^{3}\chi_{12}^{\ast}-\gamma\,\Xi_{2}^{\ast}\big)\,\Xi_{1}\,$,
$\Omega_{21x}$ $=$ $\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\big(\Xi_{2x}+\frac{1}{2}\gamma^{\ast-1}\Xi_{2}\big)\,\Xi_{1}^{\ast}\,,\qquad\Omega_{21t}=-\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\big(\alpha\|A\|^{3}\chi_{11}^{\ast}+\gamma^{\ast}\Xi_{1}^{\ast}\big)\,\Xi_{2}\,$,
which leads to
$\Omega_{12}$ $=$ $\frac{\mathrm{i}}{16\alpha^{2}\|A\|^{2}\gamma^{3}\,r}\Big(c_{11}c_{21}^{\ast}\big(1+\gamma\,(\mathrm{i}\,\alpha+2r)\big)\,\big(1+\gamma^{2}\,(\alpha^{2}+4r^{2})-4\gamma\,r\big)\,e^{2\psi}$
$-c_{12}c_{22}^{\ast}\big(1+\gamma\,(\mathrm{i}\,\alpha-2r)\big)\,\big(1+\gamma^{2}(\alpha^{2}+4r^{2})+4\gamma\,r\big)\,e^{-2\psi}\Big)$
$+\frac{\mathrm{i}\,x}{8\alpha^{2}\|A\|^{2}\gamma^{3}}\Big(c_{12}c_{21}^{\ast}\big(1+\gamma\,(\mathrm{i}\,\alpha+2r)\big)\,\big(1+\gamma^{2}\,(\alpha^{2}+4r^{2})+4\gamma\,r\big)$
$+c_{11}c_{22}^{\ast}\,\big(1+\gamma\,(\mathrm{i}\,\alpha-2r)\big)\,\big(1+\gamma^{2}(\alpha^{2}+4r^{2})-4\gamma\,r\big)\Big)$
$+\frac{\mathrm{i}\,t}{4\alpha^{2}\|A\|^{2}\gamma}\Big(c_{12}c_{21}^{\ast}\,\big(2\alpha\|A\|^{2}\,(1+\gamma\,(\mathrm{i}\,\alpha+2r)\big)+\gamma^{2}\,\big(\alpha^{2}-4r\,(\mathrm{i}\,\alpha+r)\big)-2\gamma\,(\mathrm{i}\,\alpha+2r)-1\big)$
$+c_{11}c_{22}^{\ast}\,\big(2\alpha\|A\|^{2}\,\big(1+\gamma\,(\mathrm{i}\,\alpha-2r)\big)+\gamma^{2}\,\big(\alpha^{2}+4r\,(\mathrm{i}\,\alpha-r)\big)-2\gamma\,(\mathrm{i}\,\alpha-2r)-1\big)\Big)+\tilde{c}_{12}\,$,
$\Omega_{21}$ $=$ $-\frac{\mathrm{i}}{16\alpha^{2}\|A\|^{2}\gamma^{\ast 3}\,r^{\ast}}\Big(c_{11}^{\ast}c_{21}\big(1-\gamma^{\ast}(\mathrm{i}\,\alpha+2r^{\ast})\big)\,\big(1+\gamma^{\ast 2}\,(\alpha^{2}+4r^{\ast 2})+4\gamma^{\ast}r^{\ast}\big)\,e^{2\psi^{\ast}}$
$-c_{12}^{\ast}c_{22}\big(1-\gamma^{\ast}(\mathrm{i}\,\alpha-2r^{\ast})\big)\,\big(1+\gamma^{\ast 2}(\alpha^{2}+4r^{\ast 2})-4\gamma^{\ast}r^{\ast}\big)\,e^{-2\psi^{\ast}}\Big)$
$+\frac{\mathrm{i}\,x}{8\alpha^{2}\|A\|^{2}\gamma^{\ast 3}}\Big(2\gamma^{\ast}r^{\ast}(c_{11}^{\ast}c_{22}-c_{12}^{\ast}c_{21})\,\big(\gamma^{\ast 2}(\alpha^{2}+4r^{\ast 2})-2\mathrm{i}\,\alpha\gamma^{\ast}+3\big)$
$+(c_{11}^{\ast}c_{22}+c_{12}^{\ast}c_{21})\,\big(4r^{\ast 2}(\mathrm{i}\,\alpha\gamma^{\ast 3}-3\gamma^{\ast 2})+(1+\alpha^{2}\gamma^{\ast 2})\,(\mathrm{i}\,\alpha\gamma^{\ast}-1)\big)\Big)$
$+\frac{\mathrm{i}\,t}{4\alpha^{2}\|A\|^{2}\gamma^{\ast}}\Big(4\gamma^{\ast}r^{\ast}(c_{11}^{\ast}c_{22}-c_{12}^{\ast}c_{21})\,\big(\alpha\,(\mathrm{i}\,\gamma^{\ast}+\|A\|^{2})-1\big)$
$+(c_{11}^{\ast}c_{22}+c_{12}^{\ast}c_{21})\,\big(1-\gamma^{\ast 2}(\alpha^{2}-4r^{\ast 2})+2\alpha\|A\|^{2}(\mathrm{i}\,\alpha\gamma^{\ast}-1)-2\mathrm{i}\,\alpha\gamma^{\ast}\big)\Big)+\tilde{c}_{21}\,$,
with complex constants $\tilde{c}_{12}$ and $\tilde{c}_{21}$. The remaining condition (4.5) becomes
$\gamma\,(\tilde{c}_{12}-\tilde{c}_{21}^{\ast})=c_{01}c_{02}^{\ast}+c_{11}c_{22}^{\ast}+c_{12}c_{21}^{\ast}\,.$
The resulting solution of the two-component FL equation is obtained by inserting the expressions for $\chi_{1}$, $\chi_{2}$, $\Xi$ and $\Omega$ in (6.21). In addition to the exponential dependence of the solution via $e^{\psi}$ and its complex conjugate, there is also a rational dependence on the variables $x$ and $t$. Fig. 9 displays an example from this class.
Figure 9: Plots of the absolute values of the two components of a solution from the class in Example 6.9, showing a bright, respectively dark soliton, a breather and a periodic rogue wave. We chose $\alpha=1$, $A_{1}=1/2$, $A_{2}=\sqrt{3}/2$, $\gamma=1$ (so that $r=-1/2+\mathrm{i}/2$), $c_{01}=c_{02}=c_{12}=c_{21}=c_{22}=\tilde{c}_{12}=1$, $c_{11}=\tilde{c}_{21}=-1/2+3\mathrm{i}/2$.
$\Box$
Example 6.10.
Let $\Gamma$ be the $2\times 2$ lower triangular Jordan block with eigenvalue $\gamma$. Then
$R=\pm\left(\begin{array}[]{cc}r&0\\ -(1-\mathrm{i}\,\alpha^{2}\beta\gamma)(4\gamma^{3}\,r)^{-1}&r\end{array}\right)\,,\qquad r=\frac{1}{2\gamma}\sqrt{1-\alpha^{2}\gamma\,(\gamma+2\mathrm{i}\,\beta)}\,.$ (6.42)
Let $r\neq 0$. We obtain
$\chi_{1}$ $=$ $e^{\frac{1}{2}\mathrm{i}\,\varphi}\left(\begin{array}[]{c}c_{11}\,e^{\psi}+c_{21}\,e^{-\psi}\\ (4\gamma^{3}r)^{-1}\Big(\big(-c_{11}\,\rho+4c_{12}\gamma^{3}r\big)\,e^{\psi}+\big(c_{21}\,\rho+4c_{22}\gamma^{3}r\big)\,e^{-\psi}\Big)\end{array}\right)\,$,
$\chi_{2}$ $=$ $e^{\mathrm{i}\,\varphi}\,e^{\frac{1}{2}(\gamma^{-1}x+\gamma\,t)}\left(\begin{array}[]{c}c_{01}\\ c_{02}-\frac{1}{2}c_{01}\gamma^{-1}(\gamma^{-1}x-\gamma\,t)\end{array}\right)\,$,
$\Xi$ $=$ $\frac{\|A\|}{8\gamma^{4}r}\,e^{\frac{\mathrm{i}}{2}\varphi}\left(\begin{array}[]{c}-4\gamma^{3}r\,\big(c_{11}\,b_{-}\,e^{\psi}+c_{21}\,b_{+}\,e^{-\psi}\big)\\ \big(c_{11}(b_{-}\,\rho-2\gamma\,d_{-})-4c_{12}b_{-}\gamma^{3}r\big)\,e^{\psi}-\big(c_{21}(b_{+}\,\rho-2\gamma\,d_{+})+4c_{22}b_{+}\gamma^{3}r\big)\,e^{-\psi}\end{array}\right)\,$,
where
$\rho=(1-\mathrm{i}\,\alpha^{2}\beta\gamma)\,x+\alpha\gamma^{2}(\beta-\mathrm{i}\,\gamma)\,t\,,\qquad\psi=r\,\big(x-\mathrm{i}\,\frac{\gamma}{\alpha}t\big)\,$,
$b_{\pm}=1\pm 2\gamma\,r+\mathrm{i}\,\alpha\,\gamma\,,\qquad d_{\pm}=1\pm 2\gamma\,r-\mathrm{i}\,\alpha^{2}\beta\,\gamma\,.$
It is now straightforward to elaborate the equations determining $\Omega$ and then the generated solution (6.21) of the two-component vector FL equation. If $\mathrm{Re}(\gamma)\neq 0$, then $\Omega$ is completely determined by the Lyapunov equation and given by Example A.2. If $\mathrm{Re}(\gamma)=0$, we have to employ the differential equations for $\Omega$ in addition, a more tedious task. We skip the details. $\Box$
Example 6.11.
Again, let $\Gamma$ be the $2\times 2$ lower triangular Jordan block with eigenvalue $\gamma$, so that $R$ and $r$ are given by (6.42). But now we consider the double root case $r=0$. Then the expression on the right hand side of (6.13) is nilpotent and does not possess a square root, and $\gamma$ is fixed by (6.36). The general solution of the linear system is now given by $\tilde{\chi}_{i}=(\tilde{\chi}_{i1},\tilde{\chi}_{i2})^{T}$, $i=1,2$, with
$\tilde{\chi}_{11}$ $=$ $a_{1}\,\rho+a_{0}\qquad\mbox{where}\quad\rho=x-\frac{\mathrm{i}}{2\gamma}\,t\,$,
$\tilde{\chi}_{12}$ $=$ $-\frac{(1-\mathrm{i}\,\alpha^{2}\beta\gamma)}{12\gamma^{3}}(a_{1}\,\rho^{3}+3a_{0}\,\rho^{2})+c_{2}\rho+c_{1}+\frac{6\,\mathrm{i}\,a_{1}}{12\gamma^{2}}\,t\,$,
$\tilde{\chi}_{21}$ $=$ $b_{1}\,e^{\frac{1}{2}(\gamma^{-1}x+\gamma\,t)}\,,\qquad\tilde{\chi}_{22}=\big(b_{2}-\frac{1}{2}b_{1}(\gamma^{-2}x-t)\big)\,e^{\frac{1}{2}(\gamma^{-1}x+\gamma\,t)}\,$,
with constants $a_{0},a_{1},b_{1},b_{2},c_{1},c_{2}$. Then we obtain the following components of $\Xi$,
$\Xi_{1}$ $=$ $-\frac{\|A\|}{2\gamma}\,e^{\mathrm{i}\,\varphi}\,\Big(a_{1}\,(\rho-2\gamma)+a_{0}\Big)\,$,
$\Xi_{2}$ $=$ $\|A\|\,e^{\mathrm{i}\,\varphi}\,\Big(\frac{a_{1}}{24\,\gamma^{4}}(1-\mathrm{i}\,\alpha^{2}\beta\gamma)\,\rho^{3}+\frac{a_{0}-2a_{1}\gamma}{8\gamma^{4}}(1-\mathrm{i}\,\alpha^{2}\beta\gamma)\,\rho^{2}$
$-\frac{1}{2\gamma^{3}}\big(a_{0}(1-\mathrm{i}\,\alpha^{2}\beta\gamma)-a_{1}\gamma+c_{2}\gamma^{2}\big)\,\rho+\frac{1}{4\gamma^{3}}\big(-\mathrm{i}\,a_{1}\,t+2a_{0}\gamma-2c_{1}\gamma^{2}+4c_{2}\gamma^{3}\big)\,.$
If $\alpha^{-1}>\|A\|^{2}$, then $\mathrm{Re}(\gamma)\neq 0$ (cf. Example 6.8), and $\Omega$ is completely determined by the Lyapunov equation alone,
$\Omega_{11}$ $=$ $\frac{1}{2\,\mathrm{Re}(\gamma)}\Big(\frac{\gamma^{\ast}}{\alpha^{2}\|A\|^{4}}|\Xi_{1}|^{2}+|\tilde{\chi}_{11}|^{2}+|\tilde{\chi}_{21}|^{2}\Big)\,$,
$\Omega_{12}$ $=$ $\frac{1}{4\,\mathrm{Re}(\gamma)^{2}\alpha^{2}\|A\|^{4}}\Big(\gamma\,|\Xi_{1}|^{2}+2\,\mathrm{Re}(\gamma)\,\gamma^{\ast}\,\Xi_{1}\,\Xi_{2}^{\ast}$
$-\alpha^{2}\|A\|^{4}\big(|\tilde{\chi}_{11}|^{2}+|\tilde{\chi}_{21}|^{2}-2\,\mathrm{Re}(\gamma)\,(\tilde{\chi}_{11}\tilde{\chi}_{12}^{\ast}+\tilde{\chi}_{21}\tilde{\chi}_{22}^{\ast})\big)\Big)\,$,
$\Omega_{21}$ $=$ $\frac{1}{4\,\mathrm{Re}(\gamma)^{2}\alpha^{2}\|A\|^{4}}\Big(-\gamma^{\ast}\,|\Xi_{1}|^{2}+2\,\mathrm{Re}(\gamma)\,\gamma^{\ast}\,\Xi_{1}^{\ast}\,\Xi_{2}$
$-\alpha^{2}\|A\|^{4}\big(|\tilde{\chi}_{11}|^{2}+|\tilde{\chi}_{21}|^{2}-2\,\mathrm{Re}(\gamma)\,(\tilde{\chi}_{11}^{\ast}\tilde{\chi}_{12}+\tilde{\chi}_{21}^{\ast}\tilde{\chi}_{22})\big)\Big)\,$,
$\Omega_{22}$ $=$ $\frac{1}{8\,\mathrm{Re}(\gamma)^{3}\alpha^{2}\|A\|^{4}}\Big(-2\mathrm{i}\,\mathrm{Im}(\gamma)\,|\Xi_{1}|^{2}+4\mathrm{Re}(\gamma)^{2}\,\gamma^{\ast}|\Xi_{2}|^{2}+4\mathrm{i}\,\mathrm{Re}(\gamma)\,\mathrm{Im}(\gamma\,\Xi_{1}^{\ast}\,\Xi_{2})$
$+2\alpha^{2}\|A\|^{4}\big(|\tilde{\chi}_{11}|^{2}+|\tilde{\chi}_{21}|^{2}+2\mathrm{Re}(\gamma)^{2}\,(|\tilde{\chi}_{12}|^{2}+|\tilde{\chi}_{22}|^{2})-2\,\mathrm{Re}(\gamma)\,\mathrm{Re}(\tilde{\chi}_{11}\tilde{\chi}_{12}^{\ast}+\tilde{\chi}_{21}\tilde{\chi}_{22}^{\ast})\big)\Big)\,.$
Now (6.21) determines a solution of the two-component vector FL equation, which is a second order version of the solution in Example 6.8. It has already been derived in a different way in [12]. This solution is quasi-rational if and only if $\tilde{\chi}_{2}=0$, in which case it is equivalent to a solution of the scalar FL equation. $\Box$
Remark 6.12.
The solution set of a polynomial matrix equation, like (6.13), is in general more intricate than for polynomial equations over a field. Here we take a look at an example in our context. Let $\Gamma=\mathrm{diag}(\gamma_{1},\gamma_{2})$, where $\gamma_{1},\gamma_{2}$ have the form (6.36). In this case,
$R_{1}=\left(\begin{array}[]{cc}0&z_{1}\\ 0&0\end{array}\right)\,,\qquad R_{2}=\left(\begin{array}[]{cc}0&0\\ z_{2}&0\end{array}\right)\,$,
solve (6.13), which is now $R^{2}=0$, for arbitrary $z_{1},z_{2}\in\mathbb{C}$. Assuming $z_{1}z_{2}\neq 0$, (6.24) yields
$\chi_{1}=e^{\frac{\mathrm{i}}{2}\,\varphi}\left(\begin{array}[]{c}c_{11}+c_{12}\,(x-\frac{\mathrm{i}}{\alpha}\gamma_{1}t)\\ c_{21}+c_{22}\,(x-\frac{\mathrm{i}}{\alpha}\gamma_{2}t)\end{array}\right)\,.$
Here $z_{1}$ and $z_{2}$ have been absorbed by redefinition of the constant 2-vectors $c_{\kappa}$. It is thus obvious that, in this way, we get directly a solution describing a (nonlinear) superposition of two solitons of the kind given in Example 6.8, without having to reconsider the linear system in order to find missing solutions. $\Box$
6.2 The generic plane wave case $\alpha_{1}\neq\alpha_{2}$
If $\alpha_{1}\neq\alpha_{2}$ and if the matrix in the last brackets in (6.8) is invertible, the latter equation can be solved for $\chi_{2}$,
$\chi_{2}$ $=$ $\frac{\|A\|^{2}}{(\alpha_{2}-\alpha_{1})\,A_{1}A_{2}}\Big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,\Gamma^{-1}-\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}I_{n}\Big)^{-1}\Big(\chi_{1xx}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}\,\chi_{1x}-\frac{1}{4}\Gamma^{-2}\chi_{1}$
$+\frac{\mathrm{i}}{2\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)\,\Gamma^{-1}\chi_{1}\Big)\,.$
Inserted in (6.7), this yields a third order ODE for $\chi_{1}$,
$\chi_{1xxx}-\big(\mathrm{i}\,(\alpha_{1}+\alpha_{2})I_{n}+\frac{1}{2}\Gamma^{-1}\big)\,\chi_{1xx}-\big(\frac{1}{4}\Gamma^{-2}-\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}^{2}A\,\Gamma^{-1}+\alpha_{1}\alpha_{2}I_{n}\big)\,\chi_{1x}$
$+\Big(\frac{1}{8}\Gamma^{-3}+\frac{1}{4}\mathrm{i}\,\big(\alpha_{1}+\alpha_{2}-2A^{\dagger}\boldsymbol{\alpha}^{2}A\big)\Gamma^{-2}+\alpha_{1}\alpha_{2}(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2}\big)\Gamma^{-1}\Big)\,\chi_{1}=0\,.$
(5.14) reads
$\chi_{1t}$ $=$ $-(\mathrm{i}\,\|A\|^{2}I_{n}-\frac{1}{2}\Gamma)\,\chi_{1}+\frac{\mathrm{i}}{\|A\|^{2}}\big(A^{\dagger}\boldsymbol{\beta}\,A\,\chi_{1}+(\beta_{2}-\beta_{1})\,A_{1}A_{2}\,\chi_{2}\big)-\mathrm{i}\,\frac{\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\Big(\Gamma\,\chi_{1x}$
$-\frac{\mathrm{i}}{\|A\|^{2}}\Gamma\big(A^{\dagger}\boldsymbol{\alpha}\,A\,\chi_{1}+(\alpha_{2}-\alpha_{1})\,A_{1}A_{2}\,\chi_{2}\big)-\frac{1}{2}\chi_{1}\Big)\,$,
$\chi_{2t}$ $=$ $\frac{1}{2}\Gamma\,\chi_{2}+\frac{\mathrm{i}}{\|A\|^{2}}\big((\beta_{2}-\beta_{1})\,A_{1}^{\ast}A_{2}^{\ast}\,\chi_{1}+(\beta_{1}|A_{2}|^{2}+\beta_{2}|A_{1}|^{2})\,\chi_{2}\big)\,$,
where now
$\beta_{1}=2|A_{1}|^{2}+(1+\alpha_{2}/\alpha_{1})|A_{2}|^{2}-\frac{1}{\alpha_{1}}\,,\qquad\beta_{2}=(1+\alpha_{1}/\alpha_{2})|A_{1}|^{2}+2|A_{2}|^{2}-\frac{1}{\alpha_{2}}\,.$
After elimination of $\chi_{2}$ in the equation for $\chi_{1t}$, we get
$\alpha_{1}\alpha_{2}\,\chi_{1t}=\Gamma\chi_{1xx}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})\Gamma\chi_{1x}-\big(\frac{1}{2}\alpha_{1}\alpha_{2}\,\Gamma+\frac{1}{4}\Gamma^{-1}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2})I_{n}\big)\,\chi_{1}\,$,
which determines the $t$-dependence of $\chi_{1}$. By consistency, the equation for $\chi_{2t}$ should be identically satisfied as a consequence of the others. This can indeed be verified by a tedious computation. The linear system is now (in special cases not necessarily completely) solved by
$\chi_{1}$ $=$ $\sum_{\kappa}e^{R_{\kappa}x+S_{\kappa}t}c_{\kappa}\,$, (6.48)
$\chi_{2}$ $=$ $\frac{\|A\|^{2}}{(\alpha_{2}-\alpha_{1})\,A_{1}A_{2}}\Big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,\Gamma^{-1}-\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}I_{n}\Big)^{-1}\sum_{\kappa}\Big(R_{\kappa}^{2}-\frac{\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}\,R_{\kappa}-\frac{1}{4}\Gamma^{-2}$ (6.49)
$+\frac{\mathrm{i}}{2\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)\,\Gamma^{-1}\Big)\,e^{R_{\kappa}x+S_{\kappa}t}c_{\kappa}\,$,
where $\kappa$ runs over a set numbering different roots $R_{\kappa}$, commuting with $\Gamma$, of the cubic $n\times n$ matrix equation
$R^{3}-\big(\mathrm{i}\,(\alpha_{1}+\alpha_{2})I_{n}+\frac{1}{2}\Gamma^{-1}\big)R^{2}-\big(\frac{1}{4}\Gamma^{-2}-\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}^{2}A\,\Gamma^{-1}+\alpha_{1}\alpha_{2}I_{n}\big)R$
$+\frac{1}{8}\Gamma^{-3}+\frac{1}{4}\mathrm{i}\,\big(\alpha_{1}+\alpha_{2}-2A^{\dagger}\boldsymbol{\alpha}^{2}A\big)\Gamma^{-2}+\alpha_{1}\alpha_{2}\,(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2}\big)\Gamma^{-1}=0\,.$
Further, $\{c_{\kappa}\}$ are constant $n$-component column vectors and
$S_{\kappa}=\frac{1}{\alpha_{1}\alpha_{2}}\Big(\Gamma R_{\kappa}^{2}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})\Gamma R_{\kappa}-\frac{1}{2}\alpha_{1}\alpha_{2}\Gamma-\frac{1}{4}\Gamma^{-1}+\mathrm{i}\,(\alpha_{1}+\alpha_{2})(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2})I_{n}\Big)\,.$
It remains to determine $\Omega$. Then, finally,
$u^{\prime}$ $=$ $\left(\begin{array}[]{cc}A_{1}\,e^{\mathrm{i}\,\varphi_{1}}&-A_{2}^{\ast}\,e^{\mathrm{i}\,\varphi_{1}}\\ A_{2}\,e^{\mathrm{i}\,\varphi_{2}}&A_{1}^{\ast}\,e^{\mathrm{i}\,\varphi_{2}}\end{array}\right)\left[\left(\begin{array}[]{c}1\\ 0\end{array}\right)+\frac{1}{\|A\|\,A^{\dagger}\boldsymbol{\alpha}A}\,\left(\begin{array}[]{c}\chi_{1}^{\dagger}\,\Omega^{-1}\Xi\\ \chi_{2}^{\dagger}\,\Omega^{-1}\Xi\end{array}\right)\right]$ (6.56)
$=$ $\left(\begin{array}[]{cc}e^{\mathrm{i}\,\varphi_{1}}&0\\ 0&e^{\mathrm{i}\,\varphi_{2}}\end{array}\right)\left(\begin{array}[]{cc}A_{1}+(\|A\|\,A^{\dagger}\boldsymbol{\alpha}A)^{-1}(A_{1}\chi_{1}^{\dagger}-A_{2}^{\ast}\chi_{2}^{\dagger})\,\Omega^{-1}\Xi\\[4.0pt] A_{2}+(\|A\|\,A^{\dagger}\boldsymbol{\alpha}A)^{-1}(A_{2}\chi_{1}^{\dagger}+A_{1}^{\ast}\chi_{2}^{\dagger})\,\Omega^{-1}\Xi\end{array}\right)\,$, (6.61)
with $\Xi$ given by (6.9), is a solution of the two-component vector FL equation.
Again, we have to distinguish the following two cases.
(1) If the spectrum condition for $\Gamma$ holds, then the Lyapunov equation (5.17) possesses a unique solution, irrespective of the $n\times n$ matrix on its right hand side. In particular, if $\Gamma=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n})$, with $\gamma_{i}\neq-\gamma_{j}^{\ast}$, $i,j=1,\ldots,n$, the solution $\Omega$ is given by the Cauchy-like matrix with components
$\Omega_{ij}=\frac{\mathrm{i}\,\Xi_{i}\,\Xi_{j}^{\ast}\,\gamma_{j}^{\ast}+(A^{\dagger}\boldsymbol{\alpha}A)^{2}\big(\chi_{1i}\chi_{1j}^{\ast}+\chi_{2i}\chi_{2j}^{\ast}\big)}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}\,(\gamma_{i}+\gamma_{j}^{\ast})}\,$,
where
$\Xi_{i}=\|A\|\sum_{\kappa}\Big(r_{i\kappa}-\frac{1}{2}\gamma_{i}^{-1}-\frac{\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}}\Big)\,e^{r_{i\kappa}x+s_{i\kappa}t}\,c_{\kappa i}-\frac{\mathrm{i}\,(\alpha_{2}-\alpha_{1})A_{1}A_{2}}{\|A\|}\chi_{2i}\,$,
writing $R_{\kappa}=\mathrm{diag}(r_{1\kappa},\ldots,r_{n\kappa})$ and $S_{\kappa}=\mathrm{diag}(s_{1\kappa},\ldots,s_{n\kappa})$.
(2) If the spectrum condition does not hold, the Lyapunov equation (5.17) constrains its right hand side. Moreover, it does not (fully) determine $\Omega$. In this case, we have to take (5.8) into account, i.e.,
$\Omega_{x}$ $=$ $\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\,\Xi_{x}\,\Xi^{\dagger}+\frac{1}{2}\Gamma^{-1}\chi_{1}\chi_{1}^{\dagger}\Gamma^{\dagger-1}$
$-\Big(\frac{A_{1}^{\ast}A_{2}^{\ast}}{A^{\dagger}\boldsymbol{\alpha}A}(\alpha_{2}-\alpha_{1})\big(\chi_{1x}-\frac{1}{2}\Gamma^{-1}\chi_{1}\big)-\chi_{2x}+\frac{\mathrm{i}\,\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\alpha_{1}\alpha_{2}\chi_{2}\Big)\chi_{2}^{\dagger}\,\Gamma^{\dagger-1}\,$,
$\Omega_{t}$ $=$ $-\frac{\mathrm{i}}{2(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\,\Xi\,\,\Xi^{\dagger}\,\Gamma^{\dagger}-\frac{\mathrm{i}\,\|A\|}{A^{\dagger}\boldsymbol{\alpha}A}\,\Xi\,\chi_{1}^{\dagger}+\frac{1}{2}(\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger})\,$,
and (5.7), which is (6.22).
6.2.1 $n=1$
For $n=1$, we write $\Gamma=\gamma$ and $R_{\kappa}=r_{\kappa}$, $S_{\kappa}=s_{\kappa}$, so that
$\chi_{1}$ $=$ $\sum_{\kappa}e^{r_{\kappa}x+s_{\kappa}t}\,c_{\kappa}\,$,
$\chi_{2}$ $=$ $\frac{\gamma\,\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A}{(\alpha_{2}-\alpha_{1})A_{1}A_{2}\big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,A^{\dagger}\boldsymbol{\alpha}A-\gamma\,\alpha_{1}\alpha_{2}\|A\|^{2}\big)}\sum_{\kappa}\Big(r_{\kappa}^{2}-\frac{\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}\,r_{\kappa}-\frac{1}{4}\gamma^{-2}$
$+\frac{\mathrm{i}}{2\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)\,\gamma^{-1}\Big)\,e^{r_{\kappa}x+s_{\kappa}t}\,c_{\kappa}\,$,
assuming $\gamma\neq\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,A^{\dagger}\boldsymbol{\alpha}A/(\alpha_{1}\alpha_{2}\|A\|^{2})$. Here $\{r_{\kappa}\}$ is a set of different roots of the cubic characteristic equation
$r^{3}-\big(\mathrm{i}\,(\alpha_{1}+\alpha_{2})+\frac{1}{2}\gamma^{-1}\big)r^{2}-\big(\frac{1}{4}\gamma^{-2}-\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}^{2}A\gamma^{-1}+\alpha_{1}\alpha_{2}\big)r$
$+\frac{1}{8}\gamma^{-3}+\frac{1}{4}\mathrm{i}\,\big(\alpha_{1}+\alpha_{2}-2A^{\dagger}\boldsymbol{\alpha}^{2}A\big)\gamma^{-2}+\alpha_{1}\alpha_{2}\,(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2}\big)\gamma^{-1}=0\,$, (6.62)
and
$s_{\kappa}=\frac{1}{\alpha_{1}\alpha_{2}}\Big(\gamma\,r_{\kappa}^{2}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})\gamma\,r_{\kappa}-\frac{1}{2}\alpha_{1}\alpha_{2}\gamma-\frac{1}{4}\gamma^{-1}+\mathrm{i}\,(\alpha_{1}+\alpha_{2})(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2})\Big)\,.$ (6.63)
If there are three different roots $r_{\kappa}$, $\kappa=1,2,3$, then the above solution of the linear system is the general solution. Otherwise, the missing solutions still have to be found, which is addressed in subsections 6.2.3 and 6.2.4.
The Darboux transformation yields the following solution of the two-component vector FL equation,
$u^{\prime}=\left(\begin{array}[]{cc}e^{\mathrm{i}\,\varphi_{1}}\,\big[A_{1}+(\|A\|\,A^{\dagger}\boldsymbol{\alpha}A\,\Omega)^{-1}(A_{1}\chi_{1}^{\dagger}-A_{2}^{\ast}\chi_{2}^{\dagger})\,\Xi\big]\\[4.0pt] e^{\mathrm{i}\,\varphi_{2}}\,\big[A_{2}+(\|A\|\,A^{\dagger}\boldsymbol{\alpha}A\,\Omega)^{-1}(A_{2}\chi_{1}^{\dagger}+A_{1}^{\ast}\chi_{2}^{\dagger})\,\Xi\big]\end{array}\right)\,$, (6.66)
where now
$\Xi=\|A\|\Big(\chi_{1x}-(\frac{1}{2}\gamma^{-1}+\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}})\,\chi_{1}\Big)-\mathrm{i}\,(\alpha_{2}-\alpha_{1})\frac{A_{1}A_{2}}{\|A\|}\,\chi_{2}\,.$
It only remains to determine $\Omega$.
Let $\mathrm{Re}(\gamma)\neq 0$. Then we get
$\Omega=\frac{\mathrm{i}\,\gamma^{\ast}\,|\Xi|^{2}+(A^{\dagger}\boldsymbol{\alpha}A)^{2}(|\chi_{1}|^{2}+|\chi_{2}|^{2})}{2\mathrm{Re}(\gamma)\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\,$,
which is everywhere non-zero, so that the generated solution $u^{\prime}$ is regular on $\mathbb{R}^{2}$. Fig. 10 shows plots of the absolute values of the two components of $u^{\prime}$ for an example from this class of solutions of the two-component vector FL equation.
Figure 10: Plots of the absolute value of the two components of a solution from the class in Section 6.2.1 (1), showing a breather that splits, during evolution in $t$, into two breathers. Here we chose $\alpha_{1}=1$, $\alpha_{2}=2$, $A_{1}=\sqrt{6/5}$, $A_{2}=\frac{1}{2}\sqrt{41/10}$, $\gamma=1+\frac{1}{2}\mathrm{i}$, $r_{1}=1-\mathrm{i}$, $r_{2,3}=\frac{1}{10}(-3+19\,\mathrm{i}\pm\sqrt{-32-114\,\mathrm{i}})$, $c_{1}=c_{2}=1$, $c_{3}=10$.
Let $\mathrm{Re}(\gamma)=0$. We write $\gamma=-\mathrm{i}\,k$ with real $k$. The Lyapunov equation then becomes the constraint
$(A^{\dagger}\boldsymbol{\alpha}A)^{2}\,\big(|\chi_{1}|^{2}+|\chi_{2}|^{2}\big)=k\,|\Xi|^{2}\,$,
which requires $k>0$. Since this has to hold for all $x$ and $t$, it further requires that only one of the $c_{\kappa}$ is nonzero. Hence
$\chi_{1}$ $=$ $e^{rx+st}\,c\,$,
$\chi_{2}$ $=$ $-\frac{k\,\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A}{(\alpha_{2}-\alpha_{1})A_{1}A_{2}\big((1-A^{\dagger}\boldsymbol{\alpha}A)\,A^{\dagger}\boldsymbol{\alpha}A+k\,\alpha_{1}\alpha_{2}\|A\|^{2}\big)}\Big(r^{2}-\frac{\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}\,r+\frac{1}{4}k^{-2}$
$-\frac{1}{2\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)\,k^{-1}\Big)\,e^{rx+st}\,c\,$,
where $r$ is any root of the characteristic equation (6.62) (and $s$ the corresponding $s_{\kappa}$). From (6.62) and its complex conjugate, we get
$|A_{1}|^{2}=\Big(\big(1+2k\,(\alpha_{1}-\mathrm{Im}(r))\big)^{2}+4k^{2}\,\mathrm{Re}(r)^{2}\Big)\,\frac{2\,\mathrm{Im}(r)-\alpha_{2}}{4\,k\,\alpha_{1}^{2}\,(\alpha_{1}-\alpha_{2})}\,$,
$|A_{2}|^{2}=-\Big(\big(1+2k\,(\alpha_{2}-\mathrm{Im}(r))\big)^{2}+4k^{2}\,\mathrm{Re}(r)^{2}\Big)\,\frac{2\,\mathrm{Im}(r)-\alpha_{1}}{4\,k\,\alpha_{2}^{2}\,(\alpha_{1}-\alpha_{2})}\,$,
which requires
$\frac{2\,\mathrm{Im}(r)-\alpha_{2}}{\alpha_{1}-\alpha_{2}}>0\,,\qquad\frac{2\,\mathrm{Im}(r)-\alpha_{1}}{\alpha_{1}-\alpha_{2}}<0\,.$
By using these expressions for $|A_{1}|^{2}$ and $|A_{2}|^{2}$, a lengthy calculation reveals that the above constraint is automatically satisfied. Furthermore, we have
$\Omega_{x}$ $=$ $-\frac{|c|^{2}\big(\delta_{1}-\mathrm{i}\,k\,\mathrm{Re}(r)\big)}{k^{2}}\,\Big(1-\frac{\big(4\,\delta_{2}\,\mathrm{Im}(r)+\alpha_{1}\alpha_{2}\big)\,\big(\delta_{1}^{2}+k^{2}\,\mathrm{Re}(r)^{2}\big)}{(2\,\delta_{1}\delta_{2}+{k\,\alpha_{1}\alpha_{2}})^{2}+4\,k^{2}\,\delta^{2}_{2}\,\mathrm{Re}(r)^{2}}\Big)\,e^{2\theta}\,$,
$\Omega_{t}$ $=$ $\frac{2\,k\,\delta_{2}}{\alpha_{1}\alpha_{2}}\,\Omega_{x}\,$,
where
$\theta=\mathrm{Re}(r)\,\big(x+\frac{2\,k\,\delta_{2}}{\alpha_{1}\alpha_{2}}\,t\big)\,,\qquad\delta_{1}=k\,\mathrm{Im}(r)-\frac{1}{2}\,,\qquad\delta_{2}=\mathrm{Im}(r)-\frac{1}{2}(\alpha_{1}+\alpha_{2})\,.$
(i) $\mathrm{Re}(r)=0$. Then $\theta=0$ and we obtain
[…]
with a constant $\tilde{c}$, for which (6.22) requires
$\mathrm{Im}(\tilde{c})=-\frac{k}{2\delta_{1}}=\frac{k}{1-2\,k\,\mathrm{Im}(r)}\,$,
assuming $\delta_{1}\neq 0$. Since $\mathrm{Im}(\tilde{c})\neq 0$, $\Omega$ vanishes nowhere, so that
$u_{\mu}^{\prime}=A_{\mu}\,e^{\mathrm{i}\,\varphi_{\mu}}\,\Big(1+\frac{\mathrm{i}\,|c|^{2}}{(2\,\delta_{1}\delta_{2}+k\,\alpha_{1}\alpha_{2})\,\Omega}\Big[(-1)^{\mu}\,\frac{\delta_{1}(k\,\alpha_{1}-\delta_{1})\,(k\,\alpha_{2}-\delta_{1})\,(4\,\delta_{2}\,\mathrm{Im}(r)+\alpha_{1}\alpha_{2})}{k\,|A_{\mu}|^{2}(\alpha_{1}-\alpha_{2})\,(2\,\delta_{1}\delta_{2}+k\,\alpha_{1}\alpha_{2})}-\alpha_{1}\alpha_{2}\Big]\Big)\,$,
$\mu=1,2$, is regular on $\mathbb{R}^{2}$. Along the line $x+2\,k\,\delta_{2}\,t/(\alpha_{1}\alpha_{2})+\mathrm{Re}(\tilde{c})=0$, the absolute square of $u^{\prime}_{\mu}$ is given by
$|u^{\prime}_{\mu}|^{2}=|A_{\mu}|^{2}\,\big(1-b_{\mu}\big)^{2}\,,\qquad\mu=1,2,$
where
$b_{\mu}$ $=$ $\frac{2}{(2\,\delta_{1}\delta_{2}+k\,\alpha_{1}\,\alpha_{2})^{2}-\delta_{1}^{2}\,(4\,\delta_{2}\,\mathrm{Im}(r)+\alpha_{1}\alpha_{2})}\Big(k\,\alpha_{1}\alpha_{2}\,(2\,\delta_{1}\delta_{2}+k\,\alpha_{1}\alpha_{2})$
$-(-1)^{\mu}\,\frac{\delta_{1}\,(k\,\alpha_{1}-\delta_{1})\,(k\,\alpha_{2}-\delta_{1})\,(4\,\delta_{2}\,\mathrm{Im}(r)+\alpha_{1}\alpha_{2})}{(\alpha_{1}-\alpha_{2})\,|A_{\mu}|^{2}}\Big)\,.$
If $0<b_{\mu}<2$, then $u^{\prime}_{\mu}$ represents a dark soliton. If $b_{\mu}<0$ or $b_{\mu}>2$, it represents an anti-dark soliton.
(ii) $\mathrm{Re}(r)\neq 0$. Then we have
$\Omega$ $=$ $-\frac{|c|^{2}\big(\delta_{1}-\mathrm{i}\,k\,\mathrm{Re}(r)\big)}{2\,k^{2}\,\mathrm{Re}(r)}\,\Big(1-\frac{\big(4\,\delta_{2}\,\mathrm{Im}(r)+\alpha_{1}\alpha_{2}\big)\,\big(\delta_{1}^{2}+k^{2}\,\mathrm{Re}(r)^{2}\big)}{(2\,\delta_{1}\delta_{2}+{k\,\alpha_{1}\alpha_{2}})^{2}+4\,k^{2}\,\delta^{2}_{2}\,\mathrm{Re}(r)^{2}}\Big)\,e^{2\theta}+\tilde{c}\,$,
with a constant $\tilde{c}$, which (6.22) requires to be real. The generated solution of the two-component vector FL equation is
$u_{\mu}^{\prime}$ $=$ $A_{\mu}\,e^{\mathrm{i}\,\varphi_{\mu}}\,\Big(1+\frac{|c|^{2}\,e^{2\theta}\,F_{\mu}}{k^{2}|A_{\mu}|^{2}(\alpha_{1}-\alpha_{2})A^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{2}\big(k\,\alpha_{1}\alpha_{2}\|A\|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}+A^{\dagger}\boldsymbol{\alpha}A\big)^{2}\,\Omega}\Big)\,$,
where
$F_{\mu}$ $=$ $\Big(k\,|A_{\mu}|^{2}(\alpha_{1}-\alpha_{2})\,\big(k\,\alpha_{1}\alpha_{2}\|A\|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}+A^{\dagger}\boldsymbol{\alpha}A\big)+(-1)^{\mu}\,r^{\ast}\,\|A\|^{2}\,k^{2}\big(\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}^{2}A$
$+r^{\ast}\,A^{\dagger}\boldsymbol{\alpha}A)-(-1)^{\mu}\,k\,\big((A^{\dagger}\boldsymbol{\alpha}A)^{3}-\frac{1}{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}+\frac{1}{2}(\alpha_{1}-\alpha_{2})^{2}|A_{1}|^{2}|A_{2}|^{2}\big)+\frac{1}{4}(-1)^{\mu}\,A^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{2}\Big)$
$\times\Big(\|A\|^{2}\,k^{2}\big(\mathrm{i}\,(r^{2}-\alpha_{1}\alpha_{2})A^{\dagger}\boldsymbol{\alpha}A+r\,(\alpha_{1}\alpha_{2}\,\|A\|^{2}+A^{\dagger}\boldsymbol{\alpha}^{2}A)\big)-k\,\big((\frac{1}{2}\mathrm{i}+r\,\|A\|^{2})(A^{\dagger}\boldsymbol{\alpha}A)^{2}$
$-rA^{\dagger}\boldsymbol{\alpha}A\|A\|^{2}+\frac{1}{2}\mathrm{i}\,(\alpha_{1}\alpha_{2}\,\|A\|^{4}+|A_{1}|^{2}|A_{2}|^{2}(\alpha_{1}-\alpha_{2})^{2})\big)+\frac{1}{2}\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A\|A\|^{2}(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2})\Big)\,.$
This class of solutions includes dark and bright solitons on the plane wave background. Fig. 11 shows plots of the absolute values of the two components of $u^{\prime}$ for an example from this class of solutions of the two-component vector FL equation.
Figure 11: Plots of the absolute values of the two components of a solution from the class in Section 6.2.1 (ii), showing a bright-dark soliton. Here we chose $\gamma=-\mathrm{i}$ (i.e., $k=1$), $r=-1+\mathrm{i}/3$, $\alpha_{1}=-\alpha_{2}=1$, $c=1$, $\tilde{c}=-1$.
6.2.2 $n=2$ and generic roots
Given two $n=1$ solutions from Section 6.2.1, with spectral parameters $\gamma_{2}\neq-\gamma_{1}^{\ast}$, a corresponding superposition, which also solves the vector FL equation, is easily computed. With the correspondingly generalized restriction, an arbitrary number of $n=1$ solutions can easily be superposed. If, however, the (relative part of the) spectrum condition for the Lyapunov equation does not hold, it is more involved to compute the superposition of $n=1$ solutions. We elaborate an $n=2$ case.
Example 6.13.
Let $\Gamma=\mathrm{diag}(\gamma,-\gamma^{\ast})$ with $\mathrm{Re}(\gamma)\neq 0$. The Lyapunov equation determines the diagonal entries of $\Omega$,
$\Omega_{11}$ $=$ $\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}\,\gamma^{\ast}|\Xi_{1}|^{2}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}+|\chi_{11}|^{2}+|\chi_{21}|^{2}\Big)\,$,
$\Omega_{22}$ $=$ $\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}\,\gamma|\Xi_{2}|^{2}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}-|\chi_{12}|^{2}-|\chi_{22}|^{2}\Big)\,$,
where
$\Xi=\|A\|\,\Big(\chi_{1x}-\big(\frac{1}{2}\Gamma^{-1}+\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}}\,I_{n}\big)\,\chi_{1}\Big)-\mathrm{i}\,(\alpha_{2}-\alpha_{1})\,\frac{A_{1}A_{2}}{\|A\|}\,\chi_{2}\,$,
and the constraints
$\Big(\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Xi\Xi^{\dagger}\,\Gamma^{\dagger}+\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger}\Big)_{ij}=0\qquad\mbox{for}\quad i\neq j\,.$
We observe that if $r$ is a solution of (6.62), for a given parameter $\gamma$, then $r^{\ast}$ is a solution of (6.62) with $\gamma$ replaced by $\gamma^{\ast}$. Assuming that the three roots of (6.62) are all different [^8] If there is a multiple root, we still obtain exact solutions by restricting the sums to a set of different roots., the linear system is completely solved by (6.48) and (6.49), where now $R_{\kappa}=\mathrm{diag}(r_{\kappa},-r_{\kappa}^{\ast})$, $S_{\kappa}=\mathrm{diag}(s_{\kappa},-s_{\kappa}^{\ast})$, with $s_{\kappa}$ given by (6.63), and $c_{\kappa}=(c_{1\kappa},c_{2\kappa})^{T}$. Then the above constraints boil down to
$\sum_{\kappa=1}^{3}c_{1\kappa}\,c_{2\kappa}^{\ast}\,\big(1+d_{\kappa}^{2}+\mathrm{i}\,\gamma h_{\kappa}^{2}\big)=0\,$,
where
$d_{\kappa}$ $=$ $\frac{1}{(\alpha_{2}-\alpha_{1})A_{1}A_{2}\big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,A^{\dagger}\boldsymbol{\alpha}A-\gamma\,\alpha_{1}\alpha_{2}\|A\|^{2}\big)}\Big(\gamma\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A\,r_{\kappa}^{2}-\mathrm{i}\,\gamma\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}^{2}A\,r_{\kappa}$
$-\frac{1}{4}\gamma^{-1}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A+\frac{1}{2}\mathrm{i}\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}(2A^{\dagger}\boldsymbol{\alpha}A-1)+\frac{1}{2}\mathrm{i}\,(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}\Big)\,$,
$h_{\kappa}$ $=$ $\frac{1}{2\|A\|A^{\dagger}\boldsymbol{\alpha}A}\Big(2\,(\|A\|^{2}r_{\kappa}-\mathrm{i}\,k)-\gamma^{-1}\|A\|^{2}-2\mathrm{i}\,d_{\kappa}A_{1}A_{2}\,(\alpha_{2}-\alpha_{1})\Big)\,.$
The differential equations for $\Omega$ read
$\Omega_{12x}$ $=$ $\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Xi_{1x}\Xi_{2}^{\ast}-\frac{1}{2}\gamma^{-2}\chi_{11}\chi_{12}^{\ast}+\gamma^{-1}\Big(\frac{A_{1}^{\ast}A_{2}^{\ast}}{A^{\dagger}\boldsymbol{\alpha}A}(\alpha_{2}-\alpha_{1})\,(\chi_{11x}-\frac{1}{2}\gamma^{-1}\chi_{11})-\chi_{21x}$
$+\frac{\mathrm{i}\,\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\alpha_{1}\alpha_{2}\chi_{21}\Big)\,\chi_{22}^{\ast}\,$,
$\Omega_{12t}$ $=$ $\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\big(\gamma\,\Xi_{2}^{\ast}-\|A\|A^{\dagger}\boldsymbol{\alpha}A\,\chi_{12}^{\ast}\big)\,\Xi_{1}\,$,
$\Omega_{21x}$ $=$ $\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Xi_{1}^{\ast}\Xi_{2x}-\frac{1}{2}\gamma^{\ast-2}\chi_{11}^{\ast}\chi_{12}-\gamma^{\ast-1}\Big(\frac{A_{1}^{\ast}A_{2}^{\ast}}{A^{\dagger}\boldsymbol{\alpha}A}(\alpha_{2}-\alpha_{1})\,(\chi_{12x}+\frac{1}{2}\gamma^{\ast-1}\chi_{12})-\chi_{22x}$
$+\frac{\mathrm{i}\,\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\alpha_{1}\alpha_{2}\chi_{22}\Big)\,\chi_{21}^{\ast}\,$,
$\Omega_{21t}$ $=$ $-\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\big(\gamma^{\ast}\Xi_{1}^{\ast}+\|A\|A^{\dagger}\boldsymbol{\alpha}A\chi_{11}^{\ast}\big)\,\Xi_{2}\,$,
which leads to
$\Omega_{12}$ $=$ $-\frac{\mathrm{i}}{s_{1}-s_{2}}\big(c_{11}c_{22}^{\ast}h_{1}(\gamma h_{2}+\|A\|)\,e^{(r_{1}-r_{2})x+(s_{1}-s_{2})t}-c_{12}c_{21}^{\ast}h_{2}(\gamma h_{1}+\|A\|)\,e^{-(r_{1}-r_{2})x-(s_{1}-s_{2})t}\big)$
$-\frac{\mathrm{i}}{s_{1}-s_{3}}\big(c_{11}c_{23}^{\ast}h_{1}(\gamma h_{3}+\|A\|)\,e^{(r_{1}-r_{3})x+(s_{1}-s_{3})t}-c_{13}c_{21}^{\ast}h_{3}(\gamma h_{1}+\|A\|)\,e^{-(r_{1}-r_{3})x-(s_{1}-s_{3})t}\big)$
$-\frac{\mathrm{i}}{s_{2}-s_{3}}\big(c_{12}c_{23}^{\ast}h_{2}(\gamma h_{3}+\|A\|)\,e^{(r_{2}-r_{3})x+(s_{2}-s_{3})t}-c_{13}c_{22}^{\ast}h_{3}(\gamma h_{2}+\|A\|)\,e^{-(r_{2}-r_{3})x-(s_{2}-s_{3})t}\big)$
$-\mathrm{i}\,\sum_{\kappa=1}^{3}c_{1\kappa}c_{2\kappa}^{\ast}h_{\kappa}(\gamma h_{\kappa}+\|A\|)\,t+\frac{1}{2\gamma^{2}A^{\dagger}\boldsymbol{\alpha}A}\Big((\alpha_{2}-\alpha_{1})A_{1}^{\ast}A_{2}^{\ast}\sum_{\kappa=1}^{3}c_{1\kappa}c_{2\kappa}^{\ast}d_{\kappa}(2\gamma r_{\kappa}-1)$
$-\sum_{\kappa=1}^{3}\big(A^{\dagger}\boldsymbol{\alpha}A\,c_{1\kappa}c_{2\kappa}^{\ast}(2\mathrm{i}\,\gamma^{2}r_{\kappa}h_{\kappa}^{2}+2\gamma r_{\kappa}d_{\kappa}^{2}+1)-2\mathrm{i}\,\gamma\alpha_{1}\alpha_{2}\|A\|^{2}c_{1\kappa}c_{2\kappa}^{\ast}d_{\kappa}^{2}\big)\Big)\,x+\tilde{c}_{12}\,$,
$\Omega_{21}$ $=$ $\frac{\mathrm{i}}{s_{1}^{\ast}-s_{2}^{\ast}}\big(c_{11}^{\ast}c_{22}h_{2}^{\ast}(\gamma^{\ast}h_{1}^{\ast}+\|A\|)\,e^{(r_{1}^{\ast}-r_{2}^{\ast})x+(s_{1}^{\ast}-s_{2}^{\ast})t}-c_{12}^{\ast}c_{21}h_{1}^{\ast}(\gamma^{\ast}h_{2}^{\ast}+\|A\|)\,e^{-(r_{1}^{\ast}-r_{2}^{\ast})x-(s_{1}^{\ast}-s_{2}^{\ast})t}\big)$
$+\frac{\mathrm{i}}{s_{1}^{\ast}-s_{3}^{\ast}}\big(c_{11}^{\ast}c_{23}h_{3}^{\ast}(\gamma^{\ast}h_{1}^{\ast}+\|A\|)e^{(r_{1}^{\ast}-r_{3}^{\ast})x+(s_{1}^{\ast}-s_{3}^{\ast})t}-c_{13}^{\ast}c_{21}h_{1}^{\ast}(\gamma^{\ast}h_{3}^{\ast}+\|A\|)\,e^{-(r_{1}^{\ast}-r_{3}^{\ast})x-(s_{1}^{\ast}-s_{3}^{\ast})t}\big)$
$+\frac{\mathrm{i}}{s_{2}^{\ast}-s_{3}^{\ast}}\big(c_{12}^{\ast}c_{23}h_{3}^{\ast}(\gamma^{\ast}h_{2}^{\ast}+\|A\|)\,e^{(r_{2}^{\ast}-r_{3}^{\ast})x+(s_{2}^{\ast}-s_{3}^{\ast})t}-c_{13}^{\ast}c_{22}h_{2}^{\ast}(\gamma^{\ast}h_{3}^{\ast}+\|A\|)\,e^{-(r_{2}^{\ast}-r_{3}^{\ast})x-(s_{2}^{\ast}-s_{3}^{\ast})t}\big)$
$+\mathrm{i}\,\sum_{\kappa=1}^{3}c_{1\kappa}^{\ast}c_{2\kappa}h_{\kappa}^{\ast}(\gamma^{\ast}h_{\kappa}^{\ast}+\|A\|)\,t+\frac{1}{2\gamma^{\ast 2}A^{\dagger}\boldsymbol{\alpha}A}\Big((\alpha_{2}-\alpha_{1})A_{1}^{\ast}A_{2}^{\ast}\sum_{\kappa=1}^{3}c_{1\kappa}^{\ast}c_{2\kappa}d_{\kappa}^{\ast}(2\gamma^{\ast}r_{\kappa}^{\ast}-1)$
$+\sum_{\kappa=1}^{3}\big(A^{\dagger}\boldsymbol{\alpha}A\,c_{1\kappa}^{\ast}c_{2\kappa}(2\mathrm{i}\,\gamma^{\ast 2}r_{\kappa}^{\ast}h_{\kappa}^{\ast 2}-2\gamma^{\ast}r_{\kappa}^{\ast}d_{\kappa}^{\ast 2}-1)-2\mathrm{i}\,\gamma^{\ast}\alpha_{1}\alpha_{2}\|A\|^{2}c_{1\kappa}^{\ast}c_{2\kappa}d_{\kappa}^{\ast 2}\big)\Big)\,x+\tilde{c}_{21}\,$,
with complex constants $\tilde{c}_{12}$ and $\tilde{c}_{21}$. The remaining condition (6.22) becomes
$\gamma\,(\tilde{c}_{12}-\tilde{c}_{21}^{\ast})=\sum_{\kappa=1}^{3}\big(c_{1\kappa}\,c_{2\kappa}^{\ast}+d_{1\kappa}\,d_{2\kappa}^{\ast}\big)\,.$
A class of solutions of the two-component FL equation is now obtained by inserting the expressions for $\chi_{1}$, $\chi_{2}$, $\Xi$ and $\Omega$ in (6.21). In addition to the exponential dependence on the independent variables via $e^{\pm(r_{i}-r_{j})x\pm(s_{i}-s_{j})t}$ and their complex conjugates, the solution also exhibits a rational dependence, a typical feature in those cases, where the spectrum condition for the Lyapunov equation does not hold. $\Box$
Example 6.14.
Let $\Gamma$ be the $2\times 2$ lower triangular Jordan block with eigenvalue $\gamma$. Then we have
$R_{\kappa}=\left(\begin{array}[]{cc}r_{\kappa}&0\\ \tilde{r}_{\kappa}&r_{\kappa}\end{array}\right)\,,\qquad S_{\kappa}=\left(\begin{array}[]{cc}s_{\kappa}&0\\ \tilde{s}_{\kappa}&s_{\kappa}\end{array}\right)\,$,
with
$\tilde{r}_{\kappa}$ $=$ $\frac{8\mathrm{i}\,\gamma A^{\dagger}\boldsymbol{\alpha}^{2}A\,(\gamma r_{\kappa}-1)+4\gamma^{2}\alpha_{1}\alpha_{2}\,(2A^{\dagger}\boldsymbol{\alpha}A-1)+4\mathrm{i}\,\gamma\,(\alpha_{1}+\alpha_{2})-4\gamma r_{\kappa}\,(1+\gamma r_{\kappa})+3}{2\gamma^{2}\big(4\mathrm{i}\,\gamma A^{\dagger}\boldsymbol{\alpha}^{2}A-8\mathrm{i}\,\gamma^{2}r_{\kappa}\,(\alpha_{1}+\alpha_{2})-4\gamma^{2}\alpha_{1}\alpha_{2}+12\gamma^{2}r_{\kappa}^{2}-4\gamma r_{\kappa}-1\big)}\,$,
$\tilde{s}_{\kappa}$ $=$ $-\frac{1}{4\alpha_{1}\alpha_{2}\gamma^{2}}\Big(4\mathrm{i}\,\gamma^{2}(r_{\kappa}+\gamma\tilde{r}_{\kappa})\,(\alpha_{1}+\alpha_{2})+2\gamma^{2}\alpha_{1}\alpha_{2}-8\gamma^{3}r_{\kappa}\tilde{r}_{\kappa}-4\gamma^{2}r_{\kappa}^{2}-1\Big)\,.$
Assuming that the three roots $r_{\kappa}$, $\kappa=1,2,3$, of (6.62) are all different, the linear system is completely solved by (6.48) and (6.49). Then $\Xi$ is easily computed and it remains to determine $\Omega$, in order to be able to elaborate the resulting solution of the two-component vector FL equation, which is a second order version of that in Section 6.2.1 with the same $\gamma$. If $\mathrm{Re}(\gamma)\neq 0$, then $\Omega$ is fully determined by the Lyapunov equation, see Example A.2. If $\mathrm{Re}(\gamma)=0$, the differential equations for $\Omega$ have to be solved, which is more tedious. We omit the details. $\Box$
6.2.3 The triple root case
The solution of the linear equation for $\chi_{1}$, obtained in Section 6.2 in terms of only exponential functions, is not complete in case of a multiple root of the characteristic equation (6.62), because there are then additional solutions involving polynomials of the independent variables. In this section, we address the special case of a triple root (also see [11, 16]). This leads to genuine [^9] In contrast to solutions that are equivalent to solutions of the scalar FL equation. quasi-rational solutions of the two-component vector FL equation, which turn out to describe in space and time (formally regarding $t$ as time and $x$ as space coordinate) localized perturbations of the plane wave background.
Example 6.15.
Let $n=1$. If (6.62) admits a triple root, then it is given by
$r=\frac{\mathrm{i}}{3}(\alpha_{1}+\alpha_{2})+\frac{1}{6}\gamma^{-1}\,.$ (6.68)
The triple root condition further requires
$\gamma$ $=$ $\frac{2\,\big((\alpha_{1}-\alpha_{2})^{2}+\alpha_{1}\alpha_{2}\big)}{\mathrm{i}\,(\alpha_{1}+\alpha_{2})(2\alpha_{1}-\alpha_{2})(\alpha_{1}-2\alpha_{2})\pm 3\sqrt{3}\,\alpha_{1}\alpha_{2}(\alpha_{1}-\alpha_{2})}\,$, (6.69)
$|A_{\mu}|^{2}$ $=$ $\frac{\alpha_{\mu}\,(\alpha_{1}-\alpha_{2})^{2}}{\big((\alpha_{1}-\alpha_{2})^{2}+\alpha_{1}\alpha_{2}\big)^{2}}\,,\qquad\mu=1,2\,.$ (6.70)
Introducing
$\tilde{\chi}_{1}=\chi_{1}\,e^{-\theta}\,$,
with
$\theta=r\,x+\frac{1}{\alpha_{1}\alpha_{2}}\Big(r^{2}\gamma-\mathrm{i}\,r\gamma(\alpha_{1}+\alpha_{2})-\big(\frac{1}{2}\alpha_{1}\alpha_{2}\gamma+\frac{1}{4}\gamma^{-1}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2})\big)\Big)\,t\,$, (6.71)
the linear equations for $\chi_{1}$ boil down to
$\tilde{\chi}_{1xxx}=0\,,\qquad\alpha_{1}\alpha_{2}\,\tilde{\chi}_{1t}=\gamma\,\tilde{\chi}_{1xx}+\frac{1}{3}\tilde{\gamma}\,\tilde{\chi}_{1x}\,$,
where
$\tilde{\gamma}=1-\mathrm{i}\,(\alpha_{1}+\alpha_{2})\gamma\,.$ (6.72)
The linear system is now completely solved by
$\chi_{1}$ $=$ $\big(f_{1}\,x^{2}+f_{2}\,x+f_{3}\big)\,e^{\theta}\,$,
$\chi_{2}$ $=$ $\frac{\|A\|^{2}}{(\alpha_{2}-\alpha_{1})\,A_{1}A_{2}}\Big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,\gamma^{-1}-\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\Big)^{-1}\Big(\chi_{1xx}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}\,\chi_{1x}-\frac{1}{4}\gamma^{-2}\chi_{1}$
$+\frac{\mathrm{i}}{2\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)\,\gamma^{-1}\chi_{1}\Big)$
$=$ $\frac{\|A\|^{2}}{(\alpha_{2}-\alpha_{1})\,A_{1}A_{2}}\Big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,\gamma^{-1}-\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\Big)^{-1}\Big(\Big[\frac{\mathrm{i}}{2\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}$
$-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)\,\gamma^{-1}+r^{2}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}r-\frac{1}{4}\gamma^{-2}\Big]\,(f_{1}\,x^{2}+f_{2}\,x+f_{3})$
$-\big(\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}-2r)\,(2f_{1}\,x+f_{2})+2f_{1}\Big)\,e^{\theta}\,$,
where the functions $f_{1},f_{2},f_{3}$ satisfy
$\left(\begin{array}[]{c}f_{1}\\ f_{2}\\ f_{3}\end{array}\right)_{t}=C\left(\begin{array}[]{c}f_{1}\\ f_{2}\\ f_{3}\end{array}\right)\,,\qquad C=\frac{1}{3\,\alpha_{1}\alpha_{2}}\,\left(\begin{array}[]{ccc}0&0&0\\ 2\tilde{\gamma}&0&0\\ 6\gamma&\tilde{\gamma}&0\end{array}\right)\,.$
Hence
$\left(\begin{array}[]{c}f_{1}\\ f_{2}\\ f_{3}\end{array}\right)=e^{Ct}\left(\begin{array}[]{c}c_{1}\\ c_{2}\\ c_{3}\end{array}\right)=\left(\begin{array}[]{ccc}1&0&0\\ 2\tilde{\gamma}t/(3\,\alpha_{1}\alpha_{2})&1&0\\ \tilde{\gamma}^{2}t^{2}/(3\,\alpha_{1}\alpha_{2})^{2}+2\gamma t/(\alpha_{1}\alpha_{2})&\tilde{\gamma}t/(3\,\alpha_{1}\alpha_{2})&1\end{array}\right)\left(\begin{array}[]{c}c_{1}\\ c_{2}\\ c_{3}\end{array}\right)\,$, (6.86)
with complex constants $c_{1},c_{2},c_{3}$. Since we assume $\alpha_{1}\neq\alpha_{2}$, we have
$\mathrm{Re}(\gamma)=\pm\frac{3\sqrt{3}\,\alpha_{1}\alpha_{2}\,(\alpha_{1}-\alpha_{2})}{2\,\big((\alpha_{1}-\alpha_{2})^{2}+\alpha_{1}\alpha_{2}\big)^{2}}\neq 0\,$,
so that the Lyapunov equation has the unique solution
$\Omega=\frac{\mathrm{i}\,\gamma^{\ast}|\Xi|^{2}+(A^{\dagger}\boldsymbol{\alpha}A)^{2}(|\chi_{1}|^{2}+|\chi_{2}|^{2})}{2\mathrm{Re}(\gamma)\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\,.$
The generated solution (6.66) of the two-component vector FL equation is now
$u_{1}^{\prime}$ $=$ $e^{\mathrm{i}\,\varphi_{1}}\Big(A_{1}+\frac{2\mathrm{Re}(\gamma)A^{\dagger}\boldsymbol{\alpha}A\,(A_{1}\chi_{1}^{\ast}-A_{2}^{\ast}\chi_{2}^{\ast})\,\Xi}{\|A\|\big(\mathrm{i}\,\gamma^{\ast}|\Xi|^{2}+(A^{\dagger}\boldsymbol{\alpha}A)^{2}(|\chi_{1}|^{2}+|\chi_{2}|^{2})\big)}\Big)\,$,
$u_{2}^{\prime}$ $=$ $e^{\mathrm{i}\,\varphi_{2}}\Big(A_{2}+\frac{2\mathrm{Re}(\gamma)A^{\dagger}\boldsymbol{\alpha}A\,(A_{2}\chi_{1}^{\ast}+A_{1}^{\ast}\chi_{2}^{\ast})\,\Xi}{\|A\|\big(\mathrm{i}\,\gamma^{\ast}|\Xi|^{2}+(A^{\dagger}\boldsymbol{\alpha}A)^{2}(|\chi_{1}|^{2}+|\chi_{2}|^{2})\big)}\Big)\,$,
where
$\Xi$ $=$ $\|A\|\,\Big(\big(r-(\frac{1}{2}\gamma^{-1}+\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}})\big)\,(f_{1}x^{2}+f_{2}x+f_{3})+2f_{1}x+f_{2}$
$-\Big((1-A^{\dagger}\boldsymbol{\alpha}A)\,\gamma^{-1}+\mathrm{i}\,\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\Big)^{-1}\Big[\Big(\frac{\mathrm{i}}{2\gamma\,\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}$
$-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)+r^{2}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}r-\frac{1}{4}\gamma^{-2}\Big)(f_{1}x^{2}+f_{2}x+f_{3})$
$-\big(\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}-2r)\,(2f_{1}x+f_{2})+2f_{1}\Big]\Big)\,e^{\theta}\,.$
We note that $e^{\theta}$ drops out in the expression for $u^{\prime}$, so that the solution $u^{\prime}$ is quasi-rational.
If $c_{1}=c_{2}=0$, then $u_{1}^{\prime}$ and $u_{2}^{\prime}$ reduce to the plane wave background solutions.
If $c_{1}=0$ and $c_{2}\neq 0$, so that $f_{1}=0$ and $f_{2}=c_{2}\neq 0$, then $\chi_{1}$, $\chi_{2}$ and $\Xi$ only involve polynomials (in the independent variables) that are linear. In this case we obtain the single Peregrine-type soliton solution
$u_{\mu}^{\prime}=\frac{1}{2}A_{\mu}\,e^{\mathrm{i}\,\varphi_{\mu}}\Big(-1\pm(-1)^{\mu}\,\mathrm{i}\,\sqrt{3}-\frac{4\,\mathrm{Re}(\gamma)A^{\dagger}\boldsymbol{\alpha}A\,F_{\mu}}{D\,A_{\mu}}\Big)\qquad\quad\mu=1,2\,$,
where
$F_{\mu}$ $=$ $\varrho_{3}\,\Big(A_{\mu}\big(\mathrm{i}\,\gamma^{\ast}\varrho_{2}^{\ast}\varrho_{3}\|A\|^{2}+\varrho_{1}^{\ast}\varrho_{4}(A^{\dagger}\boldsymbol{\alpha}A)^{2}\big)+\tilde{A}_{\mu}\big(\mathrm{i}\,\gamma^{\ast}\varrho_{2}^{\ast}\varrho_{3}\varrho_{4}^{\ast}\|A\|^{2}-\varrho_{1}^{\ast}(\mathrm{i}\,\gamma^{\ast}|\varrho_{3}|^{2}\|A\|^{2}+(A^{\dagger}\boldsymbol{\alpha}A)^{2})\big)\Big)\,\rho$
$+(A^{\dagger}\boldsymbol{\alpha}A)^{2}(A_{\mu}+\varrho_{4}^{\ast}\tilde{A}_{\mu})\,(\varrho_{1}\varrho_{3}\varrho_{4}^{\ast}-\varrho_{2}|\varrho_{4}|^{2}-\varrho_{2})\,\rho^{\ast}+\varrho_{3}A_{\mu}\big(\mathrm{i}\,\gamma^{\ast}|\varrho_{2}|^{2}\|A\|^{2}+|\varrho_{1}|^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}\big)$
$-\tilde{A}_{\mu}\big(\varrho_{1}^{\ast}\varrho_{2}(\mathrm{i}\,\gamma^{\ast}|\varrho_{3}|^{2}\|A\|^{2}+(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1+|\varrho_{4}|^{2}))-\varrho_{3}\varrho_{4}^{\ast}(\mathrm{i}\,\gamma^{\ast}|\varrho_{2}|^{2}\|A\|^{2}+|\varrho_{1}|^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2})\big)\,$,
$D$ $=$ $\big(\mathrm{i}\,\gamma^{\ast}|\varrho_{3}|^{2}\|A\|^{2}+(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1+|\varrho_{4}|^{2})\big)\,\Big(\big((A^{\dagger}\boldsymbol{\alpha}A)^{2}(1+|\varrho_{4}|^{2})+\mathrm{i}\,\gamma^{\ast}|\varrho_{3}|^{2}\|A\|^{2}\big)\,|\rho|^{2}$
$+\big((A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{1}^{\ast}\varrho_{4}+\mathrm{i}\,\gamma^{\ast}\varrho_{2}^{\ast}\varrho_{3}\|A\|^{2}\big)\,\rho+\big((A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{1}\varrho_{4}^{\ast}+\mathrm{i}\,\gamma^{\ast}\varrho_{2}\varrho_{3}^{\ast}\|A\|^{2}\big)\,\rho^{\ast}$
$+(A^{\dagger}\boldsymbol{\alpha}A)^{2}|\varrho_{1}|^{2}+\mathrm{i}\,\gamma^{\ast}|\varrho_{2}|^{2}\|A\|^{2}\Big)\,$,
with $\tilde{A}_{\mu}=(-1)^{\mu}A_{3-\mu}^{\ast}$, $\rho=c_{2}\,(x+(3\alpha_{1}\alpha_{2})^{-1}\tilde{\gamma}\,t)+c_{3}$, and
$\varrho_{1}$ $=\pm\frac{4\sqrt{3}\,c_{2}\,(\nu_{1}^{2}-\nu_{2}^{2})}{A_{1}A_{2}\nu_{2}^{2}\,(2\nu_{1}-\nu_{2})^{2}}\,$, $\qquad\varrho_{2}$ $=\frac{c_{2}\,(\nu_{1}^{2}-4\nu_{1}\nu_{2}+\nu_{2}^{2})}{(2\nu_{1}-\nu_{2})^{2}}\,$,
$\varrho_{3}$ $=\mp\frac{(\nu_{1}+\nu_{2})\,(\nu_{1}^{2}-4\nu_{1}\nu_{2}+\nu_{2}^{2})}{6\nu_{1}(2\nu_{1}-\nu_{2})}\,$, $\qquad\varrho_{4}$ $=\frac{2\sqrt{3}(\nu_{1}+\nu_{2})^{2}(\nu_{1}^{2}-\nu_{1}\nu_{2}+\nu_{2}^{2})}{9A_{1}A_{2}\nu_{1}^{2}\nu_{2}^{2}(2\nu_{1}-\nu_{2})}\,.$
Here $\nu_{1}=\sqrt{3}(\alpha_{1}-\alpha_{2})\pm\mathrm{i}\,(\alpha_{1}+\alpha_{2}),\,\nu_{2}=\sqrt{3}(\alpha_{1}-\alpha_{2})\mp\mathrm{i}\,(\alpha_{1}+\alpha_{2})$. The $\pm$ corresponds to that in the denominator of (6.69). This solution has been explored in particular in [10], where it has been shown that it exhibits surprisingly high amplitudes for particular parameter values, in particular as compared with the Manakov (two-component vector NLS) system (reached in the limit $\epsilon\to 0$, see Remark 6.4 for this parameter). Indeed, choosing for example $c_{2}=c_{3}=1$, we find that the absolute values of the components of $u^{\prime}$ at the point
$(x_{0},t_{0})=\Big(-1\mp\frac{\sqrt{3}\,\big((\alpha_{1}-\alpha_{2})^{4}-\alpha_{1}^{2}\alpha_{2}^{2}\big)}{6\alpha_{1}\alpha_{2}(\alpha_{1}^{3}-\alpha_{2}^{3})}\,,\mp\frac{\sqrt{3}\,\big(\alpha_{1}^{2}-\alpha_{1}\alpha_{2}+\alpha_{2}^{2}\big)^{3}}{6\alpha_{1}\alpha_{2}(\alpha_{1}^{3}-\alpha_{2}^{3})}\Big)$
are given by
$|u_{1}^{\prime}|=|A_{1}|\,\frac{|\alpha_{1}-5\alpha_{2}|}{|\alpha_{1}+\alpha_{2}|}\,,\qquad|u_{2}^{\prime}|=|A_{2}|\,\frac{|5\alpha_{1}-\alpha_{2}|}{|\alpha_{1}+\alpha_{2}|}\,.$
Note that (6.70) requires $\alpha_{\mu}>0$. The last equations show that the peak amplitude can approach nearly five times the background level, which is consistent with the results reported in [10]. An example of a rogue wave from this class of solutions is displayed in Fig. 12.
Figure 12: Plots of the absolute values of the two components of a solution from the class in Example 6.15, both showing rogue waves. Here we chose $\alpha_{1}=1$, $\alpha_{2}=2$, $A_{1}=1/3$, $A_{2}=\sqrt{2}/3$, $c_{1}=0$, $c_{2}=c_{3}=1$.
If $c_{1}\neq 0$, the asymptotically leading polynomial in $\chi_{1}$, $\chi_{2}$ and $\Xi$ is $x^{2}+\tilde{\gamma}^{2}(3\alpha_{1}\alpha_{2})^{-2}t^{2}$. Excluding $\mathrm{Re}(\tilde{\gamma})=0$, which is $(\alpha_{1}+\alpha_{2})\mathrm{Im}(\gamma)=-1$, this polynomial is different from zero for all $x,t$. Then we obtain
$u^{\prime}_{\mu}\sim\frac{1}{2}A_{\mu}\,e^{\mathrm{i}\,\varphi_{\mu}}\Big(-1\pm(-1)^{\mu}\,\mathrm{i}\,\sqrt{3}\Big)\qquad\quad\mu=1,2\,$,
as $x\to\infty$ or $t\to\infty$. Here the $\pm$ corresponds to that in the denominator of (6.69). As a consequence, asymptotically we have $|u^{\prime}_{\mu}|=|A_{\mu}|$. Hence the components of the solution $u^{\prime}$ are both localized on the respective plane wave background, which is the characteristic property of rogue waves (interpreting $x$ and $t$ as “space” and “time” coordinates, or vice versa). The components $u_{1}^{\prime}$ and $u_{2}^{\prime}$ form pairs of (elementary) rogue waves. An example is shown in Fig. 13.
Figure 13: Plots of the absolute squares of the two components of a solution from the class in Example 6.15, exhibiting rogue wave pairs. Here we chose the same data as in Fig. 12, but now $c_{1}=1$ and $c_{3}=0$.
$\Box$
Example 6.16.
Let $n=2$ and
$\Gamma=\left(\begin{array}[]{cc}\gamma&0\\ 1&\gamma\end{array}\right)\,,\qquad\chi_{i}=\left(\begin{array}[]{c}\chi_{i1}\\ \chi_{i2}\end{array}\right)\qquad i=1,2\,.$
Again, we assume that $r$ is a triple root of the cubic equation (6.62), hence given by (6.68). Then $\gamma$ and $A_{\mu}$, $\mu=1,2$, are given by (6.69) and (6.70), respectively. Now the linear system takes the form
$\tilde{\chi}_{11xxx}$ $=$ $0\,$,
$\tilde{\chi}_{12xxx}$ $=$ $-\frac{1}{2}\gamma^{-2}\big(\tilde{\chi}_{11xx}+2\gamma^{-1}(w_{r}\,\tilde{\chi}_{11x}+w\,\tilde{\chi}_{11})\big)\,$,
$\alpha_{1}\alpha_{2}\,\tilde{\chi}_{11t}$ $=$ $\gamma\,\tilde{\chi}_{11xx}+\frac{1}{3}\tilde{\gamma}\,\tilde{\chi}_{11x}\,$,
$\alpha_{1}\alpha_{2}\,\tilde{\chi}_{12t}$ $=$ $\gamma\,\tilde{\chi}_{12xx}+\frac{1}{3}\tilde{\gamma}\,\tilde{\chi}_{12x}+\gamma^{-1}\alpha_{1}\alpha_{2}\,\tilde{\chi}_{11t}+B\,\tilde{\chi}_{11}\,$,
where
$B$ $=$ $\frac{1}{18}\big(5\gamma^{-2}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})\gamma^{-1}+(4\alpha_{1}^{2}-\alpha_{1}\alpha_{2}+4\alpha_{2}^{2})\big)\,$,
$w$ $=$ $3\gamma^{2}r^{3}-\big(1+3\,\mathrm{i}\,\gamma(\alpha_{1}+\alpha_{2})\big)\gamma\,r^{2}+\big(2\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}^{2}A\gamma-3\alpha_{1}\alpha_{2}\gamma^{2}-\frac{1}{4}\big)r$
$-\frac{1}{2}\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}^{2}A+(2A^{\dagger}\boldsymbol{\alpha}A-1)\,\alpha_{1}\alpha_{2}\gamma+\frac{1}{4}\mathrm{i}\,(\alpha_{1}+\alpha_{2})\,.$
It is solved by
$\chi_{11}$ $=$ $\big(f_{1}\,x^{2}+f_{2}\,x+f_{3}\big)\,e^{\theta}\,$,
$\chi_{12}$ $=$ $\Big(f_{4}\,x^{2}+f_{5}\,x+f_{6}-\frac{(wf_{3}+w_{r}f_{2}+\gamma f_{1})\,x^{3}}{6\,\gamma^{3}}-\frac{(wf_{2}+2\,w_{r}f_{1})\,x^{4}}{24\,\gamma^{3}}-\frac{wf_{1}\,x^{5}}{60\,\gamma^{3}}\Big)\,e^{\theta}\,$,
$\chi_{2}$ $=$ $\frac{\|A\|^{2}}{(\alpha_{2}-\alpha_{1})\,A_{1}A_{2}}\Big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,\Gamma^{-1}-\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}I_{n}\Big)^{-1}\Big(\chi_{1xx}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}\,\chi_{1x}-\frac{1}{4}\Gamma^{-2}\chi_{1}$
$+\frac{\mathrm{i}}{2\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)\,\Gamma^{-1}\chi_{1}\Big)\,$,
where $\theta$ is given by (6.71) and $f=(f_{1},\ldots,f_{6})^{T}$ satisfies $f_{t}=Cf$ with the constant nilpotent matrix
$C=\frac{1}{3\,\alpha_{1}\alpha_{2}}\,\left(\begin{array}[]{cccccc}0&0&0&0&0&0\\ 2\tilde{\gamma}&0&0&0&0&0\\ 6\gamma&\tilde{\gamma}&0&0&0&0\\ 3(B-\gamma^{-2}w_{r})-\frac{1}{2}\gamma^{-2}\tilde{\gamma}&-\frac{1}{2}\gamma^{-3}(3w\gamma+\tilde{\gamma}w_{r})&-\frac{1}{2}w\gamma^{-3}\tilde{\gamma}&0&0&0\\ \gamma^{-1}(2\tilde{\gamma}-3)&3(B-\gamma^{-2}w_{r})&-3\gamma^{-2}w&2\tilde{\gamma}&0&0\\ 6&\gamma^{-1}\tilde{\gamma}&3B&6\gamma&\tilde{\gamma}&0\end{array}\right)\,.$
Consequently, $f_{1},f_{2},f_{3}$ are given by (6.86) and, in addition, we obtain
$f_{4}$ $=$ $-\frac{c_{1}w\tilde{\gamma}^{3}}{6\gamma^{3}(3\alpha_{1}\alpha_{2})^{3}}\,t^{3}-\frac{\tilde{\gamma}}{4\gamma^{3}(3\alpha_{1}\alpha_{2})^{2}}\,\big(2c_{1}\,(w_{r}\tilde{\gamma}+6w\gamma)+c_{2}w\tilde{\gamma}\big)\,t^{2}$
$+\frac{1}{6\gamma^{3}\alpha_{1}\alpha_{2}}\Big(c_{1}\gamma\,(6\,(B\gamma^{2}-w_{r})-\tilde{\gamma})-c_{2}\,(3w\gamma+w_{r}\tilde{\gamma})-c_{3}w\tilde{\gamma}\Big)\,t+c_{4}\,$,
$f_{5}$ $=$ $-\frac{c_{1}w\tilde{\gamma}^{4}}{12\gamma^{3}(3\alpha_{1}\alpha_{2})^{4}}\,t^{4}-\frac{\tilde{\gamma}^{2}}{6\gamma^{3}(3\alpha_{1}\alpha_{2})^{3}}\big(2c_{1}\,(w_{r}\tilde{\gamma}+9w\gamma)+c_{2}w\tilde{\gamma}\big)\,t^{3}$
$+\frac{1}{2\gamma^{3}(3\alpha_{1}\alpha_{2})^{2}}\Big(c_{1}\big(12\gamma\tilde{\gamma}\,(B\gamma^{2}-w_{r})-\gamma\,(18w\gamma+\tilde{\gamma}^{2})\big)-c_{2}\tilde{\gamma}\,(w_{r}\tilde{\gamma}+6w\gamma)-c_{3}w\tilde{\gamma}^{2}\Big)\,t^{2}$
$+\frac{1}{3\alpha_{1}\alpha_{2}\gamma^{2}}\Big(c_{1}\gamma\,(2\tilde{\gamma}-3)+3c_{2}\,(B\gamma^{2}-w_{r})-3c_{3}w+2c_{4}\gamma^{2}\tilde{\gamma}\Big)\,t+c_{5}\,$,
$f_{6}$ $=$ $-\frac{c_{1}w\tilde{\gamma}^{5}}{60\gamma^{3}(3\alpha_{1}\alpha_{2})^{5}}\,t^{5}-\frac{\tilde{\gamma}^{3}}{24\gamma^{3}(3\alpha_{1}\alpha_{2})^{4}}\big(2c_{1}\,(w_{r}\tilde{\gamma}+12w\gamma)+c_{2}w\tilde{\gamma}\big)\,t^{4}$
$+\frac{\tilde{\gamma}}{6\gamma^{3}(3\alpha_{1}\alpha_{2})^{3}}\Big(c_{1}\gamma\,\big(18\gamma\,(B\gamma\tilde{\gamma}-3w)-\tilde{\gamma}\,(18w_{r}+\tilde{\gamma})\big)-c_{2}\tilde{\gamma}(9w\gamma+w_{r}\tilde{\gamma})-c_{3}w\tilde{\gamma}^{2}\Big)\,t^{3}$
$+\frac{1}{2\gamma^{2}(3\alpha_{1}\alpha_{2})^{2}}\Big(2c_{1}\gamma\,\big(9\,(2B\gamma^{2}-w_{r})+\tilde{\gamma}\,(2\tilde{\gamma}-3)\big)+c_{2}\,\big(6\tilde{\gamma}\,(B\gamma^{2}-w_{r})-9w\gamma\big)-6c_{3}w\tilde{\gamma}$
$+2c_{4}\gamma^{2}\tilde{\gamma}^{2}\Big)\,t^{2}+\frac{1}{3\gamma\,\alpha_{1}\alpha_{2}}\big(6c_{1}\gamma+c_{2}\tilde{\gamma}+3c_{3}B\gamma+6c_{4}\gamma^{2}+c_{5}\gamma\tilde{\gamma}\big)\,t+c_{6}\,$,
with complex constants $c_{1},\ldots,c_{6}$, and $\tilde{\gamma}$ given by (6.72). The solution of the Lyapunov equation is
$\Omega=\left(\begin{array}[]{cc}\kappa^{-1}M_{11}&-\kappa^{-2}M_{11}+\kappa^{-1}M_{12}\\[4.0pt] -\kappa^{-2}M_{11}+\kappa^{-1}M_{21}&2\kappa^{-3}M_{11}-\kappa^{-2}(M_{12}+M_{21})+\kappa^{-1}M_{22}\end{array}\right)\,$,
with $\kappa=2\,\mathrm{Re}(\gamma)$ and
$M$ $=$ $(M_{ij})=\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Xi\,\Xi^{\dagger}\Gamma^{\dagger}+\chi\,\chi^{\dagger}$
$=$ $\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\left(\begin{array}[]{cc}\gamma^{\ast}|\Xi_{1}|^{2}&|\Xi_{1}|^{2}+\gamma^{\ast}\Xi_{1}\Xi_{2}^{\ast}\\[2.0pt] \gamma^{\ast}\Xi_{1}^{\ast}\Xi_{2}&\Xi_{1}^{\ast}\Xi_{2}+\gamma^{\ast}|\Xi_{2}|^{2}\end{array}\right)+\left(\begin{array}[]{cc}|\chi_{11}|^{2}+|\chi_{21}|^{2}&\chi_{11}\chi_{12}^{\ast}+\chi_{21}\chi_{22}^{\ast}\\[2.0pt] \chi_{11}^{\ast}\chi_{12}+\chi_{21}^{\ast}\chi_{22}&|\chi_{12}|^{2}+|\chi_{22}|^{2}\end{array}\right)\,$,
where
$\Xi=(\Xi_{1},\Xi_{2})^{T}=\|A\|\Big(\chi_{1x}-(\frac{1}{2}\Gamma^{-1}+\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}}I_{2})\,\chi_{1}-\mathrm{i}\,(\alpha_{2}-\alpha_{1})\frac{A_{1}A_{2}}{\|A\|^{2}}\,\chi_{2}\Big)\,.$
Hence
$\Omega^{-1}=\frac{1}{\det(\Omega)}\left(\begin{array}[]{cc}2\kappa^{-3}M_{11}-\kappa^{-2}(M_{12}+M_{21})+\kappa^{-1}M_{22}&\kappa^{-2}M_{11}-\kappa^{-1}M_{12}\\[4.0pt] \kappa^{-2}M_{11}-\kappa^{-1}M_{21}&\kappa^{-1}M_{11}\end{array}\right)\,$,
with
$\det(\Omega)=\kappa^{-4}M_{11}^{2}+\kappa^{-2}\det(M)\,.$
Inserting the expressions for $\chi_{1},\chi_{2}$ and $\Omega^{-1}$ in (6.61), determines a family of quasi-rational solutions of the two-component vector FL equation, which are localized on the plane wave background. These are thus rogue waves. They are of second order in the sense that here $\Gamma$ is a $2\times 2$ Jordan block, which causes the appearance of higher (than in the $n=1$ case) order polynomials of the dependent variables. $\Box$
Using an $n\times n$ Jordan block for $\Gamma$, and choosing $\gamma$, $A_{1}$, $A_{2}$ and $r$ as in the preceding examples, we obtain $n$-th order rogue waves (also see [33] for the NLS case). Results about rogue wave solutions of the two-component vector FL equation on the general plane wave background have been obtained earlier in [11]. A deeper analysis of the structure of rogue waves (location of peaks) in the triple root case appeared recently in [16].
A different class of rogue waves is obtained in the next example.
Example 6.17.
Let $\Gamma=\mathrm{diag}(\gamma,-\gamma^{\ast})$ with $\mathrm{Re}(\gamma)\neq 0$, and $r$ given by (6.68). Then we have $R=\mathrm{diag}(r,-r^{\ast})$. Using Example 6.15, the general solution of the linear system is then
$\chi_{1\kappa}$ $=$ $(f_{1\kappa}x^{2}+f_{2\kappa}x+f_{3\kappa})\,e^{\theta_{\kappa}}\,$,
$\chi_{2\kappa}$ $=$ $\frac{\|A\|^{2}}{(\alpha_{2}-\alpha_{1})\,A_{1}A_{2}}\Big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,\gamma_{\kappa}^{-1}-\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\Big)^{-1}\Big(\Big[\frac{\mathrm{i}}{2\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}$
$-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)\,\gamma_{\kappa}^{-1}+r_{\kappa}^{2}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}r_{\kappa}-\frac{1}{4}\gamma_{\kappa}^{-2}\Big]\,(f_{1\kappa}\,x^{2}+f_{2\kappa}\,x+f_{3\kappa})$
$-\big(\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}-2r_{\kappa})\,(2f_{1\kappa}\,x+f_{2\kappa})+2f_{1\kappa}\Big)\,e^{\theta_{\kappa}}\,$,
where
$\left(\begin{array}[]{c}f_{1\kappa}\\ f_{2\kappa}\\ f_{3\kappa}\end{array}\right)=\left(\begin{array}[]{ccc}1&0&0\\ 2\tilde{\gamma}_{\kappa}t/(3\,\alpha_{1}\alpha_{2})&1&0\\ \tilde{\gamma}_{\kappa}^{2}t^{2}/(3\,\alpha_{1}\alpha_{2})^{2}+2\gamma_{\kappa}t/(\alpha_{1}\alpha_{2})&\tilde{\gamma}_{\kappa}t/(3\,\alpha_{1}\alpha_{2})&1\end{array}\right)\left(\begin{array}[]{c}c_{1\kappa}\\ c_{2\kappa}\\ c_{3\kappa}\end{array}\right)\,$,
with $\kappa=1,2$, and $\gamma_{1}=\gamma$, $\gamma_{2}=-\gamma^{\ast}$, $r_{1}=r$, $r_{2}=-r^{\ast}$, $\theta_{1}=\theta$, $\theta_{2}=-\theta^{\ast}$. Here $r$ is again given by (6.68) and $\theta$ by (6.71). The Lyapunov equation determines the diagonal entries of $\Omega$,
$\Omega_{11}=\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}\,\gamma^{\ast}|\Xi_{1}|^{2}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}+|\chi_{11}|^{2}+|\chi_{21}|^{2}\Big)\,,\qquad\Omega_{22}=\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}\,\gamma|\Xi_{2}|^{2}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}-|\chi_{12}|^{2}-|\chi_{22}|^{2}\Big)\,$,
where
$\Xi_{\kappa}$ $=$ $\|A\|\,\Big(\big(r_{\kappa}-(\frac{1}{2}\gamma_{\kappa}^{-1}+\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}})\big)\,(f_{1\kappa}x^{2}+f_{2\kappa}x+f_{3\kappa})+2f_{1\kappa}x+f_{2\kappa}$
$-\Big((1-A^{\dagger}\boldsymbol{\alpha}A)\,\gamma_{\kappa}^{-1}+\mathrm{i}\,\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\Big)^{-1}\Big[\Big(\frac{\mathrm{i}}{2\gamma\,\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}$
$-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)+r_{\kappa}^{2}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}r_{\kappa}-\frac{1}{4}\gamma_{\kappa}^{-2}\Big)(f_{1\kappa}x^{2}+f_{2\kappa}x+f_{3\kappa})$
$-\big(\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}-2r_{\kappa})\,(2f_{1\kappa}x+f_{2\kappa})+2f_{1\kappa}\Big]\Big)\,e^{\theta_{\kappa}}\,$,
and the constraints
$\Big(\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Xi\Xi^{\dagger}\,\Gamma^{\dagger}+\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger}\Big)_{ij}=0\qquad\mbox{for}\quad i\neq j\,$,
which evaluate to
$c_{31}c_{32}^{\ast}+\frac{\|A\|^{4}D_{1}D_{2}}{(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}}+\frac{\mathrm{i}\,\gamma\,\|A\|^{2}\,H_{1}H_{2}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}=0\,.$
Here we introduced
$D_{1}=d(c_{11},c_{21},c_{31})\,,\qquad D_{2}=d(c_{12}^{\ast},-c_{22}^{\ast},c_{32}^{\ast})\,$,
$H_{1}=h(c_{11},c_{21},c_{31})\,,\qquad H_{2}=h(c_{12}^{\ast},-c_{22}^{\ast},c_{32}^{\ast})\,$,
and
$d(a,b,c)$ $=$ $\varrho_{3}\,\big(2a-\varrho_{1}b+\varrho_{4}c\big)\,$,
$h(a,b,c)$ $=$ $2\,\mathrm{i}\,\varrho_{3}a-(\mathrm{i}\,\varrho_{1}\varrho_{3}+1)\,b+(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})\,c\,$,
with constants
$\varrho_{1}$ $=$ $\frac{\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}-2r\,$,
$\varrho_{2}$ $=$ $r-\frac{1}{2}\gamma^{-1}-\frac{\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}}\,$,
$\varrho_{3}$ $=$ $\frac{\gamma A^{\dagger}\boldsymbol{\alpha}A}{\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A\,(1-A^{\dagger}\boldsymbol{\alpha}A)-\gamma\alpha_{1}\alpha_{2}\|A\|^{2}}\,$,
$\varrho_{4}$ $=$ $\frac{\mathrm{i}}{2\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)\,\gamma^{-1}+r^{2}-\frac{\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}r-\frac{1}{4}\gamma^{-2}\,.$
Integrating the differential equations for $\Omega$, we find
$\Omega_{12}$ $=$ $-\frac{\mathrm{i}\,c_{11}c_{12}^{\ast}\tilde{\gamma}^{4}\|A\|^{2}}{405\,\alpha_{1}^{4}\alpha_{2}^{4}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})\,\big(\gamma\,(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})-A^{\dagger}\boldsymbol{\alpha}A\big)\,t^{5}$
$+\frac{\mathrm{i}\,\tilde{\gamma}^{2}\|A\|}{108\,\alpha_{1}^{3}\alpha_{2}^{3}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Big((\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})\,\big(c_{12}^{\ast}(2c_{11}\|A\|A^{\dagger}\boldsymbol{\alpha}A\,(\tilde{\gamma}\,x-3\gamma)+3\gamma\alpha_{1}\alpha_{2}\varrho_{6})$
$+c_{11}(c_{22}^{\ast}\tilde{\gamma}\|A\|A^{\dagger}\boldsymbol{\alpha}A-3\gamma\alpha_{1}\alpha_{2}\varrho_{5})\big)-3c_{12}^{\ast}\alpha_{1}\alpha_{2}\varrho_{6}A^{\dagger}\boldsymbol{\alpha}A\Big)\,t^{4}$
$+\frac{\mathrm{i}}{27\,\alpha_{1}^{2}\alpha_{2}^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Big(c_{11}c_{12}^{\ast}\tilde{\gamma}^{2}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})\,x^{2}$
$+\tilde{\gamma}\|A\|A^{\dagger}\boldsymbol{\alpha}A\,\big(c_{11}c_{22}^{\ast}\tilde{\gamma}\|A\|\,(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})-6c_{12}^{\ast}\alpha_{1}\alpha_{2}\varrho_{6}\big)\,x$
$-\tilde{\gamma}^{2}\|A\|^{2}(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})\,\big(c_{12}^{\ast}\gamma\varrho_{7}-c_{11}(c_{32}^{\ast}A^{\dagger}\boldsymbol{\alpha}A-\gamma\,\varrho_{8})\big)+c_{12}^{\ast}\|A\|A^{\dagger}\boldsymbol{\alpha}A\,(\tilde{\gamma}^{2}\varrho_{7}\|A\|$
$+18\gamma\alpha_{1}\alpha_{2}\varrho_{6})-3\alpha_{1}\alpha_{2}\varrho_{6}(c_{22}^{\ast}\tilde{\gamma}\|A\|A^{\dagger}\boldsymbol{\alpha}A-3\gamma\alpha_{1}\alpha_{2}\varrho_{5})\Big)\,t^{3}$
$-\frac{\mathrm{i}\,\|A\|}{6\alpha_{1}\alpha_{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Big(3c_{12}^{\ast}\alpha_{1}\alpha_{2}\varrho_{6}A^{\dagger}\boldsymbol{\alpha}A\,x^{2}-A^{\dagger}\boldsymbol{\alpha}A\,(2c_{12}^{\ast}\tilde{\gamma}\varrho_{7}\|A\|-3c_{22}^{\ast}\alpha_{1}\alpha_{2}\varrho_{6})\,x$
$+\varrho_{7}\|A\|A^{\dagger}\boldsymbol{\alpha}A\,(6c_{12}^{\ast}\gamma-c_{22}^{\ast}\tilde{\gamma})+3\alpha_{1}\alpha_{2}\big(c_{32}^{\ast}\varrho_{6}A^{\dagger}\boldsymbol{\alpha}A+\gamma(\varrho_{5}\varrho_{7}-\varrho_{6}\varrho_{8})\big)\Big)\,t^{2}$
$+\frac{\mathrm{i}\,\varrho_{7}\|A\|^{2}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\big(A^{\dagger}\boldsymbol{\alpha}A\,(c_{12}^{\ast}x^{2}+c_{22}^{\ast}x+c_{32}^{\ast})-\gamma\varrho_{8}\big)\,t$
$+\frac{c_{11}c_{12}^{\ast}}{10\gamma^{2}(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A}\Big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}(2\gamma r-1)\,(\varrho_{3}\varrho_{4}\|A\|^{2}+A^{\dagger}\boldsymbol{\alpha}A)$
$+2\mathrm{i}\,\gamma\alpha_{1}\alpha_{2}\varrho_{3}^{2}\varrho_{4}^{2}\|A\|^{6}\Big)\,x^{5}$
$+\frac{1}{8\gamma^{2}(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A}\Big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}\big(2c_{11}c_{12}^{\ast}(|A|^{2}\varrho_{3}(2\gamma\,(r\varrho_{1}+\varrho_{4})-\varrho_{1})$
$+2\gamma A^{\dagger}\boldsymbol{\alpha}A)+(2\gamma r-1)\,(c_{11}c_{22}^{\ast}+c_{12}^{\ast}c_{21})\,(\varrho_{3}\varrho_{4}\|A\|^{2}+A^{\dagger}\boldsymbol{\alpha}A)\big)$
$+2\mathrm{i}\,\gamma\alpha_{1}\alpha_{2}\varrho_{3}^{2}\varrho_{4}^{2}\|A\|^{6}(c_{11}c_{22}^{\ast}+c_{12}^{\ast}c_{21})\Big)\,x^{4}$
$+\frac{1}{6\gamma^{2}(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Big($
$(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}\big(c_{12}^{\ast}\varrho_{10}+c_{22}^{\ast}((A^{\dagger}\boldsymbol{\alpha}A)^{2}(c_{21}(2\gamma r-1)+2c_{11}\gamma)$
$+\varrho_{3}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A\,(2\gamma r-1)\,(c_{11}\varrho_{1}+c_{21}\varrho_{4})+4c_{11}\gamma\varrho_{3}\varrho_{4}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A-2\mathrm{i}\,c_{11}\gamma^{2}\|A\|^{2}(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})^{2})$
$+c_{11}c_{32}^{\ast}A^{\dagger}\boldsymbol{\alpha}A\,(2\gamma r-1)(\varrho_{3}\varrho_{4}\|A\|^{2}+A^{\dagger}\boldsymbol{\alpha}A)\big)$
$+2\gamma\varrho_{3}^{2}\|A\|^{4}A^{\dagger}\boldsymbol{\alpha}A\,\big(\mathrm{i}\,\alpha_{1}\alpha_{2}\varrho_{9}\|A\|^{2}-c_{11}\varrho_{4}A^{\dagger}\boldsymbol{\alpha}A\,(2c_{12}^{\ast}\varrho_{1}+c_{22}^{\ast}\varrho_{4})\big)\Big)\,x^{3}$
$-\frac{1}{4\gamma^{2}(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}\big(4\mathrm{i}\,c_{11}\gamma^{2}H_{2}\|A\|^{2}(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})$
$-D_{2}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A\,(2\gamma\,(2c_{11}+c_{21}r)-c_{21})-2\gamma\varrho_{3}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A\,(c_{21}+c_{31}r)(2c_{12}^{\ast}\varrho_{1}+c_{22}^{\ast}\varrho_{4})$
$+c_{31}\varrho_{3}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A\,(2c_{12}^{\ast}\varrho_{1}+c_{22}^{\ast}\varrho_{4})\big)-\varrho^{(1)}+4c_{11}\gamma\varrho_{3}\varrho_{4}D_{2}\|A\|^{4}(A^{\dagger}\boldsymbol{\alpha}A)^{2}\Big)\,x^{2}$
$+\frac{1}{2\gamma^{2}(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}\big(c_{31}c_{32}^{\ast}(A^{\dagger}\boldsymbol{\alpha}A)^{2}(2\gamma r-1)$
$-\|A\|^{2}(2\gamma^{2}H_{2}(\mathrm{i}\,c_{21}(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})-2c_{11}(\mathrm{i}-\varrho_{1}\varrho_{3}))-2\gamma D_{2}A^{\dagger}\boldsymbol{\alpha}A\,(c_{21}+c_{31}r)+c_{31}D_{2}A^{\dagger}\boldsymbol{\alpha}A)\big)$
$+2\gamma D_{2}\|A\|^{4}A^{\dagger}\boldsymbol{\alpha}A\,\big(\mathrm{i}\,\alpha_{1}\alpha_{2}D_{1}\|A\|^{2}+\varrho_{3}A^{\dagger}\boldsymbol{\alpha}A\,(2c_{11}\varrho_{1}-c_{21}\varrho_{4})\big)\Big)\,x+\tilde{c}_{12}\,$,
$\Omega_{21}$ $=$ $\frac{\mathrm{i}\,c_{11}^{\ast}c_{12}\tilde{\gamma}^{\ast 4}\|A\|^{2}}{405\,\alpha_{1}^{4}\alpha_{2}^{4}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}(\mathrm{i}\,\varrho_{3}^{\ast}\varrho_{4}^{\ast}+\varrho_{2}^{\ast})\,\big(\gamma^{\ast}(\mathrm{i}\,\varrho_{3}^{\ast}\varrho_{4}^{\ast}+\varrho_{2}^{\ast})+A^{\dagger}\boldsymbol{\alpha}A\big)\,t^{5}$
$+\frac{\mathrm{i}\,\tilde{\gamma}^{\ast 2}\|A\|}{108\,\alpha_{1}^{3}\alpha_{2}^{3}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Big((\mathrm{i}\,\varrho_{3}^{\ast}\varrho_{4}^{\ast}+\varrho_{2}^{\ast})\,\big(c_{11}^{\ast}(2c_{12}\|A\|\,A^{\dagger}\boldsymbol{\alpha}A\,(\tilde{\gamma}^{\ast}x+3\gamma^{\ast})-3\gamma^{\ast}\alpha_{1}\alpha_{2}\varrho_{5}^{\ast})$
$+c_{12}(c_{21}^{\ast}\tilde{\gamma}^{\ast}\|A\|\,A^{\dagger}\boldsymbol{\alpha}A+3\gamma^{\ast}\alpha_{1}\alpha_{2}\varrho_{6}^{\ast})\big)-3c_{11}^{\ast}\alpha_{1}\alpha_{2}\varrho_{5}^{\ast}A^{\dagger}\boldsymbol{\alpha}A\Big)\,t^{4}$
$+\frac{\mathrm{i}}{27\alpha_{1}^{2}\alpha_{2}^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Big(c_{11}^{\ast}c_{12}\tilde{\gamma}^{\ast 2}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A(\mathrm{i}\,\varrho_{3}^{\ast}\varrho_{4}^{\ast}+\varrho_{2}^{\ast})\,x^{2}$
$+\tilde{\gamma}^{\ast}\|A\|\,A^{\dagger}\boldsymbol{\alpha}A\,\big(c_{12}c_{21}^{\ast}\tilde{\gamma}^{\ast}\|A\|\,(\mathrm{i}\,\varrho_{3}^{\ast}\varrho_{4}^{\ast}+\varrho_{2}^{\ast})-6c_{11}^{\ast}\alpha_{1}\alpha_{2}\varrho_{5}^{\ast}\big)\,x$
$-\tilde{\gamma}^{\ast 2}\|A\|^{2}(\mathrm{i}\,\varrho_{3}^{\ast}\varrho_{4}^{\ast}+\varrho_{2}^{\ast})\,\big(c_{11}^{\ast}\gamma^{\ast}\varrho_{8}^{\ast}-c_{12}\,(c_{31}^{\ast}A^{\dagger}\boldsymbol{\alpha}A-\gamma^{\ast}\varrho_{7}^{\ast})\big)-c_{11}^{\ast}\|A\|\,A^{\dagger}\boldsymbol{\alpha}A\,(\tilde{\gamma}^{\ast 2}\varrho_{8}^{\ast}\|A\|$
$+18\gamma^{\ast}\alpha_{1}\alpha_{2}\varrho_{5}^{\ast})-3\alpha_{1}\alpha_{2}\varrho_{5}^{\ast}\,(c_{21}^{\ast}\tilde{\gamma}^{\ast}\|A\|\,A^{\dagger}\boldsymbol{\alpha}A+3\gamma^{\ast}\alpha_{1}\alpha_{2}\varrho_{6}^{\ast})\Big)\,t^{3}$
$-\frac{\mathrm{i}\,\|A\|}{6\alpha_{1}\alpha_{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Big(3c_{11}^{\ast}\alpha_{1}\alpha_{2}\varrho_{5}^{\ast}A^{\dagger}\boldsymbol{\alpha}A\,x^{2}+A^{\dagger}\boldsymbol{\alpha}A\,(2c_{11}^{\ast}\tilde{\gamma}^{\ast}\varrho_{8}^{\ast}\|A\|+3c_{21}^{\ast}\alpha_{1}\alpha_{2}\varrho_{5}^{\ast})\,x$
$+\varrho_{8}^{\ast}\|A\|\,A^{\dagger}\boldsymbol{\alpha}A\,(6c_{11}^{\ast}\gamma^{\ast}+c_{21}^{\ast}\tilde{\gamma}^{\ast})+3\alpha_{1}\alpha_{2}\,\big(c_{31}^{\ast}\varrho_{5}^{\ast}A^{\dagger}\boldsymbol{\alpha}A-\gamma^{\ast}(\varrho_{5}^{\ast}\varrho_{7}^{\ast}-\varrho_{6}^{\ast}\varrho_{8}^{\ast})\big)\Big)\,t^{2}$
$-\frac{\mathrm{i}\,\varrho_{8}^{\ast}\|A\|^{2}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\big(A^{\dagger}\boldsymbol{\alpha}A\,(c_{11}^{\ast}x^{2}+c_{21}^{\ast}x+c_{31}^{\ast})-\gamma^{\ast}\varrho_{7}^{\ast}\big)\,t$
$+\frac{c_{11}^{\ast}c_{12}}{10\gamma^{\ast 2}(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A}\Big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}(2\gamma^{\ast}r^{\ast}-1)\,(\varrho_{3}^{\ast}\varrho_{4}^{\ast}\|A\|^{2}+A^{\dagger}\boldsymbol{\alpha}A)$
$-2\mathrm{i}\,\gamma^{\ast}\alpha_{1}\alpha_{2}\varrho_{3}^{\ast 2}\varrho_{4}^{\ast 2}\|A\|^{6}\Big)\,x^{5}$
$-\frac{1}{8\gamma^{\ast 2}(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A}\Big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}\big(2c_{11}^{\ast}c_{12}(\|A\|^{2}\varrho_{3}^{\ast}\,(2\gamma^{\ast}(r^{\ast}\varrho_{1}^{\ast}+\varrho_{4}^{\ast})-\varrho_{1}^{\ast})$
$+2\gamma^{\ast}A^{\dagger}\boldsymbol{\alpha}A)-(2\gamma^{\ast}r^{\ast}-1)\,(\varrho_{3}^{\ast}\varrho_{4}^{\ast}\|A\|^{2}+A^{\dagger}\boldsymbol{\alpha}A)\,(c_{11}^{\ast}c_{22}+c_{12}c_{21}^{\ast})\big)$
$+2\mathrm{i}\,\gamma^{\ast}\alpha_{1}\alpha_{2}\varrho_{3}^{\ast 2}\varrho_{4}^{\ast 2}\|A\|^{6}(c_{11}^{\ast}c_{22}+c_{12}c_{21}^{\ast})\Big)\,x^{4}$
$+\frac{1}{6\gamma^{\ast 2}(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Big($
$(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}\,\big(c_{12}\,(\varrho_{10}^{\ast}-3c_{21}^{\ast}A^{\dagger}\boldsymbol{\alpha}A\,(\varrho_{3}^{\ast}\|A\|^{2}(2\gamma^{\ast}(r^{\ast}\varrho_{1}^{\ast}+\varrho_{4}^{\ast})-\varrho_{1}^{\ast})+2\gamma^{\ast}A^{\dagger}\boldsymbol{\alpha}A))$
$+c_{22}((A^{\dagger}\boldsymbol{\alpha}A)^{2}(c_{21}^{\ast}(2\gamma^{\ast}r^{\ast}-1)-4c_{11}^{\ast}\gamma^{\ast})+2\mathrm{i}\,c_{11}^{\ast}\gamma^{\ast 2}\|A\|^{2}(\mathrm{i}\,\varrho_{3}^{\ast}\varrho_{4}^{\ast}+\varrho_{2}^{\ast})^{2}$
$-\varrho_{3}^{\ast}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A\,(2\gamma^{\ast}r^{\ast}-1)(2c_{11}^{\ast}\varrho_{1}^{\ast}-c_{21}^{\ast}\varrho_{4}^{\ast})-2\gamma^{\ast}c_{11}^{\ast}\varrho_{3}^{\ast}\varrho_{4}^{\ast}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A)$
$+c_{11}^{\ast}c_{32}A^{\dagger}\boldsymbol{\alpha}A\,(2\gamma^{\ast}r^{\ast}-1)(\varrho_{3}^{\ast}\varrho_{4}^{\ast}\|A\|^{2}+A^{\dagger}\boldsymbol{\alpha}A)\big)$
$-2\gamma^{\ast}\varrho_{3}^{\ast 2}\|A\|^{4}A^{\dagger}\boldsymbol{\alpha}A\,\big(\mathrm{i}\,\alpha_{1}\alpha_{2}\varrho_{9}^{\ast}\|A\|^{2}+c_{11}^{\ast}\varrho_{4}^{\ast}A^{\dagger}\boldsymbol{\alpha}A\,(2c_{12}\varrho_{1}^{\ast}+c_{22}\varrho_{4}^{\ast})\big)\Big)\,x^{3}$
$+\frac{1}{4\gamma^{\ast 2}(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}\big(4\mathrm{i}\,c_{12}\gamma^{\ast 2}H_{1}^{\ast}\|A\|^{2}(\mathrm{i}\,\varrho_{3}^{\ast}\varrho_{4}^{\ast}+\varrho_{2}^{\ast})$
$-D_{1}^{\ast}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A\,(2\gamma^{\ast}\,(2c_{12}-c_{22}r^{\ast})+c_{22})+2\gamma^{\ast}\varrho_{3}^{\ast}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A\,(c_{22}-c_{32}r^{\ast})(2c_{11}^{\ast}\varrho_{1}^{\ast}-c_{21}^{\ast}\varrho_{4}^{\ast})$
$+c_{32}\varrho_{3}^{\ast}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A\,(2c_{11}^{\ast}\varrho_{1}^{\ast}-c_{21}^{\ast}\varrho_{4}^{\ast})\big)-\varrho^{(2)^{\ast}}+4c_{12}\gamma^{\ast}\varrho_{3}^{\ast}\varrho_{4}^{\ast}D_{1}^{\ast}\|A\|^{4}(A^{\dagger}\boldsymbol{\alpha}A)^{2}\Big)\,x^{2}$
$+\frac{1}{2\gamma^{\ast 2}(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}\big(c_{31}^{\ast}c_{32}(A^{\dagger}\boldsymbol{\alpha}A)^{2}(2\gamma^{\ast}r^{\ast}-1)$
$+\|A\|^{2}(2\gamma^{\ast 2}H_{1}^{\ast}(\mathrm{i}\,c_{22}(\mathrm{i}\,\varrho_{3}^{\ast}\varrho_{4}^{\ast}+\varrho_{2}^{\ast})-2c_{12}(\mathrm{i}+\varrho_{1}^{\ast}\varrho_{3}^{\ast}))-2\gamma^{\ast}D_{1}^{\ast}A^{\dagger}\boldsymbol{\alpha}A\,(c_{22}-c_{32}r^{\ast})-c_{32}D_{1}^{\ast}A^{\dagger}\boldsymbol{\alpha}A)\big)$
$-2\gamma^{\ast}D_{1}^{\ast}\|A\|^{4}A^{\dagger}\boldsymbol{\alpha}A\,\big(\mathrm{i}\,\alpha_{1}\alpha_{2}D_{2}^{\ast}\|A\|^{2}-\varrho_{3}^{\ast}A^{\dagger}\boldsymbol{\alpha}A\,(2c_{12}\varrho_{1}^{\ast}+c_{22}\varrho_{4}^{\ast})\big)\Big)\,x+\tilde{c}_{21}\,$,
with further constants
$\varrho_{5}$ $=$ $\frac{\|A\|}{3\alpha_{1}\alpha_{2}}\Big(\big(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2}\big)\big(2c_{12}^{\ast}(\tilde{\gamma}\,x-3\gamma)+c_{22}^{\ast}\tilde{\gamma}\big)+2c_{12}^{\ast}\tilde{\gamma}\,(\mathrm{i}\,\varrho_{1}\varrho_{3}+1)\Big)\,$,
$\varrho_{6}$ $=$ $-\frac{\|A\|}{3\alpha_{1}\alpha_{2}}\Big(2c_{11}\tilde{\gamma}\,\big((\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})\,x-\mathrm{i}\,\varrho_{1}\varrho_{3}-1\big)+(6c_{11}\gamma+c_{21}\tilde{\gamma})\,\big(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2}\big)\Big)\,$,
$\varrho_{7}$ $=$ $c_{11}\big(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2}\big)\,x^{2}+\big(c_{21}(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})-2c_{11}(\mathrm{i}\,\varrho_{1}\varrho_{3}+1)\big)\,x+\mathrm{i}\,D_{1}-c_{31}\varrho_{2}-c_{21}\,$,
$\varrho_{8}$ $=$ $c_{12}^{\ast}(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})\,x^{2}+\big(c_{22}^{\ast}(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})+2c_{12}^{\ast}(\mathrm{i}\,\varrho_{1}\varrho_{3}+1)\big)\,x+\mathrm{i}\,D_{2}-c_{32}^{\ast}\varrho_{2}+c_{22}^{\ast}\,$,
$\varrho_{9}$ $=$ $\varrho_{1}\varrho_{4}(c_{12}^{\ast}c_{21}-c_{11}c_{22}^{\ast})+\varrho_{4}^{2}(c_{11}c_{32}^{\ast}+c_{12}^{\ast}c_{31}+c_{21}c_{22}^{\ast})-4c_{11}c_{12}^{\ast}(\varrho_{1}^{2}-\varrho_{4})\,$,
$\varrho_{10}$ $=$ $(A^{\dagger}\boldsymbol{\alpha}A)^{2}(c_{31}(2\gamma r-1)+2\gamma c_{21})+\varrho_{3}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A\,((2\gamma r-1)\,(2c_{11}+2c_{21}\varrho_{1}+c_{31}\varrho_{4})$
$+2\gamma\,(4c_{11}\varrho_{1}+c_{21}\varrho_{4}))-4\mathrm{i}\,c_{11}\gamma^{2}\|A\|^{2}(\mathrm{i}\,\varrho_{3}\varrho_{4}-\varrho_{2})\,(1+\mathrm{i}\,\varrho_{1}\varrho_{3})\,$,
$\varrho^{(j)}$ $=$ $-2\gamma^{2}\,(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}\|A\|^{2}\big(\mathrm{i}\,\varrho_{3}^{2}(2c_{11}\varrho_{1}-c_{21}\varrho_{4})\,(2c_{12}^{\ast}\varrho_{1}+c_{22}^{\ast}\varrho_{4})$
$+2c_{22}^{\ast}\varrho_{3}(c_{21}\varrho_{2}\varrho_{4}-c_{11}(\varrho_{1}\varrho_{2}-\varrho_{4}))+2c_{12}^{\ast}\varrho_{3}(4c_{11}\varrho_{1}+c_{21}(\varrho_{1}\varrho_{2}-\varrho_{4}))$
$-\mathrm{i}\,(2c_{11}+c_{21}\varrho_{2})\,(2c_{12}^{\ast}-c_{22}^{\ast}\varrho_{2})\big)+2\gamma\varrho_{3}^{2}\|A\|^{4}(A^{\dagger}\boldsymbol{\alpha}A)^{2}(2c_{11}\varrho_{1}-c_{21}\varrho_{4})\,(2c_{12}^{\ast}\varrho_{1}+c_{22}^{\ast}\varrho_{4})$
$-(-1)^{j}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}(2\gamma r-1)(c_{21}c_{32}^{\ast}+c_{22}^{\ast}c_{31})$
$+2\mathrm{i}\,\gamma\alpha_{1}\alpha_{2}\varrho_{3}\|A\|^{6}A^{\dagger}\boldsymbol{\alpha}A\,(D_{1}(2c_{12}^{\ast}\varrho_{1}+c_{22}^{\ast}\varrho_{4})-D_{2}(2c_{11}\varrho_{1}-c_{21}\varrho_{4}))\big)\,$,
with complex constants $\tilde{c}_{12}$ and $\tilde{c}_{21}$. The remaining condition (6.22) results in
$c_{31}c_{32}^{\ast}+\frac{\|A\|^{4}\,D_{1}D_{2}}{(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}}=\gamma\,(\tilde{c}_{12}-\tilde{c}_{21}^{\ast})\,.$
Now (6.61) determines a rather complicated class of quasi-rational solutions of the two-component vector FL equation. We emphasize that, in this case, a rational dependence on the two independent variables $x$ and $t$ enters in two different ways. First, because of the “degeneracy” of the linear system. Secondly, via integration of (5.8), since also the Lypunov equation is “degenerate”, i.e., the spectrum condition is violated.
We note that $\Omega_{11}\,e^{-2\mathrm{Re}(\theta)}$ and $\Omega_{22}\,e^{2\mathrm{Re}(\theta)}$ are polynomials of degree four in $x$ and $t$, whereas $\Omega_{12}$ and $\Omega_{21}$ are of degree five. It follows that $u_{\mu}^{\prime}\sim A_{\mu}\,e^{\mathrm{i}\,\varphi_{\mu}}$ as $x\to\infty$ or $t\to\infty$. The resulting solutions generically model rogue waves. $\Box$
6.2.4 The double root case
First, we consider the case $n=1$ and then the case $n=2$ with $\Gamma$ a $2\times 2$ Jordan block.
Example 6.18.
Let $n=1$. If the cubic characteristic equation (6.62) admits a single root $r_{1}$ and a double root $r_{2}$, then it is equivalent to
$r_{1}=\mathrm{i}\,(\alpha_{1}+\alpha_{2})+\frac{1}{2\gamma}-2\,r_{2}$ (6.93)
and
$|A_{1}|^{2}=\frac{(2\,r_{2}-\mathrm{i}\,\alpha_{2})\,(2\gamma\,r_{2}-1-2\mathrm{i}\,\alpha_{1}\gamma)^{2}}{4\alpha_{1}^{2}\gamma\,(\alpha_{2}-\alpha_{1})}\,,\quad|A_{2}|^{2}=-\frac{(2r_{2}-\mathrm{i}\,\alpha_{1})\,(2\gamma\,r_{2}-1-2\mathrm{i}\,\alpha_{2}\gamma)^{2}}{4\alpha_{2}^{2}\gamma\,(\alpha_{2}-\alpha_{1})}\,.$ (6.94)
The right hand sides of the last two equations have to be real and positive, which imposes constraints on the parameters, which we were unable to cast into a convenient and simple form. The linear system is now solved by
$\chi_{1}$ $=$ $c_{1}\,e^{\theta_{1}}+\Big(c_{2}\,x-\frac{c_{2}\,\gamma}{\alpha_{1}\alpha_{2}}\big(\mathrm{i}\,(\alpha_{1}+\alpha_{2})-2r_{2}\big)\,t+c_{3}\Big)\,e^{\theta_{2}}\,$,
$\chi_{2}$ $=$ $\frac{\|A\|^{2}}{(\alpha_{2}-\alpha_{1})\,A_{1}A_{2}}\Big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,\gamma^{-1}-\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\Big)^{-1}\Big(\chi_{1xx}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}\,\chi_{1x}-\frac{1}{4}\gamma^{-2}\chi_{1}$
$+\frac{\mathrm{i}}{2\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)\,\gamma^{-1}\chi_{1}\Big)\,$,
with complex constants $c_{1},c_{2},c_{3}$ and
$\theta_{\kappa}$ $=$ $r_{\kappa}x+\frac{1}{\alpha_{1}\alpha_{2}\gamma}\Big(\big(r_{\kappa}^{2}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})\,r_{\kappa}-\frac{1}{2}\alpha_{1}\alpha_{2}\big)\,\gamma^{2}+\mathrm{i}\,(\alpha_{1}+\alpha_{2})\,(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2})\,\gamma-\frac{1}{4}\Big)\,t\,.$
(1) Assuming $\mathrm{Re}(\gamma)\neq 0$, so that the spectrum condition for the Lyapunov equation is satisfied, the resulting solution of the two-component vector FL equation is
$u_{1}^{\prime}=e^{\mathrm{i}\,\varphi_{1}}\big(A_{1}+\frac{(A_{1}\chi_{1}^{\ast}-A_{2}^{\ast}\chi_{2}^{\ast})\,\Xi}{\|A\|\,A^{\dagger}\boldsymbol{\alpha}A\,\Omega}\big)\,,\qquad u_{2}^{\prime}=e^{\mathrm{i}\,\varphi_{2}}\big(A_{2}+\frac{(A_{2}\chi_{1}^{\ast}+A_{1}^{\ast}\chi_{2}^{\ast})\,\Xi}{\|A\|\,A^{\dagger}\boldsymbol{\alpha}A\,\Omega}\big)\,$,
where
$\Omega$ $=$ $\frac{\mathrm{i}\,\gamma^{\ast}|\Xi|^{2}+(A^{\dagger}\boldsymbol{\alpha}A)^{2}(|\chi_{1}|^{2}+|\chi_{2}|^{2})}{2\mathrm{Re}(\gamma)\,|(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\,$,
$\Xi$ $=$ $\|A\|\,\Big(\chi_{1x}-\big(\frac{1}{2}\gamma^{-1}+\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}}\big)\,\chi_{1}-\mathrm{i}\,(\alpha_{2}-\alpha_{1})\frac{A_{1}A_{2}}{\|A\|^{2}}\,\chi_{2}\Big)\,.$
If $c_{1}=0$, this solution is quasi-rational. If also $c_{2}\neq 0$ and $r_{2}\neq\frac{1}{2}\mathrm{i}\,(\alpha_{1}+\alpha_{2})$, the latter is localized in space and time on the plane wave background. It thus represents a rogue wave.
(2) If $\gamma$ is imaginary, i.e., $\gamma=-\mathrm{i}\,k$ with $k\in\mathbb{R}\setminus\{0\}$, we have
$|A_{1}|^{2}$ $=$ $\frac{(2\mathrm{i}\,r_{2}+\alpha_{2})\,(2\mathrm{i}\,k\,r_{2}+1+2\alpha_{1}k)^{2}}{4\alpha_{1}^{2}\,k\,(\alpha_{2}-\alpha_{1})}\,,\qquad|A_{2}|^{2}=-\frac{(2\mathrm{i}\,r_{2}+\alpha_{1})\,(2\mathrm{i}\,k\,r_{2}+1+2\alpha_{2}k)^{2}}{4\alpha_{2}^{2}\,k\,(\alpha_{2}-\alpha_{1})}\,.$
The imaginary parts vanish if and only if $\mathrm{Re}(r_{2})=0$. Hence
$r_{2}=\mathrm{i}\,s\,,\qquad s\in\mathbb{R}\,$,
and thus
$|A_{1}|^{2}$ $=$ $\frac{(\alpha_{2}-2s)\,(2ks-1-2\alpha_{1}k)^{2}}{4\alpha_{1}^{2}\,k\,(\alpha_{2}-\alpha_{1})}\,,\qquad|A_{2}|^{2}=-\frac{(\alpha_{1}-2s)\,(2ks-1-2\alpha_{2}k)^{2}}{4\alpha_{2}^{2}\,k\,(\alpha_{2}-\alpha_{1})}\,$,
which requires
$\frac{\alpha_{2}-2s}{k\,(\alpha_{2}-\alpha_{1})}>0\,,\qquad\frac{\alpha_{1}-2s}{k\,(\alpha_{2}-\alpha_{1})}<0\,$,
and implies $(\alpha_{1}-2s)(\alpha_{2}-2s)<0$. The Lyapunov equation can only be satisfied if either $c_{1}=0$ or $c_{2}=c_{3}=0$.
(i) $c_{1}=0$. The Lyapunov equation reduces to an equation, which does not dependent on the independent variables,
$(A^{\dagger}\boldsymbol{\alpha}A)^{2}\,\Big(|c_{3}|^{2}+\frac{\|A\|^{4}\,|\tilde{c}|^{2}}{|A_{1}|^{2}\,|A_{2}|^{2}\,(\alpha_{2}-\alpha_{1})^{2}}\Big)=k\,\|A\|^{2}\,\Big|\mathrm{i}\,c_{2}-c_{3}\,s+c_{3}\,\big(\frac{1}{2k}+\frac{A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}}\big)-\tilde{c}\,\Big|^{2}\,$,
where
$\tilde{c}$ $=$ $\Big((1-A^{\dagger}\boldsymbol{\alpha}A)\,k^{-1}+\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\Big)^{-1}\Big(2\mathrm{i}\,c_{2}\,s-c_{3}\,s^{2}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}(c_{2}+\mathrm{i}\,c_{3}\,s)+\frac{1}{4}k^{-2}c_{3}$
$-\frac{1}{2A^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{2}}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)\,k^{-1}c_{3}\Big)\,.$
Solving the differential equations for $\Omega$, we obtain
$\Omega$ $=$ $\frac{\|A\|^{2}}{k(A^{\dagger}\boldsymbol{\alpha}A)^{2}\,\varrho_{2}^{2}}\Big(\frac{\mathrm{i}\,k\,\tilde{s}\,|c_{2}|^{2}}{3\alpha_{1}\alpha_{2}}\varrho_{3}\,(\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}-\varrho_{3})\,(3xt^{2}-\frac{3\mathrm{i}\,\alpha_{1}\alpha_{2}}{k\tilde{s}}x^{2}t+\frac{\mathrm{i}\,k\tilde{s}}{\alpha_{1}\alpha_{2}}t^{3})$
$-\mathrm{i}\,\big(k|c_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}(\varrho_{2}-k\varrho_{1})+2\mathrm{i}\,\mathrm{Re}(c_{2}c_{3}^{\ast})\,\varrho_{3}(\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}-\varrho_{3})\big)\,(xt+\frac{\mathrm{i}\,k\tilde{s}}{2\alpha_{1}\alpha_{2}}t^{2})$
$-\big(c_{3}^{\ast}\,(\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}-\varrho_{3})-kc_{2}^{\ast}\,(\varrho_{2}-k\varrho_{1})\big)\,(kc_{2}\,(\varrho_{2}-k\varrho_{1})-c_{3}\varrho_{3})\,t$
$-\frac{|c_{2}|^{2}\varrho_{5}}{6kA^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{2}}\,x^{3}-\frac{1}{4kA^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{2}}\big(2\mathrm{Re}(c_{2}c_{3}^{\ast})\,\varrho_{5}-\mathrm{i}\,k|c_{2}|^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{2}\varrho_{6}\big)\,x^{2}$
$+\frac{1}{kA^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{2}\varrho_{4}^{2}}\big(\|A\|^{4}|c_{2}|^{2}\varrho_{1}^{2}\varrho_{7}+kc_{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{2}\varrho_{4}\,(A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}-k\varrho_{4})\,(\mathrm{i}\,c_{3}^{\ast}\varrho_{4}+c_{2}^{\ast}\|A\|^{2}\varrho_{1})$
$+\mathrm{i}\,k|c_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{2}\varrho_{4}^{2}\,(k\varrho_{1}-\varrho_{2})\,\big(\mathrm{i}\,ks\,(k\varrho_{1}-\varrho_{2})+\varrho_{3}\big)+\varrho_{8}\big)\,x+c\Big)\,$,
with
$\tilde{s}$ $=$ $\mathrm{i}\,(\alpha_{1}+\alpha_{2}-2s)\,$,
$\varrho_{1}$ $=$ $-2\,s\,A^{\dagger}\boldsymbol{\alpha}A+A^{\dagger}\boldsymbol{\alpha}^{2}A\,$,
$\varrho_{2}$ $=$ $A^{\dagger}\boldsymbol{\alpha}A\,(A^{\dagger}\boldsymbol{\alpha}A-1)-k\,\alpha_{1}\alpha_{2}\|A\|^{2}\,$,
$\varrho_{3}$ $=$ $-\mathrm{i}\,k^{2}\,\big(A^{\dagger}\boldsymbol{\alpha}A\,s^{2}-(\alpha_{1}\alpha_{2}\,\|A\|^{2}+A^{\dagger}\boldsymbol{\alpha}^{2}A)\,s+\alpha_{1}\alpha_{2}\,A^{\dagger}\boldsymbol{\alpha}A\big)$
$-\mathrm{i}\,k\,\big(s\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}+\frac{1}{2}\alpha_{1}\alpha_{2}\|A\|^{2}+\frac{1}{2}\varrho_{1}\big)+\frac{1}{4}\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}A\,(2A^{\dagger}\boldsymbol{\alpha}A-1)\,$,
$\varrho_{4}$ $=$ $\frac{1}{2k^{2}}\Big(\|A\|^{2}\,\big(2\mathrm{i}\,(\mathrm{i}\,ks\,\varrho_{2}+\varrho_{3})+\varrho_{2}\big)+2kA^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2}\Big)\,$,
$\varrho_{5}$ $=$ $(A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{2}\,(1-2ks)\,(k\varrho_{4}-A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2})-2\alpha_{1}\alpha_{2}\|A\|^{2}(k(A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{2}^{2}+\|A\|^{2}\varrho_{3}^{2})\,$,
$\varrho_{6}$ $=$ $\|A\|^{2}\varrho_{1}\,(1-2ks)-2(k\varrho_{4}-A^{\dagger}\boldsymbol{\alpha}A\,\varrho_{2})\,$,
$\varrho_{7}$ $=$ $(\alpha_{1}\alpha_{2}\|A\|^{2}-s\,A^{\dagger}\boldsymbol{\alpha}A)\,(k\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{2}^{2}+\|A\|^{2}\varrho_{3}^{2})\,$,
$\varrho_{8}$ $=$ $|c_{2}|^{2}A^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{4}\varrho_{1}\varrho_{3}^{2}\varrho_{4}+\frac{1}{2}\varrho_{4}\,\Big(4\mathrm{i}\,\mathrm{Im}(c_{2}c_{3}^{\ast})\,\big(k\,s\,A^{\dagger}\boldsymbol{\alpha}A\,\|A\|^{2}\varrho_{3}\varrho_{4}\,(k\varrho_{1}-\varrho_{2})+\mathrm{i}\,\|A\|^{2}\varrho_{1}\varrho_{7}\big)$
$+\mathrm{i}\,k\,c_{2}^{\ast}c_{3}\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}\|A\|^{2}\varrho_{1}\varrho_{2}\varrho_{4}\,(1-2ks)-|c_{3}|^{2}\varrho_{4}\varrho_{5}\Big)\,$,
and a constant $c$ for which (6.22) requires
$\mathrm{Im}(c)=\frac{1}{2k\varrho_{4}^{2}}\Big(\varrho_{1}\,\big(k(A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{2}^{2}+\|A\|^{2}\varrho_{3}^{2}\big)\,\big(2\mathrm{Im}(c_{2}c_{3}^{\ast})\varrho_{4}-|c_{2}|^{2}\|A\|^{2}\varrho_{1}\big)-|c_{3}|^{2}\varrho_{3}^{2}\varrho_{4}^{2}\Big)\,.$
An example from the resulting class of solutions of the two-component vector FL equation is presented in Fig. 14.
Figure 14: Plots of the absolute values of the two components of a “double pole solution” from the class in Example 6.18 (2). Here we set $\alpha_{1}=1,\alpha_{2}=2,k=1,s=3/4,c_{1}=0,c_{2}=1+\mathrm{i}$ and $c_{3}=1+(353/1365)\,\mathrm{i}$, so that the constraint from the Lyapunov equation is satisfied.
(ii) $c_{2}=c_{3}=0$. The Lyapunov equation requires $c_{1}=0$, so that $u_{1}^{\prime}$ and $u_{2}^{\prime}$ reduce to the plane wave background solutions. $\Box$
Remark 6.19.
Let $c_{1}=0$ and $r_{2}=\frac{1}{2}\mathrm{i}\,(\alpha_{1}+\alpha_{2})$ in Example 6.18. Then $\chi_{1}$, $\chi_{2}$, and thus also $u^{\prime}$, show a rational dependence only on $x$. This particular choice for $r_{2}$ actually follows from (6.94) by just assuming $\mathrm{Re}(r_{2})=0$ and $\mathrm{Re}(\gamma)\neq 0$. Rogue waves thus require $\mathrm{Re}(r_{2})\neq 0$.
In the case $\mathrm{Re}(r_{2})=0$, (6.94) further amounts to $\mathrm{Im}(\gamma)=(1-2\alpha_{2}|A_{2}|^{2})/(\alpha_{2}-\alpha_{1})$, $\mathrm{Re}(\gamma)^{2}=4\left(-\alpha_{1}|A_{1}|^{2}+\alpha_{2}|A_{2}|^{2}\left(\alpha_{2}|A_{2}|^{2}\left(3-2\alpha_{2}|A_{2}|^{2}\right)-2\right)+1\right)/((\alpha_{2}-\alpha_{1})^{2}(2\alpha_{2}|A_{2}|^{2}-1))$, and $\alpha_{1}|A_{1}|^{2}+\alpha_{2}|A_{2}|^{2}=1$. $\Box$
Next we present a second-order version of the above class of solutions in Example 6.18.
Example 6.20.
Let $n=2$ and
$\Gamma=\left(\begin{array}[]{cc}\gamma&0\\ 1&\gamma\end{array}\right)\,,\qquad\chi_{i}=\left(\begin{array}[]{c}\chi_{i1}\\ \chi_{i2}\end{array}\right)\qquad i=1,2\,.$
Again, we assume that $r_{1}$ is a single root and $r_{2}$ is a double root of the cubic equation (6.62). Then $r_{1}$ and $A_{\mu}$, $\mu=1,2$, are given by (6.93) and (6.94), respectively. Now the linear system is solved by
$\chi_{11}$ $=$ $c_{1}e^{\theta_{1}}+(f_{1}x+f_{2})\,e^{\theta_{2}}\,$,
$\chi_{12}$ $=$ $\Big(\frac{c_{1}(v_{1}-v_{2})}{2v_{2}\gamma^{2}}x+\frac{c_{1}v_{3}}{2\alpha_{1}\alpha_{2}v_{2}\gamma^{2}}t+c_{6}\Big)\,e^{\theta_{1}}$
$+\Big(\frac{f_{1}v_{5}}{24v_{4}\gamma^{3}}x^{3}+\frac{2f_{1}\gamma\,(v_{4}v_{6}+v_{5})+f_{2}v_{4}v_{5}}{8v_{4}^{2}\gamma^{3}}x^{2}+f_{3}x+f_{4}\Big)\,e^{\theta_{2}}\,$,
$\chi_{2}$ $=$ $\frac{\|A\|^{2}}{(\alpha_{2}-\alpha_{1})\,A_{1}A_{2}}\Big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,\Gamma^{-1}-\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}I_{n}\Big)^{-1}\Big(\chi_{1xx}-\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}\,\chi_{1x}-\frac{1}{4}\Gamma^{-2}\chi_{1}$
$+\frac{\mathrm{i}}{2\|A\|^{2}\,A^{\dagger}\boldsymbol{\alpha}A}\big((\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}-(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1-2A^{\dagger}\boldsymbol{\alpha}A)\big)\,\Gamma^{-1}\chi_{1}\Big)\,$,
where
$v_{1}$ $=$ $8\big(8r_{2}^{3}-8\mathrm{i}\,(\alpha_{1}+\alpha_{2})\,r_{2}^{2}-2\big(\alpha_{1}^{2}+\alpha_{2}^{2}+3\alpha_{1}\alpha_{2}\big)\,r_{2}+\mathrm{i}\,\alpha_{1}\alpha_{2}\,(\alpha_{1}+\alpha_{2})\big)\,\gamma^{3}\,$,
$v_{2}$ $=$ $4\big(9r_{2}^{2}-6\mathrm{i}\,(\alpha_{1}+\alpha_{2})\,r_{2}-(\alpha_{1}+\alpha_{2})^{2}\big)\,\gamma^{2}-4(3r_{2}-\mathrm{i}\,(\alpha_{1}+\alpha_{2}))\,\gamma+1\,$,
$v_{3}$ $=$ $\big(8r_{2}^{2}-4\mathrm{i}\,(\alpha_{1}+\alpha_{2})\,r_{2}-\alpha_{1}\alpha_{2}\big)\,v_{2}\gamma^{2}-\big(4r_{2}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})\big)\,v_{1}\gamma+v_{1}\,$,
$v_{4}$ $=$ $2(\mathrm{i}\,(\alpha_{1}+\alpha_{2})-3r_{2})\,\gamma+1)\,$,
$v_{5}$ $=$ $(2r_{2}\gamma-1)\big(4(r_{2}^{2}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})\,r_{2}-\alpha_{1}\alpha_{2})\,\gamma^{2}+2(\mathrm{i}\,(\alpha_{1}+\alpha_{2})-2r_{2})\,\gamma+1\big)\,$,
$v_{6}$ $=$ $4\big(3r_{2}^{2}-2\mathrm{i}\,(\alpha_{1}+\alpha_{2})\,r_{2}-\alpha_{1}\alpha_{2}\big)\gamma^{2}+1\,$,
and $f=(f_{1},\ldots,f_{4})^{T}$ satisfies $f_{t}=Cf$ with the constant nilpotent matrix
$C=-\frac{1}{\alpha_{1}\alpha_{2}}\,\left(\begin{array}[]{cccc}0&0&0&0\\ \tilde{s}\,\gamma&0&0&0\\ \frac{1}{4}\big(B_{1}v_{4}+2(v_{4}v_{6}+v_{5})\,\tilde{s}\gamma\big)v_{4}^{-2}\gamma^{-2}&\frac{1}{4}v_{4}^{-1}v_{5}\tilde{s}\gamma^{-2}&0&0\\ \frac{1}{2}B_{2}v_{4}^{-2}\gamma^{-1}&\frac{1}{4}B_{1}v_{4}^{-1}\gamma^{-2}&\tilde{s}\,\gamma&0\end{array}\right)\,$,
where
$\tilde{s}$ $=$ $\mathrm{i}\,(\alpha_{1}+\alpha_{2})-2r_{2}\,$,
$B_{1}$ $=$ $-2(2r_{2}^{2}-2\mathrm{i}\,(\alpha_{1}+\alpha_{2})\,r_{2}-\alpha_{1}\alpha_{2})\,v_{4}\gamma^{2}-(v_{4}+v_{5})\,$,
$B_{2}$ $=$ $2v_{4}^{2}\tilde{s}\gamma-v_{4}v_{6}-v_{5}\,.$
Consequently, we have
$f_{1}$ $=$ $c_{2}\,$,
$f_{2}$ $=$ $-\frac{c_{2}\tilde{s}\gamma}{\alpha_{1}\alpha_{2}}\,t+c_{3}\,$,
$f_{3}$ $=$ $\frac{1}{8(\alpha_{1}\alpha_{2}v_{4}\gamma)^{2}}\Big(c_{2}v_{4}v_{5}\tilde{s}^{2}\gamma\,t^{2}-2\alpha_{1}\alpha_{2}\big(((2c_{2}v_{6}\gamma+c_{3}v_{5})\,v_{4}+2c_{2}v_{5}\gamma)\,\tilde{s}+c_{2}B_{1}v_{4}\big)\,t$
$+8c_{4}(\alpha_{1}\alpha_{2}v_{4}\gamma)^{2}\Big)\,$,
$f_{4}$ $=$ $\frac{1}{24(\alpha_{1}\alpha_{2})^{3}(v_{4}\gamma)^{2}}\Big(-c_{2}v_{4}v_{5}\tilde{s}^{3}\gamma^{2}\,t^{3}+3\big((2(v_{4}v_{6}+v_{5})\,c_{2}\gamma+c_{3}v_{4}v_{5})\,\tilde{s}+2c_{2}B_{1}v_{4}\big)\,\alpha_{1}\alpha_{2}\tilde{s}\gamma\,t^{2}$
$-6\alpha_{1}^{2}\alpha_{2}^{2}\,(2c_{2}B_{2}\gamma+c_{3}B_{1}v_{4}+4c_{4}v_{4}^{2}\tilde{s}\gamma^{3})\,t+24c_{5}(\alpha_{1}\alpha_{2})^{3}(v_{4}\gamma)^{2}\Big)\,$,
with complex constants $c_{1},\ldots,c_{6}$. Assuming $\mathrm{Re}(\gamma)\neq 0$, the expressions for $\Omega$ are obtained via those in Example 6.16. Inserting the results in (6.61), then yields a class of solutions of the two-component vector FL equation. We disregard the case $\mathrm{Re}(\gamma)=0$ because of its complexity. $\Box$
Example 6.21.
Choosing the special data of Remark 6.19 in the class of solutions in Example 6.20, a corresponding example of a solution of the two-component vector FL equation is displayed in Fig. 15.
Figure 15: Plots of the absolute values of the two components of a solution from the class in Example 6.20, with data according to the special case in Remark 6.19. We chose $\alpha_{2}=2$, $r_{2}=3\mathrm{i}/2$ (so that $\alpha_{1}=1$), $A_{1}=\sqrt{3}/2$, $A_{2}=\sqrt{2}/4$, $\mathrm{Im}(\gamma)=\sqrt{3}/2$, $c_{1}=c_{6}=0$ and $c_{2}=\cdots=c_{5}=1$.
$\Box$
Also see [11] for rogue wave solutions in the double root case.
7 Conclusions
We derived a vectorial binary Darboux transformation (bDT) for matrix versions of the first negative flow of the Kaup-Newell (KN) hierarchy. For a reduction of the latter system to matrix generalizations (see (1.2)) of the Fokas-Lenells (FL) equation, we found a corresponding reduction of the bDT to a vectorial Darboux transformation of the latter and generated various classes of matrix soliton solutions of (1.2) from the trivial seed solution $u=0$.
Furthermore, we explored in detail the vectorial Darboux transformation with a plane wave seed solution in the case of the two-component vector FL equation, recovering in a different and more systematic way classes of solutions obtained previously in [7, 11, 16]. But we also presented solutions that have not been obtained before to the best of our knowledge. This concerns in particular solutions corresponding to cases, where the matrix $\Gamma$ in the Lyapunov equation does not satisfy the spectrum condition, so that the Lyapunov equation does not have a unique solution.
An interesting feature, known to show up in vector NLS equations, is “beating solitons” (see [34, 35, 36, 32], and references cited there). We presented corresponding solutions for the two-component vector FL equation in Examples 6.5 and 6.6, apparently for the first time. As in the cited work on beating soliton solutions of vector NLS equations, here the relative wave number has been restricted to zero. Similar results have been obtained in [34] for the Manakov system (i.e., the two-component vector NLS equation), also see [36, 32].
The stage is also set to generate solutions of vector FL equations with more components (see [17] for some results for the three-component vector FL equation), and of matrix versions beyond the vector case, from a non-vanishing seed solution. In the matrix (not vector) case, we then have to deal with a linear system having left as well as right matrix coefficients, so that a vectorization map is required in order to solve it. We shall return to this in a separate work.
We would like to highlight again the nonlinear superposition property of a vectorial Darboux transformation, for solutions generated from the same seed solution, as formulated in Remarks 4.4 and 5.6 for the specific integrable equations studied in this work (also see Appendix Appendix D: Superposition of two solutions from the class in Example 6.5 for a simple example). In our opinion, this is a major advantage over the classical iterative Darboux transformation method.
Finally, we have found further classes of integrable equations that fit into the bidifferential calculus framework.
For generating the plots in this work and verifying examples of solutions, we used Mathematica [37].
Appendix A: Some facts about the Lyapunov equation
We consider the Lyapunov equation
$\Gamma\,\Omega+\Omega\,\Gamma^{\dagger}=K\,$, (A.1)
with $n\times n$ matrices $\Gamma,\Omega,K$. Here we think of $\Gamma$ and $K$ as given matrices and look for a solution $\Omega$.
Let $\Gamma=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n})$, with $\gamma_{i}\neq-\gamma_{j}^{\ast}$, $i,j=1,\ldots,n$. Then (A.1) reads
$(\gamma_{i}-\gamma_{j}^{\ast})\,\Omega_{ij}=K_{ij}\qquad i,j=1,\ldots,n\,.$
If $\Gamma$ satisfies the spectrum condition, the Lyapunov equation possesses a unique solution, given by the Cauchy-like matrix
$\Omega=\left(\frac{K_{ij}}{\gamma_{i}+\gamma_{j}^{\ast}}\right)\,.$
If, for some values of $i$ and $j$, $\gamma_{i}=\gamma_{j}^{\ast}$, then a solution can only exist if the corresponding component $K_{ij}$ of $K$ vanishes, and $\Omega_{ij}$ is left undetermined.
Let $\Gamma$ be the lower triangular $n\times n$ Jordan block
$\Gamma=\left(\begin{array}[]{ccccc}\gamma&0&\cdots&\cdots&0\\ 1&\gamma&\ddots&\ddots&0\\ 0&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&0\\ 0&\cdots&0&1&\gamma\end{array}\right)\,$, (A.7)
with $\gamma\in\mathbb{C}$. If the spectrum condition holds, we have the following result [33].
Proposition A.1.
Let $\mathrm{Re}(\gamma)\neq 0$. For $1\leq k<n$, let $K_{(k)}$ and $\Gamma_{(k)}$ be the $k\times k$ upper left part of $K$, respectively $\Gamma$, and $\Omega_{(k)}$ the solution of
$\Gamma_{(k)}\,\Omega_{(k)}+\Omega_{(k)}\,\Gamma_{(k)}^{\dagger}=K_{(k)}\,.$
For $k=1,\ldots,n-1$, we have
$\Omega_{(k+1)}=\left(\begin{array}[]{cc}\Omega_{(k)}&B_{(k)}\\ B_{(k)}^{\dagger}&\omega_{(k+1)}\end{array}\right)\,$,
where
$B_{(k)}:=\hat{\Gamma}_{(k)}^{-1}\,\Big((K_{1,k+1},K_{2,k+1},\ldots,K_{k,k+1})^{T}-\Omega_{(k)}(0,\ldots,0,1)^{T}\Big)\,$,
$\omega_{(k+1)}:=\frac{1}{\kappa}\Big(K_{k+1,k+1}-2\,\mathrm{Re}[(0,\ldots,0,1)B_{(k)}]\Big)\,$,
and $\hat{\Gamma}_{(k)}$ is the Jordan block $\Gamma_{(k)}$ with $\gamma$ replaced by $\kappa=2\,\mathrm{Re}(\gamma)$. $\Box$
This proposition allows to recursively compute the solution of the Lyapunov equation (A.1) with a Jordan matrix $\Gamma_{(n)}$. The inverse of $\Omega_{(n)}$ can also be recursively computed via
$\Omega_{(k+1)}^{-1}=\left(\begin{array}[]{cc}\Omega_{(k)}^{-1}+S_{\Omega_{(k)}}^{-1}\Omega_{(k)}^{-1}B_{(k)}B_{(k)}^{\dagger}\Omega_{(k)}^{-1}&-S_{\Omega_{(k)}}^{-1}\Omega_{(k)}^{-1}B_{(k)}\\ -S_{\Omega_{(k)}}^{-1}B_{(k)}^{\dagger}\Omega_{(k)}^{-1}&S_{\Omega_{(k)}}^{-1}\end{array}\right)\,$, (A.11)
with the scalar Schur complement (a quasi-determinant)
$S_{\Omega_{(k)}}=\omega_{(k+1)}-B_{(k)}^{\dagger}\Omega_{(k)}^{-1}B_{(k)}\,.$
Example A.2.
For $n=2$, we obtain
$\Omega_{(2)}=\frac{1}{\kappa}\left(\begin{array}[]{cc}K_{11}&K_{12}-\kappa^{-1}K_{11}\\ K_{21}-\kappa^{-1}K_{11}&K_{22}-\kappa^{-1}(K_{12}+K_{21})+2\kappa^{-2}K_{11}\end{array}\right)\,.$
Its determinant is
$\det(\Omega_{(2)})=\kappa^{-4}\,K_{11}^{2}+\kappa^{-2}\,\det(K)\,.$ (A.13)
$\Box$
Proposition A.3.
Let $\Gamma$ be the $n\times n$ lower triangular Jordan block with imaginary eigenvalue $\gamma$. Then the Lyapunov equation (A.1) is equivalent to
$K_{11}=0\,$,
$\Omega_{i-1,j}=K_{i1}\qquad i=2,3,\ldots,n\,$,
$\Omega_{1,j-1}=K_{1j}\qquad i=2,3,\ldots,n\,$,
$\Omega_{i-1,j}+\Omega_{i,j-1}=K_{ij}\qquad i,j=2,3,\ldots,n\,.$
We note that $\Omega_{nn}$ drops out of the Lyapunov equation.
Proof.
Expressed in components, the Lyapunov equation reads
$\sum_{k=1}^{n}\Big(\Gamma_{ik}\Omega_{kj}+\Omega_{ik}\Gamma^{\dagger}_{kj}\Big)=\Gamma_{ii}\Omega_{ij}+\Gamma_{i,i-1}\Omega_{i-1,j}+\Omega_{ij}\Gamma^{\dagger}_{jj}+\Omega_{i,j-1}\Gamma^{\dagger}_{j-1,j}=K_{ij}\,$,
where $i,j=1,\ldots,n$. This is equivalent to
$(\gamma+\gamma^{\ast})\,\Omega_{11}=K_{11}\,$,
$(\gamma+\gamma^{\ast})\,\Omega_{i1}+\Omega_{i-1,j}=K_{i1}\qquad i=2,3,\ldots,n\,$,
$(\gamma+\gamma^{\ast})\,\Omega_{1j}+\Omega_{1,j-1}=K_{1j}\qquad i=2,3,\ldots,n\,$,
$(\gamma+\gamma^{\ast})\,\Omega_{ij}+\Omega_{i-1,j}+\Omega_{i,j-1}=K_{ij}\qquad i,j=2,3,\ldots,n\,.$
If $\gamma$ is imaginary, this reduces to the equations in the proposition. ∎
Example A.4.
For the $2\times 2$ Jordan block with imaginary eigenvalue, the Lyapunov equation imposes the constraints $K_{11}=0$, $K_{21}=K_{12}$, and only determines part of $\Omega$ via $\Omega_{11}=K_{12}$ and $\Omega_{12}+\Omega_{21}=K_{22}$. $\Box$
Appendix B: Special examples from the class of solutions in Example 6.5
Let us choose $\alpha=c_{0}=\gamma=\|A\|=1$ in the class of solutions of the two-component vector FL equation in Example 6.5. Then we have $w=1-\mathrm{i}$. If $c_{1}=c_{2}=1$, the solution is
$u_{1}^{\prime}$ $=$ $\frac{(-1+\mathrm{i})\,e^{(2+\mathrm{i})\,t+\mathrm{i}\,x}-2\mathrm{i}\,e^{t+x}+e^{\mathrm{i}\,t+(2+\mathrm{i}\,)x}\,(1-\mathrm{i}+e^{2t})}{(1+\mathrm{i})\,e^{2t}+2\,e^{(1+\mathrm{i})(t+x)}+e^{2x}\,(1+\mathrm{i}+e^{2t})}\,$,
$u_{2}^{\prime}$ $=$ $-\frac{2\,e^{t+x}\,(e^{(1+\mathrm{i})\,t+\mathrm{i}\,x}+\mathrm{i}\,e^{x})}{(1+\mathrm{i})\,e^{2t}+(1+\mathrm{i})\,e^{2x}+2\,e^{(1+\mathrm{i})(t+x)}+e^{2(t+x)}}\,.$ (B.1)
We note that the dark soliton condition in Example 6.5) is satisfied for the corresponding solution with $c_{1}=0$ and $c_{2}=1$ (which coincides with the $c_{1}=1,c_{2}=0$ solution with $w$ replaced by w-w). It is given by
$u_{1}^{\prime}=e^{\mathrm{i}\,(t+x)}\,\Big(1-\frac{2\,e^{-x}}{e^{x}+(1+\mathrm{i})\,e^{-x}}\Big)\,,\qquad u_{2}^{\prime}=-\frac{2\,e^{\mathrm{i}\,(t+x)}}{e^{x}+(1+\mathrm{i})\,e^{-x}}\,.$ (B.2)
Plots of the absolute squares of the components of $u^{\prime}$ are presented in Fig. 16, respectively Fig. 17.
If $c_{1}=1$ and $c_{2}=0$, we obtain instead
$u_{1}^{\prime}=e^{\mathrm{i}\,(t+x)}\,\frac{1+e^{2t}-\mathrm{i}}{1+e^{2t}+\mathrm{i}}\,,\qquad u_{2}^{\prime}=-\frac{2\,\mathrm{i}\,e^{t}}{1+e^{2t}+\mathrm{i}}\,$, (B.3)
so that $|u_{1}^{\prime}|=1$ and $u_{2}^{\prime}$ represents a bright soliton. This solution has a special structure, see Appendix Appendix C: An almost trivial class of solutions of the two-component vector FL equation.
In both cases, there is no “beating”. In fact, with the chosen data, we have $\mathrm{Re}\big(\gamma\,(w^{\ast}-1)\big)=0$, so that the velocity $v$, defined in Example 6.5, is infinite and the beating period $T^{\prime}$ should then be expected to be zero.
Figure 16: Plots of the absolute squares of the two components, and of $\|u^{\prime}\|^{2}$, for the solution (B.1) of the two-component vector FL equation. Comparison with Fig. 17 shows that the dark, respectively bright soliton part for $t>0$ appears since $c_{2}\neq 0$. Correspondingly, the bright soliton part of $u_{2}^{\prime}$ for $x>0$ shows up because $c_{1}\neq 0$.
Figure 17: Plots of the absolute squares of the two components of the solution (B.2). The data are the same as used in Fig. 16, except that now $c_{1}=0$.
The class of solutions in Example 6.5, with $c_{2}=0$, also includes cases where $u^{\prime}$ is localized in $t$ (instead of $x$). An example is presented in Fig. 18.
Figure 18: Plots of the absolute squares of the two components of a solution from the class in Example 6.5, with $c_{2}=0$. Here we chose $\alpha=1/4$, $c_{0}=c_{1}=\|A\|=\gamma=1$, so that $w=\frac{1}{4}\sqrt{15+4\,\mathrm{i}}$.
Appendix C: An almost trivial class of solutions of the two-component vector FL equation
Inserting the ansatz
$u_{1}(x,t)=e^{\mathrm{i}\,\alpha_{1}\,x}\,f_{1}(t)\,,\qquad u_{2}(x,t)=e^{\mathrm{i}\,\alpha_{2}\,x}\,f_{2}(t)\,$,
with real constants $\alpha_{1},\alpha_{2}$, and functions $f_{\mu}$, $\mu=1,2$, of $t$ only, reduces the two-component vector FL equation (with $\boldsymbol{\sigma}=\mathrm{diag}(1,1)$) to the ODE system
$\alpha_{1}\,f_{1t}=\mathrm{i}\,\big(2\alpha_{1}|f_{1}|^{2}-1+(\alpha_{1}+\alpha_{2})|f_{2}|^{2})\,f_{1}\,,\qquad\alpha_{2}f_{2t}=\mathrm{i}\,\big(2\alpha_{2}|f_{2}|^{2}-1+(\alpha_{1}+\alpha_{2})|f_{1}|^{2})\,f_{2}\,.$
If $\alpha_{1}\alpha_{2}\neq 0$, this implies $|f_{1}|=a_{1}$ and $|f_{2}|=a_{2}$ with non-negative real constants $a_{1},a_{2}$, so that the above system becomes equivalent to the two linear homogeneous ODEs
$\alpha_{1}\,f_{1t}=\mathrm{i}\,\big(2\alpha_{1}a_{1}^{2}-1+(\alpha_{1}+\alpha_{2})a_{2}^{2})\,f_{1}\,,\qquad\alpha_{2}\,f_{2t}=\mathrm{i}\,\big(2\alpha_{2}a_{2}^{2}-1+(\alpha_{1}+\alpha_{2})a_{1}^{2})\,f_{2}\,.$
If the brackets on the right hand sides are different from zero, we obtain
$f_{1}=A_{1}\,e^{\mathrm{i}\,\big(2a_{1}^{2}-\alpha_{1}^{-1}+(1+\alpha_{1}^{-1}\alpha_{2})a_{2}^{2}\big)\,t}\,,\qquad f_{2}=A_{2}\,e^{\mathrm{i}\,\big(2a_{2}^{2}-\alpha_{2}^{-1}+(1+\alpha_{1}\alpha_{2}^{-1})a_{1}^{2}\big)\,t}\,$,
with complex constants $A_{1},A_{2}$. (B.3) has the form of our ansatz, but $|u_{2}^{\prime}|$ is not constant. This apparent contradiction is resolved by noting that, in this example, we have $\alpha_{2}=0$ and $|f_{1}|^{2}=\alpha_{1}^{-1}$, so that only $|f_{1}|$ is conserved and left and right hand sides of the ODE for $f_{2}$ vanish separately, hence $f_{2}$ can be any function of $t$.
Appendix D: Superposition of two solutions from the class in Example 6.5
Let $(\gamma_{i},\chi_{1}^{(i)},\chi_{2}^{(i)},\Xi^{(i)},\Omega^{(i)})$, $i=1,2$, with $\gamma_{i}\neq-\gamma_{j}^{\ast}$, $i,j=1,2$, determine $n=1$ solutions of the two-component FL equation, obtained via the Darboux transformation from the plane wave seed $u=(A_{1}\,e^{\mathrm{i}\,\varphi},0)$. Setting $\Gamma=\mathrm{diag}(\gamma_{1},\gamma_{2})$ and
$\chi_{1}=\left(\begin{array}[]{c}\chi_{1}^{(1)}\\ \chi_{1}^{(2)}\end{array}\right)\,,\quad\chi_{2}=\left(\begin{array}[]{c}\chi_{2}^{(1)}\\ \chi_{2}^{(2)}\end{array}\right)\,,\quad\Xi=\left(\begin{array}[]{c}\Xi^{(1)}\\ \Xi^{(2)}\end{array}\right)\,,\quad\Omega=\left(\begin{array}[]{cc}\Omega^{(1)}&\Omega^{(12)}\\ \Omega^{(21)}&\Omega^{(2)}\end{array}\right)\,$,
where
$\Omega^{(12)}$ $=$ $\frac{1}{\gamma_{1}+\gamma_{2}^{\ast}}\Big(\frac{\mathrm{i}\,\gamma_{2}^{\ast}}{\alpha^{2}\|A\|^{4}}\Xi^{(1)}\Xi^{(2)\ast}+\chi_{1}^{(1)}\chi_{1}^{(2)\ast}+\chi_{2}^{(1)}\chi_{2}^{(2)\ast}\Big)\,$,
$\Omega^{(21)}$ $=$ $\frac{1}{\gamma_{2}+\gamma_{1}^{\ast}}\Big(\frac{\mathrm{i}\,\gamma_{1}^{\ast}}{\alpha^{2}\|A\|^{4}}\Xi^{(2)}\Xi^{(1)\ast}+\chi_{1}^{(2)}\chi_{1}^{(1)\ast}+\chi_{2}^{(2)}\chi_{2}^{(1)\ast}\Big)\,$,
then
$u^{\prime}=e^{\mathrm{i}\,\varphi}\,\boldsymbol{U}_{0}\left(\begin{array}[]{r}\|A\|+(\alpha\,\|A\|^{2})^{-1}\chi_{1}^{\dagger}\,\Omega^{-1}\,\Xi\\ (\alpha\,\|A\|^{2})^{-1}\chi_{2}^{\dagger}\,\Omega^{-1}\,\Xi\end{array}\right)$
is also a solution. Here $\boldsymbol{U}_{0}$ is the constant unitary matrix in (6.5). An example, where $\boldsymbol{U}_{0}$ is the identity matrix, is displayed in Fig. 19.
Figure 19: Plots of the absolute values of the two components of a superposition of two solution from the class in Example 6.5 with $\alpha=A_{1}=1$, $A_{2}=0$ and $c_{0}=1$. For the first of the constituting solutions we chose $\gamma=1$, $c_{1}=0$, $c_{2}=1$, for the second $\gamma=2$, $c_{1}=1$, $c_{2}=0$.